### v1

# AttackSeqBench: Benchmarking Large Language Models’ Understanding of Sequential Patterns in Cyber Attacks

Abstract.
The observations documented in Cyber Threat Intelligence (CTI) reports play a critical role in describing adversarial behaviors, providing valuable insights for security practitioners to respond to evolving threats. Recent advancements of Large Language Models (LLMs) have demonstrated significant potential in various cybersecurity applications, including CTI report understanding and attack knowledge graph construction. While previous works have proposed benchmarks that focus on the CTI extraction ability of LLMs, the sequential characteristic of adversarial behaviors within CTI reports remains largely unexplored, which holds considerable significance in developing a comprehensive understanding of how adversaries operate. To address this gap, we introduce AttackSeqBench, a benchmark tailored to systematically evaluate LLMs’ capability to understand and reason attack sequences in CTI reports. Our benchmark encompasses three distinct Question Answering (QA) tasks, each task focuses on the varying granularity in adversarial behavior. To alleviate the laborious effort of QA construction, we carefully design an automated dataset construction pipeline to create scalable and well-formulated QA datasets based on real-world CTI reports. To ensure the quality of our dataset, we adopt a hybrid approach of combining human evaluation and systematic evaluation metrics. We conduct extensive experiments and analysis with both fast-thinking and slow-thinking LLMs, while highlighting their strengths and limitations in analyzing the sequential patterns in cyber attacks. The overarching goal of this work is to provide a benchmark that advances LLM-driven CTI report understanding and fosters its application in real-world cybersecurity operations. Our dataset and code are available at https://github.com/Javiery3889/AttackSeqBench.
Cyber Security, Cyber Threat Intelligence, Large Language Models
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Security and privacyccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Machine learning approaches
1. Introduction
Figure 1. Illustration of the difference between existing CRU-related benchmarks and our AttackSeqBench based on a real-world CTI report (Koehl and Hannon, 2020). We highlight the sequential patterns within CTI reports and investigate the capability of LLMs in comprehending and reasoning such sequential patterns.
The increasing complexity and diversity of cyber attacks have imposed a challenge to defend against Advanced Persistent Threats (APTs) (Huang et al., 2018; Duo et al., 2022). Cyber Threat Intelligence (CTI) reports, which detail the adversarial behaviors observed in cyber attacks, are crucial towards addressing emerging cyber threats. However, their unstructured nature demands extensive manual analysis (Wagner et al., 2019; Sun et al., 2023). Meanwhile, recent advancements in Large Language Models (LLMs) have demonstrated promising applications in cybersecurity, such as attack knowledge graph construction (Zhang et al., 2025b; Hu et al., 2024) and threat intelligence extraction (Rahman et al., 2023; Xu et al., 2024b). These capabilities shed new light on LLM-driven CTI Report Understanding (CRU) to support cybersecurity operations, leading to the development of various CRU-related benchmarks to assess LLMs’ capability in the field.
Existing CRU-related benchmarks primarily focus on evaluating LLMs on three aspects: (1) Classifying adversarial behaviors (Alam et al., 2023) (2) Extracting threat intelligence from unstructured text (Bhusal et al., 2024), or (3) Understanding meaningful correlations between the cyber threats and malware groups (Alam et al., 2024). While these benchmarks focus on the information extraction capabilities of LLMs, their ability to understand and reason about the sequential patterns of adversarial behavior remains largely unexplored. We refer to these sequences as attack sequences (Choi et al., 2021; Al-Sada et al., 2025), which represent the temporal order of malicious actions adversaries perform to achieve their goals. The actions which constitute an attack sequence are commonly referred to as attack patterns, characterizing adversarial behavior on a tactical level (Mavroeidis and Bromander, 2021; Siracusano et al., 2023). We illustrate the key differences between existing CRU-related tasks and analyzing attack sequences in Figure 1. Although knowledge bases (KBs) such as the MITRE ATT&CK® framework (Strom et al., 2018) document real-world adversary behaviors through defined attack patterns, analyzing the patterns individually is insufficient to fully capture the progression of cyber attacks as documented in CTI reports. The sophisticated and stealthy nature of APTs requires a comprehensive understanding of how adversaries transition between the different attack phases, which are orchestrated as an attack sequence. This raises the need to consider the sequential characteristics of a cyber attack during CRU.
To address this research gap, we delve into the suitability of LLMs in analyzing attack sequences. We highlight three key perspectives when formulating our main research task:
Perspective 1: The process of mining threat intelligence from CTI reports not only requires general text comprehension; it also necessitates expert-level cybersecurity knowledge. This requirement becomes apparent in attack sequence analysis, as it requires a deep understanding of adversarial behavior to formulate the associate attack sequence. Although LLMs demonstrate significant potential in CRU, prior research has identified limitations in their factual reliability, particularly in technical and knowledge-intensive tasks (Wang et al., 2023a; Xu et al., 2024a). In real-world cybersecurity operations, these limitations can hinder precise and timely analysis of CTI reports, ultimately impacting defensive strategies.
Perspective 2: LLMs have demonstrated remarkable capabilities across a broad range of scenarios, from traditional Natural Language Processing (NLP) tasks to more even more challenging real-world problems (Zhao et al., 2023). The recent breakthrough of slow-thinking reasoning LLMs has also shed new light on complex reasoning tasks related to coding and mathematical reasoning (DeepSeek-AI, 2025; OpenAI, 2025). Yet, the extent to which LLMs can be applied to attack sequence analysis has remained under-explored.
Perspective 3: To address the constantly evolving cyber threat landscape and the advancements of LLMs, the benchmark must be extensible to incorporate newly observed attack sequences from CTI reports. Additionally, the unstructured nature of CTI reports necessitates the benchmark to be generalizable to ensure a comprehensive and realistic evaluation of LLMs’ capabilities. Most importantly, the benchmark must be reproducible to facilitate fair and consistent evaluations across the different LLMs.
Building on these perspectives, we introduce AttackSeqBench, a novel benchmark designed for extensive investigation. Specifically, considering Perspective 1, we define three distinct Question Answering (QA) benchmarking tasks that follow the three-level structure of adversary behaviors —Tactics, Techniques, and Procedures (TTPs) —as outlined in the ATT&CK KB. This hierarchical representation allows security practitioners to represent observed attack patterns in varying levels of granularity based on pre-defined adversarial behaviors. Each tactic represents a high-level objective within a cyber attack’s lifecycle, while techniques and procedures elaborate on the specific actions performed by an adversary to achieve their objectives. Following Perspective 2, we design three benchmark settings to extensively evaluate the effectiveness and robustness of AttackSeqBench. Moreover, we leverage a diverse set of fast-thinking and slow-thinking LLMs in our experiments (Min et al., 2024), providing comprehensive insights into the capabilities and limitations of LLMs in attack sequence analysis. Catering to Perspective 3, we construct attack sequences based on real-world CTI reports. Additionally, we carefully design an automated dataset generation pipeline to generate questions in our QA dataset based on the attack sequence constructed. We summarize our contributions as follow:
  • We propose AttackSeqBench, a novel benchmark designed to systematically evaluate LLMs’ capabilities in analyzing attack sequences constructed from real-world CTI reports.
  • We introduce three QA tasks in our benchmark, each aligned with the different levels of granularity in adversarial behavior representation.
  • We conduct extensive experiments and various benchmarking settings using a diverse set of fast-thinking and slow-thinking LLMs to provide insights into their capabilities and limitations.
2. Related Work
CTI Report Understanding and LLMs. CTI refers to the collection and analysis of evidence-based information related to cyber attacks and threat actors, enabling security practitioners to gain actionable insights into the adversarial operations respond to these adversaries effectively (Wagner et al., 2019; Sun et al., 2023). A crucial source of CTI comes from CTI reports published by security vendors, these reports are typically presented as unstructured textual documents and document observations based on emerging and ongoing cyber threats (Rahman et al., 2020). CTI Report Understanding (CRU) broadly encompasses the extraction, interpretation, and reasoning of key insights within CTI reports. TTP extraction is a critical aspect in CRU, enabling security practitioners to characterize adversarial behaviors into known attack patterns defined in the ATT&CK KB (Husari et al., 2017). Some works (Legoy et al., 2020; Marchiori et al., 2023) leverage multi-label text classification to extract TTPs, while more recent research utilize Language Models (LMs) to perform TTP extraction (Rani et al., 2024). Another line of work (Shin et al., 2023) explores embedding TTPs using Global Vectors for Word Representation to capture co-occurrences. Despite considerable progress, these works tend to be costly and error-prone as adversarial behavior continues evolving.
With the breakthrough of LLMs, researchers have progressively explored their applicability within the cybersecurity domain (Zhang et al., 2024a). Within CRU, existing research (Fayyazi et al., 2024) has showcased significant capabilities of LLMs in interpreting TTPs from the ATT&CK KB, surpassing the performance of some of the fine-tuned LMs with cybersecurity data. Another line of works (Huang and Xiao, 2024; Cheng et al., 2024) utilizes LLMs to extract the relevant entities to construct knowledge graphs (KGs) based on the CTI reports. AttacKG+ (Zhang et al., 2025b) constructs adversarial behavior KGs by leveraging LLMs to map extracted entities to ATT&CK TTPs. However, the capabilities of LLMs to analyze attack sequences in CTI reports remain unexplored.
Related Benchmarks. Within the cybersecurity domain, existing work (Liu, 2023; Tihanyi et al., 2024; Garza et al., 2023) have proposed several benchmarks to evaluate LLMs’ cybersecurity knowledge. However, these works primarily rely on authoritative sources (e.g., textbooks, open standards, and example TTPs from ATT&CK KB) and do not incorporate CTI reports in their dataset. CTIBench (Alam et al., 2024) designs four CTI-related tasks, including threat actor attribution and vulnerability assessment, yet only the threat actor attribution task leverages CTI reports. CTISum (Peng et al., 2024) evaluates the ability of LLMs to summarize key facts and attack processes from CTI reports. Extending this work, SevenLLM (Ji et al., 2024) designs an additional task that focuses on extracting entities and relations. To add on, prior research (Siracusano et al., 2023) assesses LLMs’ capabilities in extracting attack patterns, along with entities and relations. In the general domain, existing works (Wang et al., 2023b; Valmeekam et al., 2023) have proposed benchmarks to evaluate the reasoning capabilities of LLMs in sequence-related tasks. Furthermore, previous research has highlighted LLMs capabilities in temporal event analysis (Zhang et al., 2024b; Chang et al., 2024). However, these works may not accurately reflect the extent to which LLMs can reason about attack sequences. In contrast, AttackSeqBench delves into the suitability of LLMs in attack sequence analysis through three tasks that requires in-depth knowledge of adversarial behavior, while emphasizing the practical aspects of CRU by utilizing real-world CTI reports to construct attack sequences in our dataset.
3. Task Definition and Dataset
3.1. Task Definition
An attack sequence $S$ models the progression of adversarial behavior in a cyber attack, characterized by the logical order of the ATT&CK tactics observed in a CTI report alongside the associated techniques and procedures. Formally, $S$ is defined as a 4-tuple:
$S=(T,E,P,O),$
where:
  1. (1)
    Tactic Sequence $T=(t_{1},\ldots,t_{n})$: An ordered list of ATT&CK tactics, which outlines the phases of a cyber attack identified in a CTI report. Note that tactic $t_{k}$ represents the $k$-th tactic.
  2. (2)
    Technique Mapping $E(t_{k})=\{e_{1,k},\ldots,e_{i_{k},k}\}$: A set of ATT&CK techniques (and/or sub-techniques) associated with tactic $t_{k}$. The overall set of techniques across all tactics in the CTI report is represented as $E=\bigcup_{k=1}^{n}E(t_{k})$.
  3. (3)
    Procedure Mapping $P(e_{j,k})=\{p_{1,j,k},\ldots,p_{m_{j,k},j,k}\}$: A set of procedures describing how technique $e_{j,k}$ is performed, the overall set of procedures across all techniques in the CTI report is represented as $P=\bigcup_{e\in E}P(e)$. Each procedure $p_{l,j,k}$ is expressed as a (subject, action, object) triplet, e.g., (The Lazarus group, use, spear-phishing emails).
  4. (4)
    CTI Outline $O=(o_{1},\ldots,o_{n})$: A tactic-centric outline of the TTPs described in the CTI report, where $o_{k}$ is a textual summary of the techniques and procedures associated with $t_{k}$.
AttackSeq-Tactic. This task tests the model’s ability to deduce a tactic $t_{k}\in S$, given a question $Q$ and four shuffled candidate tactics $C=\{c_{r}:r\in[1,4]\}$. The model must select the correct tactic $c_{l}\in C$ that corresponds to $t_{k}$.
AttackSeq-Technique. This task evaluates the model’s capability to infer a technique $e_{j,k}\in E(t_{k})$, given a question $Q$ and four shuffled candidate techniques $C=\{c_{r}:r\in[1,4]\}$. The model must select the correct technique $c_{l}\in C$ that corresponds to $e_{j,k}$.
AttackSeq-Procedure. This task challenges the model to determine the plausibility of a group of procedures $P(e_{j,k})\subseteq P$. The input is a question $Q$ and two candidate choices $C=\{c_{yes},c_{no}\}$. The model must select the correct choice $c_{l}\in C$ that indicates if $P(e_{j,k})$ is plausible. We further divide the task based on the correct answer to the question, i.e. AttackSeq-Procedure-Yes and AttackSeq-Procedure-No.
Within AttackSeqBench, we create three distinct benchmark settings (refer to Section 4.1) that have varying inputs for each benchmarking task differ to the LLMs. These benchmark settings are as follows: Zero-shot, where the input is a question $Q$ and candidate choices $C$; RAG-empowered, where the input is a question $Q$ and candidate choices $C$, along with a list of relevant ATT&CK TTPs to $Q$; Regular, where the input is a question $Q$ and candidate choices $C$, along with a masked CTI outline where the textual summary of $t_{k}$ is removed from $O$, i.e., $O_{masked}=O\setminus o_{k}$.
Figure 2. The pipeline of our dataset construction process. Given the CTI reports, we first leverage a LLM to extract the attack sequence and re-write the CTI outlines. Then we generate massive Q&A pairs with a question generator. Finally, we design the tailored evaluation criteria to enable LLM to select the appropriate Q&A pairs and construct the datasets for AttackSeqBench.
3.2. Dataset Construction
As illustrated in Figure 2, the automated dataset construction pipeline is divided into three phases: (1) Attack Sequence Construction, (2) Question Generation, and (3) Self-Refinement.
3.2.1. Attack Sequence Construction
AttacKG+ (Zhang et al., 2025b) has summarized 500 real-world CTI reports from multiple intelligence sources into tactic-centric outlines, and extracted the associated TTPs by leveraging a LLM-based framework. To construct attack sequences that resemble real-world adversarial behavior, we exclude CTI outlines which contains less than two ATT&CK tactics as they are unlikely to detail attack patterns observed in cyber attacks (Zhang et al., 2025b). Afterwards, we combine the CTI outlines and the associated TTPs from 408 filtered CTI reports to form attack sequences.
3.2.2. Question Generation
We leverage the question generation capabilities of LLMs (Alam et al., 2024; Zhang et al., 2024b; Mucciaccia et al., 2025) to formulate the questions in our QA dataset based on the attack sequences constructed. More specifically, we propose an answer-aware question generation approach, by utilizing GPT-4o (OpenAI, 2024b) to generate a question based on each TTP present with respect to the scope of each benchmarking task for a given attack sequence $S$. To generate more relevant questions, we leverage LLMs’ in-context learning (ICL) ability (Dong et al., 2024) by providing the associated outline $O$ as contextual knowledge, and include few-shot examples based on the example procedures in the ATT&CK KB (refer to Section C.1). For the MCQ tasks, we adopt a rule-based approach to randomly select three distractors from the attack sequence and ATT&CK KB. Due to the limited tactics present in the ATT&CK KB, we randomly sample an adjacent tactic present in the attack sequence i.e., $t_{k+1}$ or $t_{k-1}$, along with two randomly selected tactics from the ATT&CK KB. For AttackSeq-Technique, we select the distractors with reference to the STARC annotation framework (Berzak et al., 2020). The three distractors are as follow: (1) Similar to AttackSeq-Tactic, the first distractor is associated with the same tactic but not supported by the CTI report, i.e., $e_{i,k}\in E(t_{k})$, indicating a plausible misunderstanding of the correct technique $e_{j,k}$; (2) The second distractor refers to a technique that is supported by the CTI report but belongs to another tactic $t_{j}\neq t_{k}$; (3) The third distractor refers to a technique that is not supported by the CTI report and belongs to a randomly chosen tactic.
For the boolean questions, AttackSeq-Procedure, we first generate the questions with correct answer ”Yes”, i.e., for each group of procedures $P(e_{j,k})\subseteq P$. To create the negative questions in AttackSeq-Procedure-No, we randomly sample 70% of the AttackSeq-Procedure-Yes questions from our final QA dataset and prompt the LLM to generate two types of negative questions based on each sampled question (refer to Table 15). Inspired by existing work (Rajpurkar et al., 2018), the two types of negative questions are: (1) We negate the temporal prepositions, i.e., changing ”before” to ”only after” and/or ”after” to ”only before”, such that the sequence of adversarial behavior is incorrect; (2) We replace $P(e_{j,k})$ in the question’s body with a group of procedures associated with a randomly sampled technique that is not supported by CTI report.
3.2.3. Self-Refinement
Inspired by the Self-Refine framework (Madaan et al., 2023), we adopt an iterative three-step approach using GPT-4o to improve the quality of the generated questions. We first assess the quality of the question based on the Answerability aspect (Fu et al., 2024), filtering out the questions that do not satisfy this criteria. More specifically, we check if there is direct evidence in the CTI outline that supports the correct answer, while clearly standing out as the best answer choice. Within this aspect, we also check if the correct answer can be inferred even if the associated summary to the correct answer’s tactic is removed from the outline (Huang et al., 2019), i.e., using $O_{masked}$ to answer the question. Based on the filtered answerable questions, we prompt the LLM to consider the quality of the question and its answer choices based on the following five aspects:
  • Clarity. We check if the question precise and unambiguous. More importantly, we also ensure that question avoid directly mentioning the correct answer such that the inference is required.
  • Logical. We check if the sequence described in the question follow the order of tactics present in the attack sequence.
  • Relevance. We check if the TTPs described in the question directly relate to the attack sequence.
  • Consistency. We check if the question is consistent with the associated TTP that is used for question generation.
  • Answer Consistency. We check if the question can be fully answered by the correct answer, without any contradictions and inconsistencies.
Based on these aspects, we prompt the LLM to give a numerical score out of 5 and select the questions that have a full score of 25. This filter-feedback-refine process is repeated once more to ensure better quality of questions. Refer to Appendix C.2 for the detailed prompts.
Table 1. Number of questions and proportional decrease in each benchmarking task before and after Self-Refinement.
Task Num. (Before) Num. (After) % Decrease
AttackSeq-Tactic 2,158 1,697 -21.36
AttackSeq-Technique 2,937 1,917 -34.73
AttackSeq-Procedure 4,642 2,635 -43.24
3.2.4. Dataset Statistics
The number of generated questions for each benchmark task, along with the corresponding percentage decrease, is shown in Table 1. More details of the dataset, such as the distribution of ATT&CK tactics and techniques associated with the questions, is in Appendix A.1.
3.3. Dataset Evaluation
3.3.1. Human Evaluation
We ask two postgraduate students majoring in computer science, who are knowledgeable about cybersecurity and the MITRE ATT&CK framework, to assess the quality of questions in AttackSeqBench. We randomly sample 35 questions from each of the three benchmarking tasks, including the AttackSeq-Procedure-No subtask, resulting in a total of 140 questions for manual evaluation. Each annotator evaluates the six aspects described in Section 3.2.3 using a 5-point Likert scale, where higher scores indicate better alignment (detailed definitions of each aspect are in Appendix A.2). Based on Table 8, the average accuracy score of annotators over the three tasks is 0.625, suggesting that questions in AttackSeqBench are notably challenging for domain experts. Furthermore, the human evaluation results indicate that the generated questions closely satisfy with each aspect of the evaluation criteria, with an average score of 4.620/5 across the six aspects. In particular, the Logical and Relevance aspects suggests that LLMs are able to construct questions that closely align with attack sequences described in CTI reports.
3.3.2. Automatic Evaluation
To alleviate the laborious task of manual human annotation, recent works (Zheng et al., 2023; Yao et al., 2024) have shown considerable effectiveness of LLM-as-a-Judge frameworks in aligning with human preferences within domain-specific fields, including cybersecurity (Xu et al., 2024b). We incorporate G-Eval (Liu et al., 2023), a Chain-of-Thought (CoT) (Wei et al., 2022) and form-filling paradigm, to systematically assess the quality of generated questions. Similar to Section 3.3.1, we carefully design prompts for each of the six aspects using a 5-point Likert scale (refer to Appendix C.3 for the detailed prompt). More specifically, we instruct GPT-4o to refer to the evaluation criterion’s definition, along with description of the correct answer from the ATT&CK KB, and rate the question out of 5 for each aspect. Based on Table 9, we observe that Logical and Relevance are the highest rated aspects, reinforcing the LLM’s ability to construct questions that follow the logical order of attack sequences. While the automatic evaluation scores are lower than the human evaluation scores, this implies that it is more challenging for LLMs to answer the questions correctly as compared to domain experts.
4. Experiment
Figure 3. The illustration of three benchmark settings. (a): Regular setting, (b): Zero-shot setting, and (c): RAG-empowered setting.
4.1. Benchmark Settings
As illustrated in Figure 3, we elaborate on the three proposed benchmark settings, Regular, Zero-Shot and RAG-Empowered. The detailed prompts for each setting is in Appendix C.4.
Regular Setting. Inspired by existing works (Ma et al., 2023; Jin et al., 2024), we propose this benchmark setting to assess the impact of contextual information in analyzing attack sequences. We adopt an ICL-based approach by carefully designing a well-constructed prompt using CoT. Within the prompt, we first instruct the LLM to analyze the attack sequence described in the masked CTI outline associated with the question, and subsequently deduce the correct answer based on the given attack sequence.
Zero-Shot Setting. Aside from excelling in semantic understanding, previous works have also demonstrated LLMs’ impressive zero-shot reasoning ability across various downstream tasks (Hou et al., 2024; Kojima et al., 2022). To evaluate the LLMs’ capabilities in zero-shot attack sequence analysis, we omit the masked CTI outline from the prompt in the Regular setting. Instead, we instruct the LLM to select the most plausible answer choice based on the attack sequence described in the question’s body.
RAG-Empowered Setting. Retrieval Augmented Generation (RAG) has gained traction due to its ability to enhance LLMs’ reliability and mitigate hallucinations (Zhang et al., 2025a). We introduce this benchmark setting to assess the impact of providing the relevant ATT&CK TTPs in the question’s body as contextual information, by leveraging a retrieval-based approach and the ATT&CK KB. We first crawl the description and the example procedures of each TTP from the Enterprise ATT&CK Matrix  [^1] https://attack.mitre.org/matrices/enterprise/. Next, we split and embed the text chunks alongside their associated ATT&CK tactic and technique as metadata. Afterwards, we retrieve relevant chunks based on the question’s body and extract the associated ATT&CK TTPs from the metadata. Lastly, we instruct the LLM to utilize the retrieved TTPs and deduce the correct answer. By doing so, we enable the LLM to focus on determining the correct attack sequence rather than interpreting the TTPs in the question.
Table 2. Performance (Accuracy) comparison of various open-source and proprietary LLMs (both fast-thinking and slow-thinking reasoning LLMs) under three settings (i.e., Regular, Zero-Shot, RAG) on three datasets (i.e., Tactic, Technique, Procedure). For each column, bold number indicates the best performance, and underlined number corresponds to the second best performance.
LLMs AttackSeq-Tactic AttackSeq-Technique AttackSeq-Procedure
Regular Zero-Shot RAG Regular Zero-Shot RAG Regular Zero-Shot RAG
Fast-thinking LLMs
[…]
Qwen-2.5-7B-Instruct 0.5121 0.4903 0.4761 0.6693 0.6568 0.6067 0.6584 0.5184 0.4941
Llama-3.1-8B-Instruct 0.4744 0.5085 0.4926 0.6260 0.6333 0.5827 0.6577 0.5328 0.5230
ChatGLM-4-9B-Chat 0.4885 0.4979 0.5009 0.6275 0.6109 0.6030 0.641 0.5408 0.5131
Llama-3.3-70B-Instruct 0.6588 0.5551 0.5681 0.7058 0.6797 0.7037 0.6903 0.5469 0.5279
Qwen-2.5-72B-Instruct 0.5793 0.5863 0.5657 0.5430 0.7162 0.6959 0.7188 0.6285 0.6030
GPT-4o-mini 0.6517 0.6005 0.5692 0.7387 0.7058 0.7021 0.6968 0.5491 0.5340
GPT-4o 0.6093 0.5740 0.5787 0.6755 0.6995 0.7188 0.7359 0.6755 0.6353
Slow-thinking Reasoning LLMs
[…]
DeepSeek-R1-Distill-Qwen-32B 0.5421 0.5698 0.5504 0.5879 0.5816 0.5816 0.6945 0.6258 0.5852
QWQ-32B-Preview 0.5345 0.3377 0.4638 0.5112 0.3918 0.5342 0.7036 0.5696 0.5457
GPT-o3-mini 0.4643 0.5445 0.5215 0.5373 0.5915 0.5822 0.6854 0.6877 0.6459
4.2. Experimental Setting
4.2.1. Implementation Details
To demonstrate the robustness of AttackSeqBench, we evaluate eight fast-thinking LLMs: Mistral-7B-Instruct-v0.3, Qwen-2.5-7B-Instruct, Llama-3.1-8B-Instruct, ChatGLM-4-9B-Chat, Llama-3.3-70B-Instruct, Qwen-2.5-72B-Instruct, GPT-4o-mini and GPT-4o. Additionally, we assess four slow-thinking reasoning LLMs: DeepSeek-R1-Distill-Llama-8B, DeepSeek-R1-Distill-Qwen-32B, QWQ-32B-Preview and GPT-o3-mini. For the RAG-empowered benchmark setting, we utilize a hybrid retriever due to its considerable performance in existing work (Zhang et al., 2024b). We combine Okapi BM25 (Robertson and Zaragoza, 2009) as the sparse retriever and a dense retriever based on BGE-EN-ICL (Li et al., 2024) (shortened as BGE) and Chroma DB  [^2] https://www.trychroma.com/, along with a re-ranker. The detailed implementation is in Appendix B.1.
4.2.2. Evaluation Metrics
We evaluate the Accuracy (abbreviated as Acc in the following sections) for all three tasks in AttackSeqBench (Liu et al., 2024; Zhang et al., 2024b; Alam et al., 2024). Specifically, given the total number of questions $M$ and the number of questions $n$ with the correct answer choice chosen by the LLM, the accuracy can be calculated as $Acc=n/M$.
4.3. Performance Comparison
We conduct comprehensive analysis using eight fast-thinking LLMs and four slow-thinking reasoning LLMs in Table 2 and elaborate on three key findings: (1) None of the LLMs clearly outperforms other LLMs across all the tasks. While GPT-4o demonstrate strong performance in AttackSeq-Technique and AttackSeq-Procedure, its performance falls behind Llama-3.3-70B-Instruct and GPT-4o-mini in AttackSeq-Tactic. This implies the complexity and the diversity of our three sub-tasks to some extent. (2) Among these three tasks, we observe that LLMs collectively demonstrate the lowest overall performance for the AttackSeq-Tactic task. This implies LLMs struggle to determine the high-level objectives in an attack sequence, even with the CTI outline given in the context. This may arise due to the abstract nature of the descriptions of tactics which causes the LLM to misinterpret the correct tactic. (3) The AttackSeq-Procedure task yields the highest performance results across all the LLMs. This suggests that LLMs are able to leverage its procedural knowledge to deduce the plausibility of a group of procedures in an attack sequence. However, this may also be attributed by the LLMs’ tendency to favor ”yes” in boolean questions when given insufficient context (Wen et al., 2024), which causes high accuracy for AttackSeq-Procedure-Yes. We explore this phenomenon further in Section 4.4.1.
4.4. Model Analysis
Table 3. Performance comparison of four representative LLMs between AttackSeq-Procedure-Yes and AttackSeq-Procedure-No sub-tasks in each benchmark setting. For each column, the bold value indicates the best performance.
LLMs AttackSeq-Procedure-Yes AttackSeq-Procedure-No
Regular Zero-Shot RAG Regular Zero-Shot RAG
Llama-3.1-8B 0.7899 0.9518 0.9460 0.5432 0.1700 0.1565
GPT-4o 0.7146 0.9403 0.9542 0.7542 0.4462 0.3591
R1 (Llama-8B) 0.8332 0.9526 0.9477 0.4341 0.1161 0.1062
GPT-o3-mini 0.5789 0.7760 0.8242 0.7776 0.6112 0.4915
4.4.1. Impact of Contextual Information
To comprehensively investigate the impact of the contextual knowledge in the given CTI outlines for analyzing attack sequences in LLMs, we compare the performance of LLMs between the Regular and Zero-Shot benchmark settings. Based on in Table 2, we observe that both fast-thinking and slow-thinking reasoning LLM’s accuracy generally increases across each task in the  Regular setting. Among the fast-thinking LLMs, the performance of models with smaller number of parameters, e.g., Llama-3.1-8B-Instruct (shortened as Llama-3.1-8B), have minimal or no improvement for AttackSeq-Tactic and Attack-Technique tasks. This indicates that the baseline LLMs with smaller parameters may not fully leverage the capabilities of ICL within these benchmarking tasks. Similarly for slow-thinking reasoning models, DeepSeek-R1-Distill-Llama-8B (shortened as R1 (Llama-8B)) has no improvement in the Regular setting as compared to the Zero-Shot setting. On the other hand, Llama-3.3-70B-Instruct achieves comparable performance against the closed source models, i.e., GPT-4o and GPT-4o-mini. Among the three tasks, AttackSeq-Procedure has the highest improvement in LLMs from the Zero-Shot setting to the Regular setting. Based on Table 3, we observe a significant increase in the AttackSeq-Procedure-No sub-task. This suggests that the contextual information found in the masked CTI outline is essential in justifying procedures that are unlikely to happen.
Table 4. Performance comparison between three embedding models. Eeb. M. refers to the embedding model used, while Tac, Tec and Pro refers to the three tasks in AttackSeqBench (i.e., Tactic, Technique, and Procedures) respectively. For each row, the bold value indicates the best performance.
#Eeb. M. Task Llama-3.1-8B GPT-4o R1 (Llama-8B) GPT o3-mini
BGE Tac 0.4926 0.5787 0.4532 0.5215
Tec 0.5827 0.7188 0.5138 0.5822
Pro 0.5230 0.6353 0.4968 0.6459
OpenAI Tac 0.4915 0.4814 0.4632 0.5333
Tec 0.5790 0.7178 0.5039 0.5962
Pro 0.5195 0.6330 0.4979 0.6353
ATT&CK- BERT Tac 0.4985 0.5622 0.4590 0.5292
Tec 0.5790 0.7100 0.5018 0.5874
Pro 0.5245 0.6209 0.4953 0.6398
4.4.2. Effectiveness of RAG Strategies
We analyze the capabilities of the RAG-empowered setting based on the results of AttackSeqBench in Table 2. Within the set of fast-thinking LLMs, GPT-4o achieves exceptional performance when leveraging RAG in AttackSeq-Tactic and AttackSeq-Technique tasks. Moreover, the LLMs exhibit the best performance in the AttackSeq-Technique among the three tasks in Table 2. This indicates that LLMs are able to leverage the retrieved TTPs to determine the correct technique. However, the RAG-empowered setting does not consistently yield effective results, as models that have less than or equal to 9B parameters have no improvement as compared to the Zero-Shot setting. This suggests that the information retrieved, i.e., the semantically similar ATT&CK TTPs based on the question’s body, is insufficient for smaller models. Furthermore, we evaluate the performance of four representative LLMs using two additional embedding models in Table 4. The two models are: OpenAI’s text-embedding-3-large (OpenAI, 2024c); ATT&CK BERT (Abdeen et al., 2023) which is an open-source embedding model fine-tuned for cybersecurity. Based on Table 4, BGE acheives the best performance across the three benchmarking tasks.
Figure 4. Performance and model size comparison of nine representative LLMs in the AttackSeq-Tactic task. The size of bubble corresponds to inference time, where horizontal lines denote fast-thinking LLMs and vertical lines indicate slow-thinking reasoning LLMs.
4.4.3. Computational Complexity Analysis
As illustrated in Figure 4, we extensively compare the performance, model size and inference cost of nine open-source LLMs in the AttackSeq-Tactic task. We have the following observations: (1) In general, these LLMs exhibit the performance trend in sequential attack pattern understanding which aligns with the scaling law (Kaplan et al., 2020). (2) Comparing the performance of LLMs with different thinking types, we observe that slow-thinking LLMs cannot significantly and consistently outperform the other. One possible reason is that given the question and masked CTI outlines, fast-thinking LLMs can already deduce the correct answer without any elaborate reasoning. (3) Regarding the inference cost, we notice that slow-thinking LLMs do not incur more reasoning costs compared to fast-thinking LLMs. This prompts an intuitive insight: the small-scale slow-thinking reasoning LLM fine-tuned using AttackSeqBench may enhance its capability in attack sequence analysis significantly while maintaining the lower reasoning cost.
4.4.4. Sampling Parameters Analysis
We assess the impact of two sampling parameters on the performance of LLMs in AttackSeq-Tactic under the Regular setting, namely temperature and maximum output tokens as illustrated in Figure 5. We observe that increasing the temperature at different three intervals has minimal improvements for Llama-3.3-70B-Instruct, while resulting in negative performance for the remaining three LLMs. This indicates that the most optimal performance occurs when the temperature is kept at 0.0. On the other hand, we observe a considerable improvement when increasing the number of tokens for slow-thinking reasoning models. The accuracy of Deepseek-R1-Distill-Llama-8B improved by 8.46% as the maximum output tokens increased from 1024 to 4096, while Deepseek-R1-Distill-Qwen-32B achieve a 24.9% improvement for the same increment. However, further increasing the token limit to 8192 yields diminishing returns for both models. This highlights the importance of carefully tuning the maximum output tokens parameter to optimize performance in slow-thinking models while considering the associated costs incurred (Wang et al., 2024).
Figure 5. Parameter analysis on temperature (a) and maximum output tokens (b) in AttackSeq-Tactic task using the Regular setting, we perform analysis on four representative LLMs using different colors to represent them.
Figure 6. The performance comparison of four representative LLMs on each tactic in each benchmarking task.
Table 5. Reasoning path comparison between slow-thinking reasoning and fast-thinking LLMs in Zero-Shot setting based on the AttackSeq-Procedure task in Zero-Shot setting.
Question: After delivering spearphishing emails with spyware disguised as a PDF file to gain initial access, is it likely the authors of Hacking Team’s spyware established communication with the identified C&C servers before using VMProtect for obfuscation to evade detection? Answer Choices: A) Yes  B) No (Correct Answer)
DeepSeek-R1-Distill-Qwen-32B Qwen-2.5-72B-Instruct
First, spearphishing emails are emails that look legitimate, which are actually malicious. In this case, the PDF is a disguise for spyware. So the first step is the delivery of the malicious payload via email. (Knowledge Retrieval) Once the victim opens the PDF, the spyware is executed. The next step in an attack chain is usually establishing communication with a command and control (C&C) server. Now, the question mentions VMProtect being used for obfuscation. However, sometimes obfuscation can be applied after deployment to maintain persistence or evade detection during operation. (Breakdown of Attack Sequence) But in the context of the attack sequence described, the spyware has already been delivered and executed. The next logical step would be for the spyware to reach out to the C&C server to report back and receive instructions. (Self-Correction) If the attackers were to use VMProtect at this point, it might be to re-obfuscate the spyware or to protect it further, but that seems a bit out of order. (Reasoning using Attack Sequence) So the answer would be No. (Correct Conclusion) The sequence includes: Delivering spearphishing emails with spyware disguised as a PDF file to gain initial access. Establishing communication with C&C (Command and Control) servers. Using VMProtect for obfuscation to evade detection. (Breakdown of Attack Sequence) The spearphishing emails with a PDF file are used to gain initial access to the target system. (T1566.002: Spearphishing Attachment). Once the initial access is gained, the malware typically establishes communication with the C&C servers to receive further instructions or exfiltrate data. (T1071: Application Layer Protocol). VMProtect is a tool used to obfuscate the malware to evade detection by security software. (T1027: Obfuscated Files or Information). (Knowledge Extraction) The logical sequence of events would be: 1. **Initial Access:**, 2. **C&C Communication:**, 3. **Obfuscation:**. (Incorrect Attack Sequence Order) Based on the logical sequence of TTPs, it is likely that the authors of Hacking Team’s spyware established communication with the C&C servers before using VMProtect for obfuscation to evade detection. Final Answer: Yes (Incorrect Conclusion)
4.5. In-depth Analysis
4.5.1. Impact of TTP Temporal Position
To better understand the relationship between LLMs and attack sequences, we conduct fine-grained analysis on the performance of four representative LLMs under the Regular benchmark setting and across the ATT&CK tactics (refer to Figure 6). Additionally, we identify four categories to group the tactics into overarching attack phases as shown in Figure 6: Initial Intrusion, Exploitation, Stealth Expansion, Objective Orchestration. Our findings are: (1) Based on the standard deviation range across various tactics in each LLM, we observe that the range is the largest for AttackSeq-Tactic, indicating that the LLM’s performance largely varies for this task across different tactics. In contrast, the SD range is the smallest for AttackSeq-Procedure which suggests more stability in the LLMs’ performance. This highlights the challenge for LLMs to determine the correct tactics in attack sequences, as compared to procedural knowledge needed to correctly determine techniques and procedures. (2) GPT-4o consistently outperforms other models across most tasks, particularly in the in the Exploitation and Stealth Expansion phases. On the other hand, Llama-3.1-8B-Instruct and GPT o3-mini achieve comparable performance against GPT-4o in Initial Intrusion and Objective Orchestration phases, especially at the start and end of the sequence (i.e., Reconnaissance and Impact respectively). This inconsistency suggests that TTPs within certain ATT&CK tactics are more challenging for the LLM to interpret, likely due to the rarely observed tactics in real-world CTI reports. More detailed distribution of the performance can be found in Section B.2 of the Appendix.
4.5.2. Inference Path Analysis on Slow-thinking Reasoning LLMs
We perform a case study based on a question from AttackSeq-Procedure-No and compare the difference in the reasoning process in the output of two open-source LLMs: DeepSeek-R1-Distill-Qwen-32B and Qwen-2.5-72B-Instruct in the Zero-Shot setting. We choose not to use GPT-o3-mini in our analysis as the Batch API does not return the reasoning tokens. Based on Table 5, we observe that both models are able to breakdown the attack sequence and extract the relevant attack patterns. The reasoning output tokens suggest that DeepSeek-R1-Distill-Qwen-32B is able to retrieve relevant knowledge based on the attack sequence described in the question. However, Qwen-2.5-72B-Instruct outputs the incorrect order of attack patterns as stated in the question. Furthermore, it highlights the slow-thinking reasoning LLMs’ ability to perform Self-Correction (Kumar et al., 2024), enabling the model to elicit stronger reasoning capabilities and justify why the attack sequence stated in the question is incorrect. Notably, the slow-thinking reasoning LLM considers both incorrect and correct attack sequences before deciding on a conclusion, highlighting the potential in slow-thinking reasoning LLMs in zero-shot tasks.
5. Outlook and Conclusion
Outlook. AttackSeqBench extends beyond existing CRU-related benchmarks by emphasizing sequential patterns in cyber attacks, opening up numerous research avenues for related researchers. For instance, AttackSeqBench can be seamlessly utilized for the complex attack path modeling task as we provided in Section 4.5.1. Moreover, AttackSeqBench can also be leveraged to recognize the attack intention due to its temporal characteristic. Inspired by the ubiquitous causality in real-world scenarios, we believe that integrating the logical attack sequence with the causal inference is a promising research direction, discovering the critical causal relationships within the attack sequences and enabling more accurate predictions of the attackers’ potential action. Given the predicted attack intentions and next action, applying our AttackSeqBench to optimize the defense strategy is not only intuitive but also reasonable. Finally, the constructed dataset can be reformulated as the instruction-tuning data, facilitating the post-training phase of future security-related LLMs.
Conclusion. In this work, we introduced AttackSeqBench, a novel benchmark designed to assess the LLMs’ ability to analyze attack sequences from real-world CTI reports. We propose an automated dataset construction pipeline, enabling our benchmark to be easily extended as cyber threats continue to evolve. Our benchmark incorporates three tasks that encompasses the varying aspects of adversarial behavior, allowing a comprehensive analysis of the LLMs’ understanding and reasoning capabilities in attack sequence analysis. We performed extensive experiments on a diverse set of fast-thinking and slow-thinking reasoning LLMs, offering crucial insights into the abilities and limitations of LLMs in understanding attack sequences. Our work opens up a meaningful new direction towards LLM-driven CRU in real-world cybersecurity operations.
Appendix A Dataset
A.1. Dataset Distribution
The distribution of questions in AttackSeqBench is shown in Table 6, we construct an almost equal proportions of questions in each task to allow fair evaluation in our benchmark.
Based on Figure 7, we observe that the top three most frequent tactics occur in the middle of attack sequences, while the bottom two tactics occurs at the start and end of the attack sequence. The most frequent technique belongs to the Command and Control tactic, this is likely due to the common operations of APTs. Similarly, the most frequent ATT&CK technique is T1071-Application Layer Protocol  [^3] https://attack.mitre.org/techniques/T1071/, which is associated with the Command and Control tactic.
Figure 7. Visualization of the distribution of top 10 tactics and techniques in our AttackSeqBench.
A.2. Dataset Evaluation
Inspired by existing work (Fu et al., 2024), we evaluate the quality of our QA dataset based on six dimensions below:
  • Answerability: We check if there is direct evidence in the CTI outline that supports the correct answer, while clearly standing out as the best answer choice. Within this aspect, we also check if the correct answer can be inferred even if the associated summary to the correct answer’s tactic is removed from the CTI outline.
  • Clarity. We check if the question precise and unambiguous. More importantly, we also ensure that question avoid directly mentioning the correct answer such that the inference is required.
  • Logical. We check if the sequence described in the question follow the order of tactics present in the attack sequence.
  • Relevance. We check if the TTPs described in the question directly relate to the attack sequence.
  • Consistency. We check if the question is consistent with the associated TTP that is used for question generation.
  • Answer Consistency. We check if the question can be fully answered by the correct answer, without any contradictions and inconsistencies.
We give detailed instructions to the two human evaluators by defining each score in the 5-point Likert scale for each aspect (refer to Table 7). The same Likert scale definitions are used during automatic evaluations in the prompt. The detailed results of human evaluation and automatic evaluation are shown in Table 6 and Table 7, respectively. While the automatic evaluation results are lower than human evaluation, the human evaluation shows that most QA pairs in the dataset satisfy the requirements of all aspects. This suggests that automatic evaluation is still limited in knowledge-intensive domains such as in cybersecurity. Note that for AttackSeq-Procedure-No, we only assess the questions in the following four aspects: Answerability; Clarity, Consistency; Answer Consistency.
A.3. Ethics Statement
This study adheres to ethical research practices by focusing on publicly available CTI reports and ensuring that no proprietary or sensitive information is used in the benchmark construction. The automated dataset generation process is designed to maintain the integrity and accuracy of adversarial behavior sequences without fabricating or misrepresenting cyber threats. Furthermore, human evaluation is conducted with careful consideration of annotator expertise and potential biases, ensuring fairness and reliability in assessment. Our work aims to contribute to the cybersecurity community by advancing LLM capabilities while upholding ethical standards in data usage, evaluation, and responsible AI development.
Appendix B Experiments
B.1. Baseline LLMs
To demonstrate the robustness of our proposed datasets, we evaluate eight fast-thinking LLMs and four slow-thinking reasoning LLMs on our benchmark tasks. We leverage vLLM (Kwon et al., 2023) to run all open-source LLMs locally, utilizing two Nvidia H100 GPUs. For the proprietary LLMs, we utilize OpenAI’s Batch API  [^4] https://platform.openai.com/docs/guides/batch to reduce inference cost and perform inference in batches. In our experiments, we set the following sampling parameters while keeping the default value for the remaining parameters: temperature to 0, maximum output tokens to 2048, and top_p to 1. Below is the list of LLMs used in our experiments:
Fast-thinking LLMs:
  • Mistral-7B-Instruct-v0.3 Mistral-7B-Instruct-v0.3 (Jiang et al., 2023) is an instruct fine-tuned version of the Mistral-7B-v0.3 which exhibits extended vocabulary and supports the v3 tokenizer.
  • Qwen2.5-7B-Instruct Qwen2.5-7B-Instruct (Team, 2024a) is one of the advanced Qwen series LLMs. It exhibits better instruction following and long text generating capability.
  • LLaMa-3.1-8B-Instruct LLaMa-3.1-8B-Instruct (Grattafiori et al., 2024) is an instruction-tuned LLM from Meta, balancing performance and efficiency for textual understanding tasks.
  • LLaMa-3.2-3B-Instruct LLaMa-3.2-3B-Instruct (Grattafiori et al., 2024) is a lightweight 3B version of LLaMA-3.2, prioritizing the rapid response and the deployment on low-resource devices.
  • Qwen2.5-72B-Instruct Qwen2.5-72B-Instruct (Team, 2024a) is one of the latest LLMs from Qwen and serves as the biggest Qwen-2.5 series LLM with 72B parameters.
  • LLaMa-3.3-70B-Instruct LLaMa-3.3-70B-Instruct (Grattafiori et al., 2024) is an auto-regressive language model which is instruction-tuned in 70B with supervised fine-tuning (SFT) and reinforcement learning with human feedback (RLHF).
  • GPT-4o GPT-4o (OpenAI, 2024b) is one of the most advanced closed-source LLMs, which is a multi-lingual and multi-modal language model developed and functions well in real-time processing  [^5] To assess its superiority, we leverage the most recent GPT-4o-2024-11-20 and GPT-4o-mini-2024-07-18 (at the time point when we did this work) as the benchmark evaluator, respectively..
  • GPT-4o-mini gpt-4o-mini (OpenAI, 2024a) is a compact variant of the GPT-4o (OpenAI, 2024b), which is optimized for faster inference and cost efficiency 5.
Slow-thinking reasoning LLMs:
  • DeepSeek-R1-Distill-Llama-8B DeepSeek-R1-Distill-Llama-8B (DeepSeek-AI, 2025) is fine-tuned from the open-source Llama-3.1-8B with 800k samples curated with DeepSeek-R1, aiming to equip smaller models with reasoning capabilities like DeepSeek-R1.
  • DeepSeek-R1-Distill-Qwen-32B DeepSeek-R1-Distill-Qwen-32B (DeepSeek-AI, 2025) denotes the distilled version from the well-trained Qwen2.5-32B without any Reinforce Learning stage.
  • QWQ-32B-Preview QWQ-32B-Preview (Team, 2024b) is a preview release which gives LLM time to ponder, to question, and to reflect, enabling the deeper insight into complex problems.
  • GPT-o3-mini GPT-o3-mini (OpenAI, 2025) is designed with a focus on enhancing LLMs’ reasoning capabilities. It leverages the Chain of Thought (CoT) to break down complex problems into several simpler steps to achieve this objective.
The retrievers in our experiments are built using LlamaIndex (Liu, 2022). We set the number of retrieved chunks $u$ and its size $l$ to be $3$ and $512$ respectively.
B.2. TTP Temporal Position Analysis
We conduct a detailed breakdown of the performance in four indicative LLMs based on each ATT&CK tactic associated to the questions in AttackSeqBench. Based on Table 10, the mean and standard deviation (SD) for each tactic and between each benchmark task is shown. We observe that the overall highest standard deviation occurs within the AttackSeq-Tactic task for all models, indicating that it is challenging for LLMs to correctly answer questions in AttackSeq-Tactic.
Table 10. Results of performance of four representative models on all the 14 tactics, and the associative statistics of mean and standard deviation (SD).
Tactic
Reconna-
issance
Resource
Devel.
Initial
Access
Execu. Persist.
Privilege
Escala.
Defense
Evasion
Credent.
Access
Discov.
Lateral
Move.
Collec.
Cmd. &
Control
Exfiltr. Impact Mean SD
Tactic Llama-3.1-8B 0.6170 0.3077 0.4406 0.5573 0.4260 0.3023 0.3867 0.4211 0.2551 0.5778 0.4390 0.6444 0.5915 0.4194 0.4561 0.1239
GPT-4o 0.6383 0.5000 0.4368 0.7470 0.5207 0.4884 0.5430 0.5000 0.7245 0.7111 0.6951 0.7155 0.7042 0.6774 0.6144 0.1095
R1 (Llama-8B) 0.4681 0.2692 0.3448 0.5455 0.3609 0.2791 0.2188 0.3421 0.2857 0.4667 0.3415 0.7238 0.5070 0.3548 0.3934 0.1350
GPT-o3-mini 0.5319 0.4615 0.4269 0.5635 0.3018 0.1395 0.3711 0.3289 0.4490 0.5111 0.3537 0.7071 0.493 0.6774 0.4512 0.1499
Mean 0.5638 0.3846 0.4123 0.6033 0.4024 0.3023 0.3799 0.3980 0.4286 0.5667 0.4573 0.6977 0.5739 0.5323 - -
SD 0.0786 0.1132 0.0454 0.0961 0.0938 0.1434 0.1325 0.0792 0.2149 0.1066 0.1643 0.0362 0.0971 0.1697 - -
Technique Llama-3.1-8B 0.6556 0.4878 0.5327 0.5890 0.7324 0.4667 0.5914 0.5625 0.6839 0.5179 0.6557 0.6884 0.7500 0.7241 0.6170 0.0943
GPT-4o 0.6222 0.6341 0.6262 0.6568 0.7254 0.6444 0.6512 0.6000 0.7513 0.6429 0.6148 0.7329 0.7632 0.6897 0.6682 0.0541
R1 (Llama-8B) 0.5667 0.4634 0.4393 0.5551 0.5775 0.4444 0.4651 0.5250 0.6321 0.3929 0.5246 0.6027 0.7368 0.4828 0.5292 0.0911
GPT-o3-mini 0.5222 0.5366 0.4836 0.5085 0.6241 0.4889 0.4333 0.3875 0.5389 0.6250 0.5328 0.6473 0.7237 0.6552 0.5505 0.0933
Mean 0.5917 0.5305 0.5205 0.5774 0.6649 0.5111 0.5353 0.5188 0.6516 0.5447 0.5820 0.6678 0.7434 0.6380 - -
SD 0.0591 0.0755 0.0802 0.0624 0.0764 0.0907 0.1031 0.0927 0.0896 0.1153 0.0638 0.0557 0.0170 0.1072 - -
Procedure Llama-3.1-8B 0.6634 0.5319 0.6489 0.6552 0.6667 0.5556 0.6809 0.6331 0.6987 0.6633 0.6027 0.6606 0.6906 0.6140 0.6404 0.0489
GPT-4o 0.7030 0.6170 0.8351 0.7586 0.7469 0.6852 0.7325 0.7554 0.7067 0.7041 0.6712 0.7515 0.777 0.7193 0.7260 0.0523
R1 (Llama-8B) 0.6535 0.5957 0.6064 0.6681 0.6728 0.6111 0.6474 0.5540 0.6027 0.5918 0.5023 0.6485 0.6547 0.5439 0.6109 0.0510
GPT-o3-mini 0.7030 0.7234 0.8085 0.6853 0.6852 0.7037 0.6991 0.6835 0.6640 0.6633 0.6256 0.6707 0.7338 0.5439 0.6852 0.0585
Mean 0.6807 0.6170 0.7247 0.6918 0.6929 0.6389 0.6900 0.6565 0.6680 0.6556 0.6005 0.6828 0.7140 0.6053 - -
SD 0.0260 0.0796 0.1139 0.0462 0.0368 0.0684 0.0355 0.0848 0.0473 0.0467 0.0714 0.0467 0.0530 0.0829 - -
B.3. Error Analysis
To better understand the limitations of LLMs in analyzing attack sequences, we examine a specific cases where GPT-4o successfully and fails to answer the given question correctly (refer to Table 11). In the first example, the model incorrectly associates the technique in answer choice A, i.e., T1574-Hijack Execution Flow, with the Execution tactic. While this technique involves the execution of malicious payloads based on the ATT&CK KB, it is associated to only the following tactics, Persistence, Privilege Escalation and Defense Evasion. This incorrect interpretation suggests that the model struggles in distinguishing the inherent ambiguity found in TTP descriptions, which affects their ability to analyze attack sequences.
Table 11. Error analysis of GPT-4o in AttackSeq-Technique. The LLM incorrectly associates technique T1574 with Execution tactic
Question & Answer Choices LLM Response
Question: Before establishing persistence through creating a Run key in the registry, which ATT&CK technique most likely facilitated the execution of the backdoor FinalBot.exe after the WMI discovery process identified the presence of antivirus software? Answer Choices: A) T1574-Hijack Execution Flow B) T1003.001-LSASS Memory C) T1059.008-Network Device CLI D) T1129-Shared Modules (Correct Answer) **A: T1574-Hijack Execution Flow** This technique involves hijacking the execution flow of a legitimate process to execute malicious code. DLL sideloading is a common method for this technique. The CTI report explicitly mentions that Sidewinder used DLL sideloading (‘cmpbk32.dll‘ to execute ‘Srvstr.exe‘), which aligns with this technique. This likely facilitated the execution of the backdoor ‘FinalBot.exe‘ after the WMI discovery process. **B: T1003.001-LSASS Memory**: This technique involves dumping credentials from the LSASS process memory. There is no mention in the CTI report of credential dumping or any activity related to LSASS. This choice is irrelevant. **C: T1059.008-Network Device CLI**: This technique involves using a command-line interface (CLI) to interact with network devices. The CTI report does not mention any interaction with network devices or the use of a CLI. This choice is irrelevant. **D: T1129-Shared Modules** This technique involves executing malicious code by loading shared modules. While DLL sideloading involves shared modules, the specific ATT&CK technique for DLL sideloading is T1574-Hijack Execution Flow. This choice is less precise than option A. The most plausible sequence of TTPs involves the use of **T1574-Hijack Execution Flow** (via DLL sideloading) to execute the backdoor ‘FinalBot.exe‘ after the WMI discovery process and before establishing persistence (Wrong Conclusion).
Appendix C Prompts Strategy
C.1. Question Generation Prompts
We utilize few-shot prompting for each of our benchmark tasks, the prompts for AttackSeq-Tactic, AttackSeq-Technique, AttackSeq-Procedure-Yes, AttackSeq-Procedure-No is shown in Tables 12, 13, 14, 15 respectively.
Table 12. Few-shot prompt for question generation for AttackSeq-Tactic. We insturct the LLM to generate a question based on the given CTI outline, where the answer to the question is the given ATT&CK TTP.
You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive two parts: 1. A **CTI Report** that describe a cyber attack ordered by MITRE ATT&CK tactics. Note that additional information labeled as “Others” provides context about the threat actor but is secondary. 2. A **MITRE ATT&CK Tactic** present in the CTI report. [End of Inputs]
Your task is to generate a question about the attack sequence based on the MITRE ATT&CK tactics found in the CTI report, where the answer to the question is the given MITRE ATT&CK tactic. The question should focus on inferring the given tactic by using the attack sequence based on the remaining tactics in the CTI report. [Steps]: Please follow these steps: 1. Analyze the CTI report: - Read the report carefully. - Identify and list the attack sequence in the order presented by the MITRE ATT&CK tactics. 2. Construct the Question: - Design a question that emphasizes the order of the attack sequence in the CTI report. - The question should exclude the TTPs under the given MITRE ATT&CK tactic that are described in the CTI report. Instead, include the TTPs in the tactic that precedes before and/or follows after the given MITRE ATT&CK tactic based on the order of tactics in the CTI report. - Ensure that the answer to the question is the given MITRE ATT&CK tactic. - The question should be concise, clear, and targeted towards experienced cybersecurity professionals. - Please refer to the example questions below for guidance. [Examples]: Example Questions: - Question: After gaining initial access through compromised VPN accounts, which ATT&CK tactic most likely occurred before Ke3chang achieved persistence by adding a Run key? Answer: Execution - Question: Which ATT&CK tactic most likely occurred before Axiom gained initial access to the victim’s network using SQL injection? Answer: Resource Development - Question: Which ATT&CK tactic most likely occurred after Ke3chang establishes connection with the C2 server through Internet Explorer (IE) by using the COM interface IWebBrowser2? Answer: Exfiltration - Question: After using stolen code signing certificates to sign DUSTTRAP malware and components, which ATT&CK tactic most likely occurred before APT41 used Windows Services with names such as Windows Defend for persistence of DUSTPAN? Answer: Execution [End of Examples]
3. Provide the Question-Answer Pair: - Please follow the output format: ”Question: ¡insert question here¿ Answer: ¡insert answer here¿”.
[End of Steps] Following the steps above, please generate a question based on the CTI report and ATT&CK tactic given below.
Table 13. Few-shot prompt for question generation for AttackSeq-Technique.
You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive three parts: 1. A **CTI Report** that describe a cyber attack ordered by MITRE ATT&CK tactics. Note that additional information labeled as “Others” provides context about the threat actor but is secondary. 2. A **MITRE ATT&CK Tactic** present in the CTI report. 3. A **MITRE ATT&CK Technique** present in the CTI report. [End of Inputs]
Your task is to generate a question about the attack sequence based on the MITRE ATT&CK tactics found in the CTI report, where the answer to the question is the given MITRE ATT&CK technique that belongs to the given ATT&CK tactic. The question should focus on inferring the given technique by using the attack sequence based on the remaining tactics in the CTI report. [Steps]: Please follow these steps: 1. Analyze the CTI report: - Read the report carefully. - Identify and list the attack sequence in the order presented by the MITRE ATT&CK tactics. 2. Construct the Question: - Design a question that emphasizes the order of the attack sequence in the CTI report. - The question should exclude the TTPs under the given MITRE ATT&CK tactic that are described in the CTI report. Instead, include the TTPs in the tactic that precedes before and/or follows after the given MITRE ATT&CK tactic based on the order of tactics in the CTI report. - Ensure that the answer to the question is the given MITRE ATT&CK technique. - The question should be concise, clear, and targeted towards experienced cybersecurity professionals. - Please refer to the example questions below for guidance. [Examples]: Example Questions: - Question: After gaining initial access through compromised VPN accounts, which ATT&CK technique most likely occurred before Ke3chang achieved persistence by adding a Run key? Answer: T1059-Command and Scripting Interpreter - Question: Which ATT&CK technique most likely occurred before Axiom gained initial access to the victim’s network using SQL injection? Answer: T1583.002-DNS Server - Question: Which ATT&CK technique most likely occurred after Ke3chang establishes connection with the C2 server through Internet Explorer (IE) by using the COM interface IWebBrowser2? Answer: T1020-Automated Exfiltration - Question: After using stolen code signing certificates to sign DUSTTRAP malware and components, which ATT&CK technique most likely occurred before APT41 used Windows Services with names such as Windows Defend for persistence of DUSTPAN? Answer: T1569.002-Service Execution [End of Examples]
3. Provide the Question-Answer Pair: - Please follow the output format: ”Question: ¡insert question here¿ Answer: ¡insert answer here¿”.
[End of Steps] Following the steps above, please generate a question based on the CTI report and ATT&CK tactic and technique given below.
Table 14. Few-shot prompt for question generation for AttackSeq-Procedure-Yes.
You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive four parts: 1. A **CTI Report** that describe a cyber attack ordered by MITRE ATT&CK tactics. Note that additional information labeled as “Others” provides context about the threat actor but is secondary. 2. A **MITRE ATT&CK Tactic** present in the CTI report. 3. A **MITRE ATT&CK Technique** present in the CTI report. 4. 4. A list of **Procedures** present in the CTI report, where each procedure is represented as a (Subject, Relation, Object) triplet. [End of Inputs]
Your task is to generate a question about the attack sequence based on the MITRE ATT&CK tactics found in the CTI report, the question should focus on inferring the given list of procedures based on the given MITRE ATT&CK tactic and technique. The answer to the question is ”Yes”, indicating that the given list of procedures is likely to occur in the attack sequence. [Steps]: Please follow these steps: 1. Analyze the CTI report: - Read the report carefully. - Identify and list the attack sequence in the order presented by the MITRE ATT&CK tactics. 2. Construct the Question: - Design a question that emphasizes the order of the attack sequence in the CTI report. - The question should exclude the TTPs under the given MITRE ATT&CK tactic that are described in the CTI report. Instead, include the TTPs in the tactic that precedes before and/or follows after the given MITRE ATT&CK tactic based on the order of tactics in the CTI report. - Ensure that the answer to the question is ”Yes”. - The question should be concise, clear, and targeted towards experienced cybersecurity professionals. - Please refer to the example questions below for guidance. [Examples]: Example Questions: - Question: After gaining initial access through compromised VPN accounts, is it likely that the Ke3chang malware will run commands on the command-line interface before achieving persistence by adding a Run key? Answer: Yes - Question: Is it likely that Axiom will acquire dynamic DNS services for use in the targeting of intended victims before gaining initial access to the victim’s network using SQL injection? Answer: Yes - Question: Is Ke3chang likely to perform frequent and scheduled data exfiltration from compromised networks after establishing connection with the C2 server through Internet Explorer (IE) by using the COM interface IWebBrowser2? Answer: Yes - Question: After using stolen code signing certificates to sign DUSTTRAP malware and components, is APT41 likely to use Windows services to execute DUSTPAN before using Windows Services with names such as Windows Defend for persistence of DUSTPAN? Answer: Yes [End of Examples]
3. Provide the Question-Answer Pair: - Please follow the output format: ”Question: ¡insert question here¿ Answer: ¡insert answer here¿”.
[End of Steps] Following the steps above, please generate a question based on the CTI report and ATT&CK tactic, technique and procedures given below.
Table 15. Few-shot prompt for question generation for AttackSeq-Procedure-No.
You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive two parts: 1. A **Reference Question-Answer Pair** that focuses on the logical sequence of TTPs in a CTI report, note that the answer to the this question is always ”Yes”. 2. A **Reference MITRE TTP** that is NOT supported in the CTI report. [End of Inputs]
Your task is to generate two questions based on the given reference question, such that attack sequence described in the question is modified and the correct answer to the two questions is ”No”. The definitions of the two questions are as follows: 1. Question 1 should negate the ”before” and/or ”after” clauses of the reference question, such that the attack sequence contradicts the original sequence of TTPs in the reference question. 2. Question 2 should replace the main procedures in the reference question with the provided reference MITRE TTP, such that the replaced main procedures are not found in the CTI report. [Steps]: Please follow these steps: 1. Analyze the Reference Question-Answer Pair: - Identify and outline the attack sequence in the order presented in the reference Question-Answer Pair. 2. Construct the Questions: - Design two questions that modify the attack sequence described in the reference question. - Ensure that the answer to the questions is ”No”. - The question should be concise, clear, and targeted towards experienced cybersecurity professionals. - Please refer to the examples below for guidance. [Examples]: Example Questions: - Example 1: Reference Question: After gaining initial access through compromised VPN accounts, will the Ke3chang malware most likely run commands on the command-line interface before achieving persistence by adding a Run key? Reference Answer: Yes Reference TTP: Tactic: Initial Access, Technique: T1651-Cloud Administration Command, Example Procedures: AADInternals can execute commands on Azure virtual machines using the VM agent. APT29 has used Azure Run Command and Azure Admin-on-Behalf-of (AOBO) to execute code on virtual machines. Pacu can run commands on EC2 instances using AWS Systems Manager Run Command. Question 1: After achieving persistence by adding a Run key, will the Ke3chang malware run commands on the command-line interface only after gaining initial access through compromised VPN accounts? Answer: No Question 2: After gaining initial access through compromised VPN accounts, will the Ke3chang malware most likely execute commands on Azure virtual machines using the VM agent before achieving persistence by adding a Run key? Answer: No - Example 2: Reference Question: Will Axiom acquire dynamic DNS services for use in the targeting of intended victims before gaining initial access to the victim’s network using SQL injection? Reference Answer: Yes Reference TTP: Tactic: Resource Development, Technique: T1585.001-Social Media Accounts, Example Procedures: APT32 has set up Facebook pages in tandem with fake websites. Cleaver has created fake LinkedIn profiles that included profile photos, details, and connections. EXOTIC LILY has established social media profiles to mimic employees of targeted companies. Question 1: Will Axiom acquire dynamic DNS services for use in the targeting of intended victims only after gaining initial access to the victim’s network using SQL injection? Reference Answer: No Question 2: Will Axiom set up Facebook pages in tandem with fake websites before gaining initial access to the victim’s network using SQL injection? Answer: No - Example 3: Reference Question: Will Ke3chang perform frequent and scheduled data exfiltration from compromised networks after establishing connection with the C2 server through Internet Explorer (IE) by using the COM interface IWebBrowser2? Reference Answer: Yes Reference TTP: Tactic: Exfiltration, Technique: T1030-Data Transfer Size Limits, Example Procedures: AppleSeed has divided files if the size is 0x1000000 bytes or more. APT28 has split archived exfiltration files into chunks smaller than 1MB. APT41 transfers post-exploitation files dividing the payload into fixed-size chunks to evade detection. Question 1: Will Ke3chang perform frequent and scheduled data exfiltration from compromised networks only before establishing connection with the C2 server through Internet Explorer (IE) by using the COM interface IWebBrowser2? Answer: No Question 2: Will Ke3chang divide files if the size is 0x1000000 bytes or more after establishing connection with the C2 server through Internet Explorer (IE) by using the COM interface IWebBrowser2? Answer: No [End of Examples]
3. Provide the Question-Answer Pairs: - Please follow the output format: ”Question 1: ¡insert question 1 here¿ Answer: ¡insert answer to question 1 here¿ Question 2: ¡insert question 2 here¿ Answer: ¡insert answer to question 2 here¿”.
[End of Steps] Following the steps above, please generate two questions based on the Reference Question-Answer Pair and Reference MITRE TTP given below. Please only provide the final output of the two questions.
C.2. Dataset Refinement Prompts
The prompt templates to filter based on the Answerability criteria is in Table 16, while the feedback and refinement prompts are in Table 17 and Table 18 respectively.
Table 16. Prompt template for verifying Answerability during Self-Refinement.
You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive three parts: 1. **CTI Outline**: A structured account of a cyber attack, ordered by MITRE ATT&CK tactics. Additional context under ”Others” provides background on the threat actor but is secondary. 2. **TTP Description**: A reference description of the correct answer corresponding to the question. 3. **Question with Answer Choices**: A question aimed at inferring a TTP from the attack sequence described in the CTI report, along with one correct answer and distractors among the answer choices. [End of Inputs] Your task is to evaluate the answerability of the given question using the provided information in the CTI Outline and TTP description. We define answerability based on three factors below: 1. The correct answer must be supported by the CTI outline. 2. The correct answer must clearly stand out as the best answer choice to the question based on the CTI outline. 3. Suppose the masked_tactic paragraph is removed from the CTI outline, the correct answer must be deducible from the answer choices by using the information provided in remaining tactics of the CTI outline and TTP description. You may also refer to your external cybersecurity knowledge to determine if the correct answer is deducible. [Steps]: Please follow these steps: 1. Analyze the CTI report: - Read the report carefully. - Identify and list the attack sequence in the order presented by the MITRE ATT&CK tactics. 2. Analyze the TTP Description: - Read the TTP description of the correct answer carefully. 3. Evaluate the Question with Answer Choices: - Read the question and the provided answer choices carefully. - Match the correct answer with the provided TTP description. - Determine step-by-step if the question is answerable based on the definition above. 3. Output evaluation result: - Output one of the following: - ”A”: Indicates that the question is answerable. - ”B”: Indicates that the question is not answerable. - ”C”: Indicates that you do not know/cannot determine if the question is answerable. - Please also include a short and concise explanation of your evaluation result. - Please follow the output format: ”Explanation: ¡insert explanation here¿ Evaluation Result: ¡insert letter here¿”. [End of Steps] Following the steps above, please evaluate the question using the CTI report and description below and only output the evaluation result.
Table 17. Prompt template for feedback on question quality based on evaluation criteria.
You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the Tactics, Techniques and Procedures (TTPs) in MITRE ATT&CK framework. [Inputs]: You will receive three parts: 1. **CTI Outline**: A structured account of a cyber attack, ordered by MITRE ATT&CK tactics. Additional context under ”Others” provides background on the threat actor but is secondary. 2. **TTP Description**: A reference description of the correct answer to the question. 3. **Question with Answer Choices**: A question aimed at inferring a TTP from the attack sequence described in the CTI outline, along with one correct answer and distractors among the answer choices. [End of Inputs]
[Evaluation Criteria]: Your task is to evaluate the QA pair and provide your feedback for each of the criteria defined below: 1. **Clarity**: Is the question precise, unambiguous, and free of vague phrasing? Does it avoid directly mentioning the correct answer, ensuring the respondent must infer the correct answer rather than having it stated in the question? 2. **Logical**: Does the question align with the logical sequence of MITRE ATT&CK tactics in the CTI outline? Does the question reference TTPs from the preceding or subsequent tactics in the CTI outline such that it logically leads to the correct answer? 3. **Relevance**: Does the question directly relate to the CTI outline? 4. **Consistency**: Does the question align with the provided TTP Description? 5. **Answer Consistency**: Can the question be fully answered using the correct answer, without any contradictions or inconsistencies? [End of Evaluation Criteria] [Steps]: Please follow these steps: 1. Analyze the CTI outline: - Read the CTI outline carefully. - Identify and outline the attack sequence in the order presented by the MITRE ATT&CK tactics. 2. Analyze the TTP Description: - Read the TTP description of the correct answer carefully. 3. Evaluate the Question with Answer Choices: - Read the question and the provided answer choices carefully. - Assess each criterion step by step, rating it on a scale of 1 to 5 (1 = poor, 5 = excellent). - Provide a short and concise feedback for each rating. 4. Output Feedback Scores: - Please follow the output format: Feedback Scores: - Clarity: ¡Your feedback¿ (¡Score¿/5) - Logical: ¡Your feedback¿ (¡Score¿/5) - Relevance: ¡Your feedback¿ (¡Score¿/5) - Consistency: ¡Your feedback¿ (¡Score¿/5) - Answer Consistency: ¡Your feedback¿ (¡Score¿/5) Total Score: ¡Total Score¿/25 [End of Steps] Following the steps above, please evaluate the Question with Answer Choices below using the provided CTI report and TTP Description. Please only output the Feedback Scores.
Table 18. Prompt template for question refinement.
You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive three parts: 1. **CTI Outline**: A structured account of a cyber attack, ordered by MITRE ATT&CK tactics. Additional context under ”Others” provides background on the threat actor but is secondary. 2. **Question with Answer Choices**: A question aimed at inferring a TTP from the attack sequence described in the CTI report, along with one correct answer and distractors among the answer choices. 3. **Feedback Results**: A list of feedback scores and explanations for each desired criterion of the question defined below. [End of Inputs] [Evaluation Criteria]: Please refer to the definition of each feedback criterion: 1. **Clarity**: Is the question precise, unambiguous, and free of vague phrasing? Does it avoid directly mentioning the correct answer, ensuring the respondent must infer the correct answer rather than having it stated in the question? 2. **Logical**: Does the question align with the logical sequence of MITRE ATT&CK tactics in the CTI outline? Does the question reference TTPs from the preceding or subsequent tactics in the CTI outline such that it logically leads to the correct answer? 3. **Relevance**: Does the question directly relate to the CTI outline? 4. **Consistency**: Does the question align with the provided TTP Description? 5. **Answer Consistency**: Can the question be fully answered using the correct answer, without any contradictions or inconsistencies? Your task is to iteratively refine the quality of the given question based on the feedback provided in the Feedback Results. [End of Evaluation Criteria] [Steps]: Please follow these steps: 1. Analyze the CTI report: - Read the report carefully. 2. Analyze the Question with Answer Choices: - Read the question and the provided answer choices carefully. 3. Analyze the Feedback Results: - Based on the feedback given in each criterion, refine the question to improve the each aspect. - Please ensure that the correct answer to the refined question is the same as the original question. - Please also ensure that the question avoids hinting at the correct answer. 4. Output the Refined Question: - Please follow the output format: ”Refined Question: ¡Your refined question here¿”. [End of Steps] Following the steps above, please refine the question based on the Feedback Results and CTI Outline provided below. Please only output the refined question.
C.3. Automatic Evaluation Prompt
We utilize the definitions in the evaluation criteria in Table 7 to create prompts. We show an example prompt template for evaluating the Logical aspect of the question shown in Table 19.
Table 19. Prompt template for evaluating the Logical aspect.
You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive three parts: 1. **CTI Outline**: A structured account of a cyber attack, ordered by MITRE ATT&CK tactics. Additional context under ”Others” provides background on the threat actor but is secondary. 2. **Question with Answer Choices**: A question aimed at inferring a TTP from the attack sequence described in the CTI report, along with one correct answer and distractors among the answer choices. 3. **Description to Correct Answer**: The description of the correct answer from the MITRE ATT&CK framework. [End of Inputs] Your task is to rate the question on one metric below. [Definition of Logical Aspect] Evaluation Criteria: Logical (1-5): Does the question align with the logical sequence of MITRE ATT&CK tactics in the CTI outline? Does the question reference TTPs from the preceding and/or subsequent tactics in the CTI outline such that it logically leads to the correct answer? The scale is defined as follows: 1 - Not Logical: The question does not align with the logical sequence of MITRE ATT&CK tactics in the CTI outline. It ignores or contradicts the natural order of tactics and TTPs. 2 - Weak Logical Alignment: The question shows minimal alignment with the MITRE ATT&CK sequence. It may reference unrelated tactics or disrupt the logical flow. 3 - Moderately Logical: The question has some logical alignment, but it may not reference preceding or subsequent tactics clearly. The sequence could be improved. 4 - Strong Logical Alignment: The question follows the expected sequence of MITRE ATT&CK tactics and references preceding or subsequent TTPs in a logical manner. 5 - Perfect Logical Alignment: The question perfectly aligns with the MITRE ATT&CK framework, referencing relevant TTPs in a way that naturally leads to the correct answer. [Steps:] Evaluation Steps: 1. Analyze the CTI report and Description to Correct Answer: - Read the report and the provided description carefully. - Identify and list the attack sequence in the order presented by the MITRE ATT&CK tactics. 2. Evaluate the Question: - Read the question and the provided answer choices carefully. - Using the CTI outline and provided description to the correct answer, Rate the question on a scale of 1-5 according to the evaluation criteria above. 3. Output evaluation score: - Please only output the numerical evaluation score based on the defined criteria. [End of Steps] Following the steps above, please evaluate the question and only output the numerical evaluation score.
C.4. Answering Prompts
The prompt templates for the three benchmarking tasks, Regular, Zero-Shot and RAG-empowered are shown in Tables 20, 21, 22 respectively.
Table 20. Prompt template for Regular setting.
You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive two parts: 1. A **CTI Report** that describe a cyber attack ordered by MITRE ATT&CK tactics. Note that additional information labeled as “Others” provides context about the threat actor but is secondary. 2. A **Question** about a sequence of TTPs with several answer choices. [End of Inputs]
Your task is to determine which answer choice forms the most plausible sequence of TTPs based on the attack sequence described in the CTI report. Note that the CTI report contains key details required for your analysis, but it may not directly state the answer. Your evaluation of the answer choices is essential to arrive at the correct answer. [Steps:] Please follow these steps: 1. Analyze the CTI report: - Read the report carefully. - Identify and list the attack sequence in the order presented by the MITRE ATT&CK tactics. 2. Analyze the Question: - Read the question and its answer choices. - Identify the sequence of TTPs mentioned in the question. 3. Compare and Evaluate: - Match the extracted attack sequence from the CTI report with the details in the question. - Evaluate each answer choice to determine which one aligns best with the attack sequence and any critical contextual information. 4. Provide a Step-by-Step Reasoning and Final Answer: - Outline your reasoning step-by-step. - Conclude with the final answer in the following format: ”Final Answer: ¡insert answer choice here¿”. [End of Steps] Following the steps above, please answer the question below using the provided CTI report.
Table 21. Prompt template for Zero-Shot setting.
You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive a question about a sequence of TTPs with several answer choices. Your task is to determine which answer choice forms the most plausible sequence of TTPs based on the attack sequence described in the question. [End of Inputs]
[Steps:] Please follow these steps: 1. Analyze the Question: - Read the question and its answer choices. - Identify the sequence of TTPs mentioned in the question. 2. Compare and Evaluate: - Evaluate each answer choice to determine which one aligns best with the attack sequence in the question. 3. Provide a Step-by-Step Reasoning and Final Answer: - Outline your reasoning step-by-step. - Conclude with the final answer in the following format: ”Final Answer: ¡insert answer choice here¿”. [End of Steps] Following the steps above, please answer the question below.
Table 22. Prompt template for RAG-empowered setting.
You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework. [Inputs]: You will receive two parts: 1. A **Question** about a sequence of TTPs with several answer choices. 2. A list of **Related TTPs** that are relevant to the question. [End of Inputs]
Your task is to determine which answer choice forms the most plausible sequence of TTPs based on the attack sequence described in the question.
[Steps:] Please follow these steps: 1. Analyze the Question: - Carefully read the question and its answer choices. 2. Analyze the Related TTPs: - Analyze the list of Related TTPs to understand the context of the question. 3. Compare and Evaluate: - Based on the related TTPs, evaluate each answer choice to determine which one aligns best with the attack sequence in the question. 4. Provide a Step-by-Step Reasoning and Final Answer: - Outline your reasoning step-by-step. - Conclude with the final answer in the following format: ”Final Answer: ¡insert answer choice here¿”. [End of Steps] Following the steps above, please answer the question below.

### v2

@@
-AttackSeqBench: Benchmarking Large Language Models’ Understanding of Sequential Patterns in Cyber Attacks
-Abstract.
-The observations documented in Cyber Threat Intelligence (CTI) reports play a critical role in describing adversarial behaviors, providing valuable insights for security practitioners to respond to evolving threats.
-Recent advancements of Large Language Models (LLMs) have demonstrated significant potential in various cybersecurity applications, including CTI report understanding and attack knowledge graph construction.
-While previous works have proposed benchmarks that focus on the CTI extraction ability of LLMs, the sequential characteristic of adversarial behaviors within CTI reports remains largely unexplored, which holds considerable significance in developing a comprehensive understanding of how adversaries operate.
-To address this gap, we introduce AttackSeqBench, a benchmark tailored to systematically evaluate LLMs’ capability to understand and reason attack sequences in CTI reports.
-Our benchmark encompasses three distinct Question Answering (QA) tasks, each task focuses on the varying granularity in adversarial behavior.
-To alleviate the laborious effort of QA construction, we carefully design an automated dataset construction pipeline to create scalable and well-formulated QA datasets based on real-world CTI reports.
-To ensure the quality of our dataset, we adopt a hybrid approach of combining human evaluation and systematic evaluation metrics.
-We conduct extensive experiments and analysis with both fast-thinking and slow-thinking LLMs, while highlighting their strengths and limitations in analyzing the sequential patterns in cyber attacks.
-The overarching goal of this work is to provide a benchmark that advances LLM-driven CTI report understanding and fosters its application in real-world cybersecurity operations.
-Our dataset and code are available at https://github.com/Javiery3889/AttackSeqBench.
-Cyber Security, Cyber Threat Intelligence, Large Language Models
-copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Security and privacyccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Machine learning approaches
-1. Introduction
-Figure 1.
-Illustration of the difference between existing CRU-related benchmarks and our AttackSeqBench based on a real-world CTI report (Koehl and Hannon, 2020).
-We highlight the sequential patterns within CTI reports and investigate the capability of LLMs in comprehending and reasoning such sequential patterns.
-The increasing complexity and diversity of cyber attacks have imposed a challenge to defend against Advanced Persistent Threats (APTs) (Huang et al., 2018; Duo et al., 2022).
-Cyber Threat Intelligence (CTI) reports, which detail the adversarial behaviors observed in cyber attacks, are crucial towards addressing emerging cyber threats.
-However, their unstructured nature demands extensive manual analysis (Wagner et al., 2019; Sun et al., 2023).
-Meanwhile, recent advancements in Large Language Models (LLMs) have demonstrated promising applications in cybersecurity, such as attack knowledge graph construction (Zhang et al., 2025b; Hu et al., 2024) and threat intelligence extraction (Rahman et al., 2023; Xu et al., 2024b).
-These capabilities shed new light on LLM-driven CTI Report Understanding (CRU) to support cybersecurity operations, leading to the development of various CRU-related benchmarks to assess LLMs’ capability in the field.
-Existing CRU-related benchmarks primarily focus on evaluating LLMs on three aspects: (1) Classifying adversarial behaviors (Alam et al., 2023) (2) Extracting threat intelligence from unstructured text (Bhusal et al., 2024), or (3) Understanding meaningful correlations between the cyber threats and malware groups (Alam et al., 2024).
-While these benchmarks focus on the information extraction capabilities of LLMs, their ability to understand and reason about the sequential patterns of adversarial behavior remains largely unexplored.
-We refer to these sequences as attack sequences (Choi et al., 2021; Al-Sada et al., 2025), which represent the temporal order of malicious actions adversaries perform to achieve their goals.
-The actions which constitute an attack sequence are commonly referred to as attack patterns, characterizing adversarial behavior on a tactical level (Mavroeidis and Bromander, 2021; Siracusano et al., 2023).
-We illustrate the key differences between existing CRU-related tasks and analyzing attack sequences in Figure 1.
-Although knowledge bases (KBs) such as the MITRE ATT&CK® framework (Strom et al., 2018) document real-world adversary behaviors through defined attack patterns, analyzing the patterns individually is insufficient to fully capture the progression of cyber attacks as documented in CTI reports.
-The sophisticated and stealthy nature of APTs requires a comprehensive understanding of how adversaries transition between the different attack phases, which are orchestrated as an attack sequence.
-This raises the need to consider the sequential characteristics of a cyber attack during CRU.
-To address this research gap, we delve into the suitability of LLMs in analyzing attack sequences. We highlight three key perspectives when formulating our main research task:
-Perspective 1: The process of mining threat intelligence from CTI reports not only requires general text comprehension; it also necessitates expert-level cybersecurity knowledge.
-This requirement becomes apparent in attack sequence analysis, as it requires a deep understanding of adversarial behavior to formulate the associate attack sequence.
-Although LLMs demonstrate significant potential in CRU, prior research has identified limitations in their factual reliability, particularly in technical and knowledge-intensive tasks (Wang et al., 2023a; Xu et al., 2024a).
-In real-world cybersecurity operations, these limitations can hinder precise and timely analysis of CTI reports, ultimately impacting defensive strategies.
-Perspective 2: LLMs have demonstrated remarkable capabilities across a broad range of scenarios, from traditional Natural Language Processing (NLP) tasks to more even more challenging real-world problems (Zhao et al., 2023).
-The recent breakthrough of slow-thinking reasoning LLMs has also shed new light on complex reasoning tasks related to coding and mathematical reasoning (DeepSeek-AI, 2025; OpenAI, 2025).
-Yet, the extent to which LLMs can be applied to attack sequence analysis has remained under-explored.
-Perspective 3: To address the constantly evolving cyber threat landscape and the advancements of LLMs, the benchmark must be extensible to incorporate newly observed attack sequences from CTI reports.
-Additionally, the unstructured nature of CTI reports necessitates the benchmark to be generalizable to ensure a comprehensive and realistic evaluation of LLMs’ capabilities.
-Most importantly, the benchmark must be reproducible to facilitate fair and consistent evaluations across the different LLMs.
-Building on these perspectives, we introduce AttackSeqBench, a novel benchmark designed for extensive investigation.
-Specifically, considering Perspective 1, we define three distinct Question Answering (QA) benchmarking tasks that follow the three-level structure of adversary behaviors —Tactics, Techniques, and Procedures (TTPs) —as outlined in the ATT&CK KB.
-This hierarchical representation allows security practitioners to represent observed attack patterns in varying levels of granularity based on pre-defined adversarial behaviors.
-Each tactic represents a high-level objective within a cyber attack’s lifecycle, while techniques and procedures elaborate on the specific actions performed by an adversary to achieve their objectives.
-Following Perspective 2, we design three benchmark settings to extensively evaluate the effectiveness and robustness of AttackSeqBench.
-Moreover, we leverage a diverse set of fast-thinking and slow-thinking LLMs in our experiments (Min et al., 2024), providing comprehensive insights into the capabilities and limitations of LLMs in attack sequence analysis.
-Catering to Perspective 3, we construct attack sequences based on real-world CTI reports.
-Additionally, we carefully design an automated dataset generation pipeline to generate questions in our QA dataset based on the attack sequence constructed.
-We summarize our contributions as follow:
-  • We propose AttackSeqBench, a novel benchmark designed to systematically evaluate LLMs’ capabilities in analyzing attack sequences constructed from real-world CTI reports.
-  • We introduce three QA tasks in our benchmark, each aligned with the different levels of granularity in adversarial behavior representation.
-  • We conduct extensive experiments and various benchmarking settings using a diverse set of fast-thinking and slow-thinking LLMs to provide insights into their capabilities and limitations.
-2. Related Work
-CTI Report Understanding and LLMs.
-CTI refers to the collection and analysis of evidence-based information related to cyber attacks and threat actors, enabling security practitioners to gain actionable insights into the adversarial operations respond to these adversaries effectively (Wagner et al., 2019; Sun et al., 2023).
-A crucial source of CTI comes from CTI reports published by security vendors, these reports are typically presented as unstructured textual documents and document observations based on emerging and ongoing cyber threats (Rahman et al., 2020).
-CTI Report Understanding (CRU) broadly encompasses the extraction, interpretation, and reasoning of key insights within CTI reports.
-TTP extraction is a critical aspect in CRU, enabling security practitioners to characterize adversarial behaviors into known attack patterns defined in the ATT&CK KB (Husari et al., 2017).
-Some works (Legoy et al., 2020; Marchiori et al., 2023) leverage multi-label text classification to extract TTPs, while more recent research utilize Language Models (LMs) to perform TTP extraction (Rani et al., 2024).
-Another line of work (Shin et al., 2023) explores embedding TTPs using Global Vectors for Word Representation to capture co-occurrences.
-Despite considerable progress, these works tend to be costly and error-prone as adversarial behavior continues evolving.
-With the breakthrough of LLMs, researchers have progressively explored their applicability within the cybersecurity domain (Zhang et al., 2024a).
-Within CRU, existing research (Fayyazi et al., 2024) has showcased significant capabilities of LLMs in interpreting TTPs from the ATT&CK KB, surpassing the performance of some of the fine-tuned LMs with cybersecurity data.
-Another line of works (Huang and Xiao, 2024; Cheng et al., 2024) utilizes LLMs to extract the relevant entities to construct knowledge graphs (KGs) based on the CTI reports.
-AttacKG+ (Zhang et al., 2025b) constructs adversarial behavior KGs by leveraging LLMs to map extracted entities to ATT&CK TTPs.
-However, the capabilities of LLMs to analyze attack sequences in CTI reports remain unexplored.
-Related Benchmarks.
-Within the cybersecurity domain, existing work (Liu, 2023; Tihanyi et al., 2024; Garza et al., 2023) have proposed several benchmarks to evaluate LLMs’ cybersecurity knowledge.
-However, these works primarily rely on authoritative sources (e.g., textbooks, open standards, and example TTPs from ATT&CK KB) and do not incorporate CTI reports in their dataset.
-CTIBench (Alam et al., 2024) designs four CTI-related tasks, including threat actor attribution and vulnerability assessment, yet only the threat actor attribution task leverages CTI reports.
-CTISum (Peng et al., 2024) evaluates the ability of LLMs to summarize key facts and attack processes from CTI reports.
-Extending this work, SevenLLM (Ji et al., 2024) designs an additional task that focuses on extracting entities and relations.
-To add on, prior research (Siracusano et al., 2023) assesses LLMs’ capabilities in extracting attack patterns, along with entities and relations.
-In the general domain, existing works (Wang et al., 2023b; Valmeekam et al., 2023) have proposed benchmarks to evaluate the reasoning capabilities of LLMs in sequence-related tasks.
-Furthermore, previous research has highlighted LLMs capabilities in temporal event analysis (Zhang et al., 2024b; Chang et al., 2024).
-However, these works may not accurately reflect the extent to which LLMs can reason about attack sequences.
-In contrast, AttackSeqBench delves into the suitability of LLMs in attack sequence analysis through three tasks that requires in-depth knowledge of adversarial behavior, while emphasizing the practical aspects of CRU by utilizing real-world CTI reports to construct attack sequences in our dataset.
-3. Task Definition and Dataset
-3.1. Task Definition
-An attack sequence $S$ models the progression of adversarial behavior in a cyber attack, characterized by the logical order of the ATT&CK tactics observed in a CTI report alongside the associated techniques and procedures.
-Formally, $S$ is defined as a 4-tuple:
-$S=(T,E,P,O),$
-where:
-  1. (1)
-    Tactic Sequence $T=(t_{1},\ldots,t_{n})$: An ordered list of ATT&CK tactics, which outlines the phases of a cyber attack identified in a CTI report.
-Note that tactic $t_{k}$ represents the $k$-th tactic.
-  2. (2)
-    Technique Mapping $E(t_{k})=\{e_{1,k},\ldots,e_{i_{k},k}\}$: A set of ATT&CK techniques (and/or sub-techniques) associated with tactic $t_{k}$.
-The overall set of techniques across all tactics in the CTI report is represented as $E=\bigcup_{k=1}^{n}E(t_{k})$.
-  3. (3)
-    Procedure Mapping $P(e_{j,k})=\{p_{1,j,k},\ldots,p_{m_{j,k},j,k}\}$: A set of procedures describing how technique $e_{j,k}$ is performed, the overall set of procedures across all techniques in the CTI report is represented as $P=\bigcup_{e\in E}P(e)$.
-Each procedure $p_{l,j,k}$ is expressed as a (subject, action, object) triplet, e.g., (The Lazarus group, use, spear-phishing emails).
-  4. (4)
-    CTI Outline $O=(o_{1},\ldots,o_{n})$: A tactic-centric outline of the TTPs described in the CTI report, where $o_{k}$ is a textual summary of the techniques and procedures associated with $t_{k}$.
-AttackSeq-Tactic.
-This task tests the model’s ability to deduce a tactic $t_{k}\in S$, given a question $Q$ and four shuffled candidate tactics $C=\{c_{r}:r\in[1,4]\}$.
-The model must select the correct tactic $c_{l}\in C$ that corresponds to $t_{k}$.
-AttackSeq-Technique.
-This task evaluates the model’s capability to infer a technique $e_{j,k}\in E(t_{k})$, given a question $Q$ and four shuffled candidate techniques $C=\{c_{r}:r\in[1,4]\}$.
-The model must select the correct technique $c_{l}\in C$ that corresponds to $e_{j,k}$.
-AttackSeq-Procedure.
-This task challenges the model to determine the plausibility of a group of procedures $P(e_{j,k})\subseteq P$.
-The input is a question $Q$ and two candidate choices $C=\{c_{yes},c_{no}\}$.
-The model must select the correct choice $c_{l}\in C$ that indicates if $P(e_{j,k})$ is plausible.
-We further divide the task based on the correct answer to the question, i.e.
-AttackSeq-Procedure-Yes and AttackSeq-Procedure-No.
-Within AttackSeqBench, we create three distinct benchmark settings (refer to Section 4.1) that have varying inputs for each benchmarking task differ to the LLMs.
-These benchmark settings are as follows: Zero-shot, where the input is a question $Q$ and candidate choices $C$; RAG-empowered, where the input is a question $Q$ and candidate choices $C$, along with a list of relevant ATT&CK TTPs to $Q$; Regular, where the input is a question $Q$ and candidate choices $C$, along with a masked CTI outline where the textual summary of $t_{k}$ is removed from $O$, i.e., $O_{masked}=O\setminus o_{k}$.
-Figure 2.
-The pipeline of our dataset construction process.
-Given the CTI reports, we first leverage a LLM to extract the attack sequence and re-write the CTI outlines.
-Then we generate massive Q&A pairs with a question generator.
-Finally, we design the tailored evaluation criteria to enable LLM to select the appropriate Q&A pairs and construct the datasets for AttackSeqBench.
-3.2. Dataset Construction
-As illustrated in Figure 2, the automated dataset construction pipeline is divided into three phases: (1) Attack Sequence Construction, (2) Question Generation, and (3) Self-Refinement.
-3.2.1. Attack Sequence Construction
-AttacKG+ (Zhang et al., 2025b) has summarized 500 real-world CTI reports from multiple intelligence sources into tactic-centric outlines, and extracted the associated TTPs by leveraging a LLM-based framework.
-To construct attack sequences that resemble real-world adversarial behavior, we exclude CTI outlines which contains less than two ATT&CK tactics as they are unlikely to detail attack patterns observed in cyber attacks (Zhang et al., 2025b).
-Afterwards, we combine the CTI outlines and the associated TTPs from 408 filtered CTI reports to form attack sequences.
-3.2.2. Question Generation
-We leverage the question generation capabilities of LLMs (Alam et al., 2024; Zhang et al., 2024b; Mucciaccia et al., 2025) to formulate the questions in our QA dataset based on the attack sequences constructed.
-More specifically, we propose an answer-aware question generation approach, by utilizing GPT-4o (OpenAI, 2024b) to generate a question based on each TTP present with respect to the scope of each benchmarking task for a given attack sequence $S$.
-To generate more relevant questions, we leverage LLMs’ in-context learning (ICL) ability (Dong et al., 2024) by providing the associated outline $O$ as contextual knowledge, and include few-shot examples based on the example procedures in the ATT&CK KB (refer to Section C.1).
-For the MCQ tasks, we adopt a rule-based approach to randomly select three distractors from the attack sequence and ATT&CK KB.
-Due to the limited tactics present in the ATT&CK KB, we randomly sample an adjacent tactic present in the attack sequence i.e., $t_{k+1}$ or $t_{k-1}$, along with two randomly selected tactics from the ATT&CK KB.
-For AttackSeq-Technique, we select the distractors with reference to the STARC annotation framework (Berzak et al., 2020).
-The three distractors are as follow: (1) Similar to AttackSeq-Tactic, the first distractor is associated with the same tactic but not supported by the CTI report, i.e., $e_{i,k}\in E(t_{k})$, indicating a plausible misunderstanding of the correct technique $e_{j,k}$; (2) The second distractor refers to a technique that is supported by the CTI report but belongs to another tactic $t_{j}\neq t_{k}$; (3) The third distractor refers to a technique that is not supported by the CTI report and belongs to a randomly chosen tactic.
-For the boolean questions, AttackSeq-Procedure, we first generate the questions with correct answer ”Yes”, i.e., for each group of procedures $P(e_{j,k})\subseteq P$.
-To create the negative questions in AttackSeq-Procedure-No, we randomly sample 70% of the AttackSeq-Procedure-Yes questions from our final QA dataset and prompt the LLM to generate two types of negative questions based on each sampled question (refer to Table 15).
-Inspired by existing work (Rajpurkar et al., 2018), the two types of negative questions are: (1) We negate the temporal prepositions, i.e., changing ”before” to ”only after” and/or ”after” to ”only before”, such that the sequence of adversarial behavior is incorrect; (2) We replace $P(e_{j,k})$ in the question’s body with a group of procedures associated with a randomly sampled technique that is not supported by CTI report.
-3.2.3. Self-Refinement
-Inspired by the Self-Refine framework (Madaan et al., 2023), we adopt an iterative three-step approach using GPT-4o to improve the quality of the generated questions.
-We first assess the quality of the question based on the Answerability aspect (Fu et al., 2024), filtering out the questions that do not satisfy this criteria.
-More specifically, we check if there is direct evidence in the CTI outline that supports the correct answer, while clearly standing out as the best answer choice.
-Within this aspect, we also check if the correct answer can be inferred even if the associated summary to the correct answer’s tactic is removed from the outline (Huang et al., 2019), i.e., using $O_{masked}$ to answer the question.
-Based on the filtered answerable questions, we prompt the LLM to consider the quality of the question and its answer choices based on the following five aspects:
-  • Clarity. We check if the question precise and unambiguous. More importantly, we also ensure that question avoid directly mentioning the correct answer such that the inference is required.
-  • Logical. We check if the sequence described in the question follow the order of tactics present in the attack sequence.
-  • Relevance. We check if the TTPs described in the question directly relate to the attack sequence.
-  • Consistency. We check if the question is consistent with the associated TTP that is used for question generation.
-  • Answer Consistency. We check if the question can be fully answered by the correct answer, without any contradictions and inconsistencies.
-Based on these aspects, we prompt the LLM to give a numerical score out of 5 and select the questions that have a full score of 25.
-This filter-feedback-refine process is repeated once more to ensure better quality of questions.
-Refer to Appendix C.2 for the detailed prompts.
-Table 1. Number of questions and proportional decrease in each benchmarking task before and after Self-Refinement.
-Task Num. (Before) Num. (After) % Decrease
-AttackSeq-Tactic 2,158 1,697 -21.36
-AttackSeq-Technique 2,937 1,917 -34.73
-AttackSeq-Procedure 4,642 2,635 -43.24
-3.2.4. Dataset Statistics
-The number of generated questions for each benchmark task, along with the corresponding percentage decrease, is shown in Table 1.
-More details of the dataset, such as the distribution of ATT&CK tactics and techniques associated with the questions, is in Appendix A.1.
-3.3. Dataset Evaluation
-3.3.1. Human Evaluation
-We ask two postgraduate students majoring in computer science, who are knowledgeable about cybersecurity and the MITRE ATT&CK framework, to assess the quality of questions in AttackSeqBench.
-We randomly sample 35 questions from each of the three benchmarking tasks, including the AttackSeq-Procedure-No subtask, resulting in a total of 140 questions for manual evaluation.
-Each annotator evaluates the six aspects described in Section 3.2.3 using a 5-point Likert scale, where higher scores indicate better alignment (detailed definitions of each aspect are in Appendix A.2).
-Based on Table 8, the average accuracy score of annotators over the three tasks is 0.625, suggesting that questions in AttackSeqBench are notably challenging for domain experts.
-Furthermore, the human evaluation results indicate that the generated questions closely satisfy with each aspect of the evaluation criteria, with an average score of 4.620/5 across the six aspects.
-In particular, the Logical and Relevance aspects suggests that LLMs are able to construct questions that closely align with attack sequences described in CTI reports.
-3.3.2. Automatic Evaluation
-To alleviate the laborious task of manual human annotation, recent works (Zheng et al., 2023; Yao et al., 2024) have shown considerable effectiveness of LLM-as-a-Judge frameworks in aligning with human preferences within domain-specific fields, including cybersecurity (Xu et al., 2024b).
-We incorporate G-Eval (Liu et al., 2023), a Chain-of-Thought (CoT) (Wei et al., 2022) and form-filling paradigm, to systematically assess the quality of generated questions.
-Similar to Section 3.3.1, we carefully design prompts for each of the six aspects using a 5-point Likert scale (refer to Appendix C.3 for the detailed prompt).
-More specifically, we instruct GPT-4o to refer to the evaluation criterion’s definition, along with description of the correct answer from the ATT&CK KB, and rate the question out of 5 for each aspect.
-Based on Table 9, we observe that Logical and Relevance are the highest rated aspects, reinforcing the LLM’s ability to construct questions that follow the logical order of attack sequences.
-While the automatic evaluation scores are lower than the human evaluation scores, this implies that it is more challenging for LLMs to answer the questions correctly as compared to domain experts.
-4. Experiment
-Figure 3. The illustration of three benchmark settings. (a): Regular setting, (b): Zero-shot setting, and (c): RAG-empowered setting.
-4.1. Benchmark Settings
-As illustrated in Figure 3, we elaborate on the three proposed benchmark settings, Regular, Zero-Shot and RAG-Empowered. The detailed prompts for each setting is in Appendix C.4.
-Regular Setting.
-Inspired by existing works (Ma et al., 2023; Jin et al., 2024), we propose this benchmark setting to assess the impact of contextual information in analyzing attack sequences.
-We adopt an ICL-based approach by carefully designing a well-constructed prompt using CoT.
-Within the prompt, we first instruct the LLM to analyze the attack sequence described in the masked CTI outline associated with the question, and subsequently deduce the correct answer based on the given attack sequence.
+AttackSeqBench: Benchmarking Large Language Models in Analyzing Attack Sequences within Cyber Threat Intelligence
+Abstract
+Cyber Threat Intelligence (CTI) reports document observations of cyber threats, synthesizing evidence about adversaries’ actions and intent into actionable knowledge that informs detection, response, and defense planning.
+However, the unstructured and verbose nature of CTI reports poses significant challenges for security practitioners to manually extract and analyze such sequences.
+Although large language models (LLMs) exhibit promise in cybersecurity tasks such as entity extraction and knowledge graph construction, their understanding and reasoning capabilities towards behavioral sequences remains underexplored.
+To address this, we introduce AttackSeqBench, a benchmark designed to systematically evaluate LLMs’ reasoning abilities across the tactical, technical, and procedural dimensions of adversarial behaviors, while satisfying Extensibility, Reasoning Scalability, and Domain-dpecific Epistemic Expandability.
+We further benchmark 7 LLMs, 5 LRMs and 4 post-training strategies across the proposed 3 benchmark settings and 3 benchmark tasks within our AttackSeqBench to identify their advantages and limitations in such specific domain.
+Our findings contribute to a deeper understanding of LLM-driven CTI report understanding and foster its application in cybersecurity operations.
+1 Introduction
+Amid rapid digital transformation, the increasing sophistication and diversity of cyber attacks have become a pervasive concern for cybersecurity globally (Duo et al., 2022).
+Cyber Threat Intelligence (CTI) reports, which document observations of these threats, have emerged as a crucial resource in proactive defenses (Wagner et al., 2019).
+However, they are often lengthy and unstructured, resulting in a labor-intensive task for practitioners to manually analyze and extract insights (Sun et al., 2023).
+Recently, Large Language Models (LLMs) have demonstrated promising potential in several cybersecurity applications (Zhang et al., 2024a).
+This sheds new light towards incorporating LLMs into CTI Report Understanding (CRU) task, where we define CRU as a broad concept encompassing tasks that derive and reason threat intelligence from CTI reports.
+However, existing benchmarks primarily assess LLMs on threat intelligence extraction and attack attribution, while their potential for understanding adversarial behaviors dependencies in CTI reports remains largely unexplored (cf.
+Appendix A.7).
+Such ability is crucial in anticipating future malicious attack actions, particularly in multi-stage cyber attacks launched by Advanced Persistent Threats (APTs) (Li et al., 2022).
+As illustrated in Figure 1, we define the sequence of adversary behaviors as attack sequence (Al-Sada et al., 2025) to represent the execution flow of malicious actions across different stages of a cyber attack under the MITRE ATT&CK® framework (Strom et al., 2018).
+Building on this definition and the following key perspectives, we further delve into the suitability of LLMs in analyzing attack sequences.
+1) Extensibility: To address the ever-evolving threat landscape and the advancements of LLMs, the proposed benchmark must be extensible to incorporate attack sequences from newly observed CTI reports.
+2) Reasoning Scalability: Recently, Large Reasoning Models (LRMs) have demonstrated substantial advantages over conventional LLMs in multi-step reasoning tasks, such as coding and mathematical reasoning.
+However, existing CRU works have primarily focused on addressing CTI–related tasks via LLM, leaving the necessity of reasoning for attack sequence analysis largely unexplored.
+3) Domain-Specific Epistemic Expandability: LLMs exhibited limitations in factual reliability on knowledge-intensive tasks (Xu et al., 2024b), analogously, LLM-driven CRU, which requires specialized cybersecurity knowledge, is also subject to such limitations.
+This requirement becomes particularly pronounced in attack sequence analysis, which necessitates a comprehensive understanding of adversarial behaviors to effectively reason multi-stage cyber attacks.
+Figure 1: Illustration an example cyber attack sequence and our AttackSeqBench.
+Building upon these perspectives, we introduce AttackSeqBench, a novel benchmark designed for comprehensive evaluation of LLMs in attack sequence analysis.
+Catering to Extensibility, we first construct attack sequences based on extensive real-world CTI reports, ensuring that the benchmark accurately reflects the complexity and diversity of Tactics, Techniques, and Procedures (TTPs) in cyber attacks performed by APTs.
+Moreover, we design three Question Answering (Q&A) tasks under the adversary behaviors hierarchy in MITRE ATT&CK® and develop an automated Q&A generation pipeline that converts newly-collected CTI reports into the pre-defined format, enabling its extensibility on the corpus side.
+Following Reasoning Scalability, we further evaluate several LRMs and reasoning distillation strategies, which function well in general domains, to identify their strengths and limitations on the specialized attack sequence analysis task, providing helpful insights for future research in this area.
+To achieve Domain-Specific Epistemic Expandability, we aggregate cybersecurity-related knowledge from some existing benchmarks and embed it into LLMs via several post-training strategies to examine their epistemic expandability at the model level.
+Moreover, we also extend beyond the conventional zero-shot setting by introducing context-based and RAG-empowered settings, which pertinently assess LLMs’ epistemic expandability when injecting domain-specific cybersecurity knowledge at the semantic and representation levels.
+Our contribution are as follows: (I) We introduce AttackSeqBench, the pioneering benchmark that systematically evaluates the ability of existing LLMs, LRMs, and post-training strategies to analyze attack sequences across diverse inference settings and multi-level tasks.
+(cf.
+Section 2) (II) We quantitatively demonstrate that existing LRMs fail to substantially outperform LLMs on attack sequence analysis and perform markedly worse in most cases, a contrast to their advantages observed in domains such as mathematics and coding.
+(cf.
+Section 3.3) (III) We offer a comprehensive analysis of how parameterization and parameter scale affect existing models’ attack sequence analysis, and further examine why current LRMs and RAG underperform on this specialized task.
+This work uncovers the fundamental limitations of current models in attack sequence analysis and provides actionable insights to guide future research in this domain.
+(cf.
+Section 3.4 and Section 3.5)
+2 Dataset Construction and Verification
+Goal of our AttackSeqBench Our goal is to explore the capability of diverse types of LLMs in attack sequence understanding.
+Through comprehensive evaluation across various tasks and settings, we emphasize the strengths and limitations of existing LLMs, offering the promising yet underexplored directions.
+2.1 Problem Definitions
+CTI report understanding aims to convert the unstructural report into the structural formulation and further comprehend the sequential attack patterns of the structured threat intelligence knowledge.
+To achieve this, we define the attack sequence $S$ as the progression of adversarial behaviors described in a given CTI report, characterized by the logical order of TTPs based on their associated tactics within the ATT&CK KB.
+Formally, we utilize a 4-tuple to represent $S$ as $S=(T,E,P,O)$, where:
+    Tactic Sequence $T$: An ordered list of ATT&CK tactics, such that $T=(t_{1},\ldots,t_{n})$, where $t_{k}$ is the $k$-th tactic in the sequence.
+    Technique Mappings $E$: The set of ATT&CK techniques / sub-techniques in $S$, where $E(t_{k})=\{e_{1,k},\ldots,e_{i_{k},k}\}$ denote all the techniques / sub-techniques that belong to tactic $t_{k}$.
+[…]
+    CTI Outline $O$: A textual summary of the organized TTPs based on the order of Tactic Sequence $T$, such that $O=(o_{1},\ldots,o_{n})$, where $o_{k}$ refers to the summarized text associated with tactic $t_{k}$.
+Figure 2: Overview of our automated QA dataset construction pipeline.
+2.2 Dataset Construction
+As illustrated in Figure 2, we first construct attack sequences using the extracted TTPs and CTI outline from CTI reports.
+Then, we generate Q&A pairs based on the constructed attack sequences and refine them based on a tailored evaluate criteria before populating the QA dataset.
+Attack Sequence Construction.
+To efficiently and massively extract threat intelligence from these unstructured reports, we utilize a set of 408 CTI reports from various security vendors (Cisco Talos Intelligence Group, 2025; Microsoft, 2025) to construct attack sequences that accurately reflect the behaviors of real-world APTs.
+Specifically, we utilize a LLM-based KG construction framework (Zhang et al., 2025c) to automatically parse CTI reports, extract TTPs from each chunk into three level, generate CTI outlines, and combin them to construct the attack sequences $S$.
+Notably, we exclude CTI outlines which contains less than two ATT&CK tactics in attack sequence construction as they are unlikely to detail attack patterns observed in real-world cyber attacks.
+Q&A Generation.
+Inspired by the remarkable question generation abilities of LLMs across multiple domains (Alam et al., 2024; Zhang et al., 2024b; Mucciaccia et al., 2025), we adopt an answer-aware question generation approach using GPT-4o (OpenAI, 2024a).
+To elaborate, we first instruct the LLM to generate a seed Q&A pair for each tactic, technique, and group of procedures with the given attack sequence.
+Furthermore, we utilize the model’s In-Context Learning ability to generate the more relevant Q&A pairs (Dong et al., 2024), by including the CTI outline and few-shot Q&A examples in the question generation prompt (cf.
+Appendix C.1).
+For the Multiple-Choice Question (MCQ) tasks, we adopt a rule-based approach to select three choices as distractors.
+Specifically, we select a adjacent tactic of tactic $t_{k}$ within the Tactic Sequence $T$ (i.e., $t_{k+1}$ or $t_{k-1}$) and randomly select two tactics from the ATT&CK KB in AttackSeq-Tactic.
+Regarding AttackSeq-Technique, we follow the STARC annotation framework (Berzak et al., 2020) to define the selection rules with the given technique $e_{j,k}$: (1) The first technique belongs to the same tactic $t_{k}$ but not present in the given attack sequence, i.e., $e_{i,k}\notin E(t_{k})$; (2) The second technique is supported by the given attack sequence but belongs to another tactic $e_{j,j}\notin E(t_{k})$; (3) The third technique comes from a randomly chosen tactic that is not supported by the given attack sequence.
+Regarding the Yes-No Question tasks, we first instruct LLM to generate questions for each group of procedures within the attack sequence to construct the AttackSeq-Procedure-Yes.
+Next, we randomly sample its 70% questions to generate the negative question samples.
+Specifically, we design two types of Yes-to-No transferring strategies as follows: (1) Negation of temporal prepositions, i.e., changing “before” to “only after” and/or “after” to “only before”, such that the modified question contradicts the given attack sequence (Rajpurkar et al., 2018); (2) Replacement of the procedures in the question with another procedures that is not supported by the given attack sequence.
+Q&A Refinement.
+While LLMs possess remarkable text generation capabilities, these models may deviate from the requirements specified in users’ instructions (Joshi et al., 2025), resulting in the conflict between the generated questions and the order of TTPs in attack sequences.
+Inspired by the Self-Refine framework (Madaan et al., 2023), we design a refinement criteria to iteratively refine the initial questions via the same LLM.
+To perform a holistic evaluation, we introduce six aspects below that emphasizes the question’s linguistic (i.e., Clarity) and task-oriented properties (Fu et al., 2024).
+Here, we divide the task-oriented aspects into three categories: (1) Question Complexity (i.e., Answerability); (2) Content Alignment (i.e., Relevance, Consistency, Answer Consistency); (3) Attack Sequence Alignment (i.e., Logical) (cf.
+Appendix A.2).
+Considering the foundational role of Answerability in benchmark design, we first instruct the LLM to assess whether each question satisfies this criterion—specifically, whether the CTI report provides direct evidence supporting a correct answer that is clearly preferable to alternatives.
+Questions failing this requirement are discarded from the next step.
+Secondly, the LLM is instructed to evaluate the questions based on the remaining five aspects, providing a numerical score (out of five) and feedback for each aspect.
+Lastly, the LLM is prompted to refine the questions based on the feedback given (cf.
+Appendix C.2).
+We repeat this three-step process once more to improve the quality of the questions, the questions with full numerical scores are added to our final QA dataset.
+After the Q&A refinement, the data volume of four sub-tasks in our AttackSeqBench reduce from 2,158/2,937/1,393/3,249 to 1,697/1,917/1,223/1,412, filtering out 35.82% of the original samples that cannot satisfy the defined selection criteria.
+Additionally, we further illustrate the top-10 ATT&CK tactics and techniques within our dataset in Figure 7(a) and 7(b) respectively.
+The most frequent tactic and technique in the figure is associated with a key objective of APTs, highlighting the relevance of our Q&A dataset in capturing attack sequences based on real-world cyber attacks.
+Table 1: Evaluation results of four sub-tasks on human evaluation and automatic evaluation.
+Task Human Evaluation Automatic Evaluation
+Hum. Perf. Ans. Cla. Log Rel. Con. Ans. Con. Ans. Cla. Log Rel. Con. Ans. Con.
+Tactic 0.51 4.30 4.36 4.45 4.56 4.46 4.44 4.52 4.65 4.79 4.84 4.65 4.76
+Technique 0.71 4.09 4.21 4.40 4.45 4.44 4.41 4.10 4.40 4.62 4.63 4.39 4.59
+Procedure-Yes 0.74 4.88 4.70 4.88 5.00 4.81 4.94 4.02 4.06 4.61 4.47 3.78 3.89
+Procedure-No 0.56 4.55 4.84 - - 4.82 4.66 3.29 3.66 - - 2.77 3.25
+Average 0.63 4.45 4.53 4.57 4.67 4.63 4.61 3.98 4.19 4.67 4.64 3.90 4.12
+2.3 Dataset Evaluation
+LLMs demonstrate strong potential in solving complex tasks, but they inevitably exhibit even severer hallucinations, which has become a widely recognized concern in the research community.
+To address this, we adopt a hybrid approach towards evaluating the quality of the constructed QA dataset using the criteria defined in our Q&A refinement (cf.
+Section 2.2).
+We design 5-point Likert scales for each of the evaluation criterion, where higher scores indicate better alignment.
+Human Evaluation.
+We first randomly sample 35 questions from each sub-task to construct a question set for human evaluation.
+Three cybersecurity experts are then invited to answer and evaluate the quality of our Q&A dataset based on the six aspects defined in Section 2.2.
+Based on Table 1, we observe that the average Human Performance (abbreviated as Hum.
+Perf.) equals 0.63, suggesting that these questions is challenging and deducible even for individuals with domain expertise.
+Notably, AttackSeq-Procedure-No is derived from AttackSeq-Procedure-Yes through Yes-to-No transferring strategies, that is, its Logical and Relevance are inherently misaligned with the attack sequence, and we therefore do not evaluate these two aspects.
+Furthermore, the human evaluation shows consistently high average scores across all aspects, ranging from $4.45$ to $4.61$ out of $5$, indicating that the generated Q&A are easy to comprehend and well aligned with the attack sequences.
+Automatic Evaluation.
+To alleviate the laborious task of human evaluation, recent works (Zheng et al., 2023; Yao et al., 2024) have shown considerable effectiveness of LLM-as-a-Judge framework in aligning with human preferences within specific domain, including cybersecurity (Xu et al., 2024a).
+We incorporate G-Eval (Liu et al., 2023), a Chain-of-Thought (CoT) (Wei et al., 2022) and form-filling paradigm, to systematically assess the quality of generated Q&A pairs.
+Specifically, we design individual prompts for each aspect in the evaluation criteria that includes its definition and the scoring guideline based on the same 5-point Likert scale in Human Evaluation (cf.
+Table LABEL:tab.annotation_instructions in Appendix A.2).
+Then we instruct GPT-4o rate the Q&A for each aspect based on the evaluation criterion and the correct answer from the ATT&CK KB (cf.
+Appendix C.3).
+Based on Table 1, we observe that Logical and Relevance are the highest rated aspects, reinforcing the LLM’s ability to construct questions that follow the logical order of attack sequences (cf.
+).
+The fact that automatic evaluation scores are lower than human evaluation scores further indicates that answering questions correctly in our AttackSeqBench is more challenging for LLMs than for domain experts.
+3 Benchmark and Experiments
+Figure 3: Overview of the three benchmark settings that exhibit varying levels of contextual information given to the LLM.
+3.1 Benchmark Settings
+As illustrated in Figure 3, we elaborate on three benchmark settings with varying levels of contextual information: (1) Zero-shot setting, (2) Context setting, and (3) RAG-empowered setting.
 Zero-shot Setting.
-Aside from excelling in semantic understanding, previous works have also demonstrated LLMs’ impressive zero-shot reasoning ability across various downstream tasks (Hou et al., 2024; Kojima et al., 2022).
-To evaluate the LLMs’ capabilities in zero-shot attack sequence analysis, we omit the masked CTI outline from the prompt in the Regular setting.
-Instead, we instruct the LLM to select the most plausible answer choice based on the attack sequence described in the question’s body.
+Motivated by the significant zero-shot reasoning abilities of LLMs across several downstream tasks (Hou et al., 2024; Kojima et al., 2022), we directly utilize the system prompt and the Q&A pairs to evaluate the LLMs’ performance on three tasks based on their inherent knowledge.
+Context Setting.
+Considering the existing context-aware work (Ma et al., 2023; Jin et al., 2024), we also organize this context setting to evaluate LLMs’ Domain-Specific Epistemic Scalability at the semantic level.
+Here, we remove the corresponding summary of the ground truth tactic $t_{k}$ from the CTI outline $O$ to construct the masked CTI outline $O_{m}$, where $O_{m}\!=\!O\setminus o_{k}$.
+Afterwards, the LLM will be instructed to use the corresponding $O_{m}$ to answer the question, highlighting its potential to perform abductive reasoning to determine the most plausible TTP in attack sequence.
 RAG-empowered Setting.
-Retrieval Augmented Generation (RAG) has gained traction due to its ability to enhance LLMs’ reliability and mitigate hallucinations (Zhang et al., 2025a).
-We introduce this benchmark setting to assess the impact of providing the relevant ATT&CK TTPs in the question’s body as contextual information, by leveraging a retrieval-based approach and the ATT&CK KB.
-We first crawl the description and the example procedures of each TTP from the Enterprise ATT&CK Matrix  [^1] https://attack.mitre.org/matrices/enterprise/.
-Next, we split and embed the text chunks alongside their associated ATT&CK tactic and technique as metadata.
-Afterwards, we retrieve relevant chunks based on the question’s body and extract the associated ATT&CK TTPs from the metadata.
-Lastly, we instruct the LLM to utilize the retrieved TTPs and deduce the correct answer.
-By doing so, we enable the LLM to focus on determining the correct attack sequence rather than interpreting the TTPs in the question.
-Table 2.
-Performance (Accuracy) comparison of various open-source and proprietary LLMs (both fast-thinking and slow-thinking reasoning LLMs) under three settings (i.e., Regular, Zero-Shot, RAG) on three datasets (i.e., Tactic, Technique, Procedure).
-For each column, bold number indicates the best performance, and underlined number corresponds to the second best performance.
+Previous studies have demonstrated the Retrieval Augmented Generation (RAG) can significantly enhance the reliability pf LLMs and mitigate hallucinations (Zhang et al., 2025a).
+Here, we also design the RAG-empowered setting to evaluate the Domain-Specific Epistemic Scalability of LLMs at the representation level.
+This leverages the LLMs’ in-context learning ability to learn the associations between the entity in the question’s body and the relevant TTPs, thereby decomposing the problem and eliciting its stronger reasoning ability (Wu et al., 2022).
+3.2 Implementation Details
+To investigate the CRU capability of existing models, we evaluate seven LLMs (i.e., LLaMa-3.1-8B (Grattafiori et al., 2024), ChatGLM-4-9B (GLM et al., 2024), Qwen-2.5-3B, Qwen-2.5-14B, Qwen-2.5-32B (QwenTeam, 2024), Llama-3.3-70B (Grattafiori et al., 2024), and GPT-4o (OpenAI, 2024a) ) and five LRMs (i.e., R1 (Llama-3.1-8B), R1(Qwen-2.5-14B), R1(Qwen-2.5-32B) (DeepSeek-AI, 2025), QWQ-32B (Team, 2024) and GPT-o3-mini (OpenAI, 2025)) on AttackSeqBench.
+We also utilize four post-training strategies (i.e., SFT (Zhang et al., 2023), RD (Huang et al., 2024), RLIF (Zhao et al., 2025) and RLVR (DeepSeek-AI, 2025)) to embed security knowledge into LLMs to evaluate the Domain-Specific Epistemic Scalability of our AttackSeqBench (cf.
+Appendix A.4).
+Here, we measure the performance with accuracy Acc=n/MAcc\!=\!n/M, where $n$ is the correctly-answered number of questions and $M$ is the total number.
+[^1] Due to the page limitation, we introduce the complete implementation details in Appendix A.5 and A.6.
+Table 2: Performance comparison of various LLMs, LRMs and post-training strategies across three benchmark tasks and settings. Bold and underlined denote best and second-best in each column.
 LLMs AttackSeq-Tactic AttackSeq-Technique AttackSeq-Procedure
-Regular Zero-Shot RAG Regular Zero-Shot RAG Regular Zero-Shot RAG
-Fast-thinking LLMs
+Zero-Shot Context RAG Zero-Shot Context RAG Zero-Shot Context RAG
 […]
-Qwen-2.5-7B-Instruct 0.5121 0.4903 0.4761 0.6693 0.6568 0.6067 0.6584 0.5184 0.4941
-Llama-3.1-8B-Instruct 0.4744 0.5085 0.4926 0.6260 0.6333 0.5827 0.6577 0.5328 0.5230
-ChatGLM-4-9B-Chat 0.4885 0.4979 0.5009 0.6275 0.6109 0.6030 0.641 0.5408 0.5131
-Llama-3.3-70B-Instruct 0.6588 0.5551 0.5681 0.7058 0.6797 0.7037 0.6903 0.5469 0.5279
-Qwen-2.5-72B-Instruct 0.5793 0.5863 0.5657 0.5430 0.7162 0.6959 0.7188 0.6285 0.6030
-GPT-4o-mini 0.6517 0.6005 0.5692 0.7387 0.7058 0.7021 0.6968 0.5491 0.5340
-GPT-4o 0.6093 0.5740 0.5787 0.6755 0.6995 0.7188 0.7359 0.6755 0.6353
-Slow-thinking Reasoning LLMs
 […]
-DeepSeek-R1-Distill-Qwen-32B 0.5421 0.5698 0.5504 0.5879 0.5816 0.5816 0.6945 0.6258 0.5852
-QWQ-32B-Preview 0.5345 0.3377 0.4638 0.5112 0.3918 0.5342 0.7036 0.5696 0.5457
-GPT-o3-mini 0.4643 0.5445 0.5215 0.5373 0.5915 0.5822 0.6854 0.6877 0.6459
-4.2. Experimental Setting
-4.2.1. Implementation Details
-To demonstrate the robustness of AttackSeqBench, we evaluate eight fast-thinking LLMs: Mistral-7B-Instruct-v0.3, Qwen-2.5-7B-Instruct, Llama-3.1-8B-Instruct, ChatGLM-4-9B-Chat, Llama-3.3-70B-Instruct, Qwen-2.5-72B-Instruct, GPT-4o-mini and GPT-4o.
-Additionally, we assess four slow-thinking reasoning LLMs: DeepSeek-R1-Distill-Llama-8B, DeepSeek-R1-Distill-Qwen-32B, QWQ-32B-Preview and GPT-o3-mini.
-For the RAG-empowered benchmark setting, we utilize a hybrid retriever due to its considerable performance in existing work (Zhang et al., 2024b).
-We combine Okapi BM25 (Robertson and Zaragoza, 2009) as the sparse retriever and a dense retriever based on BGE-EN-ICL (Li et al., 2024) (shortened as BGE) and Chroma DB  [^2] https://www.trychroma.com/, along with a re-ranker.
-The detailed implementation is in Appendix B.1.
-4.2.2. Evaluation Metrics
-We evaluate the Accuracy (abbreviated as Acc in the following sections) for all three tasks in AttackSeqBench (Liu et al., 2024; Zhang et al., 2024b; Alam et al., 2024).
-Specifically, given the total number of questions $M$ and the number of questions $n$ with the correct answer choice chosen by the LLM, the accuracy can be calculated as $Acc=n/M$.
-4.3. Performance Comparison
-We conduct comprehensive analysis using eight fast-thinking LLMs and four slow-thinking reasoning LLMs in Table 2 and elaborate on three key findings: (1) None of the LLMs clearly outperforms other LLMs across all the tasks.
-While GPT-4o demonstrate strong performance in AttackSeq-Technique and AttackSeq-Procedure, its performance falls behind Llama-3.3-70B-Instruct and GPT-4o-mini in AttackSeq-Tactic.
-This implies the complexity and the diversity of our three sub-tasks to some extent.
-(2) Among these three tasks, we observe that LLMs collectively demonstrate the lowest overall performance for the AttackSeq-Tactic task.
-This implies LLMs struggle to determine the high-level objectives in an attack sequence, even with the CTI outline given in the context.
-This may arise due to the abstract nature of the descriptions of tactics which causes the LLM to misinterpret the correct tactic.
-(3) The AttackSeq-Procedure task yields the highest performance results across all the LLMs.
-This suggests that LLMs are able to leverage its procedural knowledge to deduce the plausibility of a group of procedures in an attack sequence.
-However, this may also be attributed by the LLMs’ tendency to favor ”yes” in boolean questions when given insufficient context (Wen et al., 2024), which causes high accuracy for AttackSeq-Procedure-Yes.
-We explore this phenomenon further in Section 4.4.1.
-4.4. Model Analysis
-Table 3.
-Performance comparison of four representative LLMs between AttackSeq-Procedure-Yes and AttackSeq-Procedure-No sub-tasks in each benchmark setting.
-For each column, the bold value indicates the best performance.
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+3.3 Performance Comparison
+Comparison between diverse groups of LLMs.
+As shown in Table 2, we notice that: Although LLMs generally follow the scaling laws in our AttackSeqBench, none of the LLMs consistently outperforms the others, and the optimal LLM varies diverse tasks.
+For instance, the best-performing models under the zero-shot setting across the three benchmark tasks are Qwen-2.5-32B, Qwen-2.5-32B, and GPT-o3-mini, respectively.
+This suggests that current models may not possess explicit security-specific knowledge, as relevant information in pretraining corpus is likely overshadowed by general-domain content.
+Moreover, most models consistently perform worst in AttackSeq-Tactic compared to the other two tasks, mirroring the human evaluation results in Section 2.3 and underscoring the common challenge faced by both human experts and LLMs in tactical inference.
+Furthermore, from Table 3, we can observe that compared to the zero-shot setting, all models exhibit substantial performance gains on AttackSeq-Procedure-No under the context setting, indicating the importance of contextual information in identifying highly implausible actions within attack sequences.
+As defined in Appendix A.3, AttackSeq-Procedure-No is inherently more complex and reasoning-demanding than AttackSeq-Procedure-Yes, as it requires models to overcome the helpful-only bias and explicitly answer ‘No’ to disprove the plausibility of procedures occurring within the attack sequence.
+This explains why LRMs with stronger reasoning ability outperform in AttackSeq-Procedure-No compared to other tasks, underscoring the benchmark’s emphasis on Reasoning Scalability.
+Finally, most post-training strategies substantially improve the performance of its base LLM, particularly in zero-shot settings that rely solely on internal knowledge.
+However, their performance still lags behind instructive LLMs equipped with task-adapted prompts.
+This highlights a promising direction: Designing specialized post-training strategy to embed security-related knowledge into existing LLMs, thereby advancing the development of domain-specific models for cybersecurity.
+Table 3: Performance comparison of two LLMs and two LRMs on AttackSeq-Procedure-Yes and AttackSeq-Procedure-No in each benchmark setting, where the bold value indicates the best performance of each column.
 LLMs AttackSeq-Procedure-Yes AttackSeq-Procedure-No
-Regular Zero-Shot RAG Regular Zero-Shot RAG
-Llama-3.1-8B 0.7899 0.9518 0.9460 0.5432 0.1700 0.1565
-GPT-4o 0.7146 0.9403 0.9542 0.7542 0.4462 0.3591
-R1 (Llama-8B) 0.8332 0.9526 0.9477 0.4341 0.1161 0.1062
-GPT-o3-mini 0.5789 0.7760 0.8242 0.7776 0.6112 0.4915
-4.4.1. Impact of Contextual Information
-To comprehensively investigate the impact of the contextual knowledge in the given CTI outlines for analyzing attack sequences in LLMs, we compare the performance of LLMs between the Regular and Zero-Shot benchmark settings.
-Based on in Table 2, we observe that both fast-thinking and slow-thinking reasoning LLM’s accuracy generally increases across each task in the  Regular setting.
-Among the fast-thinking LLMs, the performance of models with smaller number of parameters, e.g., Llama-3.1-8B-Instruct (shortened as Llama-3.1-8B), have minimal or no improvement for AttackSeq-Tactic and Attack-Technique tasks.
-This indicates that the baseline LLMs with smaller parameters may not fully leverage the capabilities of ICL within these benchmarking tasks.
-Similarly for slow-thinking reasoning models, DeepSeek-R1-Distill-Llama-8B (shortened as R1 (Llama-8B)) has no improvement in the Regular setting as compared to the Zero-Shot setting.
-On the other hand, Llama-3.3-70B-Instruct achieves comparable performance against the closed source models, i.e., GPT-4o and GPT-4o-mini.
-Among the three tasks, AttackSeq-Procedure has the highest improvement in LLMs from the Zero-Shot setting to the Regular setting.
-Based on Table 3, we observe a significant increase in the AttackSeq-Procedure-No sub-task.
-This suggests that the contextual information found in the masked CTI outline is essential in justifying procedures that are unlikely to happen.
-Table 4.
-Performance comparison between three embedding models.
-Eeb.
-M.
-refers to the embedding model used, while Tac, Tec and Pro refers to the three tasks in AttackSeqBench (i.e., Tactic, Technique, and Procedures) respectively.
-For each row, the bold value indicates the best performance.
-#Eeb. M. Task Llama-3.1-8B GPT-4o R1 (Llama-8B) GPT o3-mini
-BGE Tac 0.4926 0.5787 0.4532 0.5215
-Tec 0.5827 0.7188 0.5138 0.5822
-Pro 0.5230 0.6353 0.4968 0.6459
-OpenAI Tac 0.4915 0.4814 0.4632 0.5333
-Tec 0.5790 0.7178 0.5039 0.5962
-Pro 0.5195 0.6330 0.4979 0.6353
-ATT&CK- BERT Tac 0.4985 0.5622 0.4590 0.5292
-Tec 0.5790 0.7100 0.5018 0.5874
-Pro 0.5245 0.6209 0.4953 0.6398
-4.4.2. Effectiveness of RAG Strategies
-We analyze the capabilities of the RAG-empowered setting based on the results of AttackSeqBench in Table 2.
-Within the set of fast-thinking LLMs, GPT-4o achieves exceptional performance when leveraging RAG in AttackSeq-Tactic and AttackSeq-Technique tasks.
-Moreover, the LLMs exhibit the best performance in the AttackSeq-Technique among the three tasks in Table 2.
-This indicates that LLMs are able to leverage the retrieved TTPs to determine the correct technique.
-However, the RAG-empowered setting does not consistently yield effective results, as models that have less than or equal to 9B parameters have no improvement as compared to the Zero-Shot setting.
-This suggests that the information retrieved, i.e., the semantically similar ATT&CK TTPs based on the question’s body, is insufficient for smaller models.
-Furthermore, we evaluate the performance of four representative LLMs using two additional embedding models in Table 4.
-The two models are: OpenAI’s text-embedding-3-large (OpenAI, 2024c); ATT&CK BERT (Abdeen et al., 2023) which is an open-source embedding model fine-tuned for cybersecurity.
-Based on Table 4, BGE acheives the best performance across the three benchmarking tasks.
-Figure 4.
-Performance and model size comparison of nine representative LLMs in the AttackSeq-Tactic task.
-The size of bubble corresponds to inference time, where horizontal lines denote fast-thinking LLMs and vertical lines indicate slow-thinking reasoning LLMs.
-4.4.3. Computational Complexity Analysis
-As illustrated in Figure 4, we extensively compare the performance, model size and inference cost of nine open-source LLMs in the AttackSeq-Tactic task.
-We have the following observations: (1) In general, these LLMs exhibit the performance trend in sequential attack pattern understanding which aligns with the scaling law (Kaplan et al., 2020).
-(2) Comparing the performance of LLMs with different thinking types, we observe that slow-thinking LLMs cannot significantly and consistently outperform the other.
-One possible reason is that given the question and masked CTI outlines, fast-thinking LLMs can already deduce the correct answer without any elaborate reasoning.
-(3) Regarding the inference cost, we notice that slow-thinking LLMs do not incur more reasoning costs compared to fast-thinking LLMs.
-This prompts an intuitive insight: the small-scale slow-thinking reasoning LLM fine-tuned using AttackSeqBench may enhance its capability in attack sequence analysis significantly while maintaining the lower reasoning cost.
-4.4.4. Sampling Parameters Analysis
-We assess the impact of two sampling parameters on the performance of LLMs in AttackSeq-Tactic under the Regular setting, namely temperature and maximum output tokens as illustrated in Figure 5.
-We observe that increasing the temperature at different three intervals has minimal improvements for Llama-3.3-70B-Instruct, while resulting in negative performance for the remaining three LLMs.
-This indicates that the most optimal performance occurs when the temperature is kept at 0.0.
-On the other hand, we observe a considerable improvement when increasing the number of tokens for slow-thinking reasoning models.
-The accuracy of Deepseek-R1-Distill-Llama-8B improved by 8.46% as the maximum output tokens increased from 1024 to 4096, while Deepseek-R1-Distill-Qwen-32B achieve a 24.9% improvement for the same increment.
-However, further increasing the token limit to 8192 yields diminishing returns for both models.
-This highlights the importance of carefully tuning the maximum output tokens parameter to optimize performance in slow-thinking models while considering the associated costs incurred (Wang et al., 2024).
-Figure 5.
-Parameter analysis on temperature (a) and maximum output tokens (b) in AttackSeq-Tactic task using the Regular setting, we perform analysis on four representative LLMs using different colors to represent them.
-Figure 6. The performance comparison of four representative LLMs on each tactic in each benchmarking task.
-Table 5. Reasoning path comparison between slow-thinking reasoning and fast-thinking LLMs in Zero-Shot setting based on the AttackSeq-Procedure task in Zero-Shot setting.
-Question: After delivering spearphishing emails with spyware disguised as a PDF file to gain initial access, is it likely the authors of Hacking Team’s spyware established communication with the identified C&C servers before using VMProtect for obfuscation to evade detection?
-Answer Choices: A) Yes  B) No (Correct Answer)
-DeepSeek-R1-Distill-Qwen-32B Qwen-2.5-72B-Instruct
-First, spearphishing emails are emails that look legitimate, which are actually malicious.
-In this case, the PDF is a disguise for spyware.
-So the first step is the delivery of the malicious payload via email.
-(Knowledge Retrieval) Once the victim opens the PDF, the spyware is executed.
-The next step in an attack chain is usually establishing communication with a command and control (C&C) server.
-Now, the question mentions VMProtect being used for obfuscation.
-However, sometimes obfuscation can be applied after deployment to maintain persistence or evade detection during operation.
-(Breakdown of Attack Sequence) But in the context of the attack sequence described, the spyware has already been delivered and executed.
-The next logical step would be for the spyware to reach out to the C&C server to report back and receive instructions.
-(Self-Correction) If the attackers were to use VMProtect at this point, it might be to re-obfuscate the spyware or to protect it further, but that seems a bit out of order.
-(Reasoning using Attack Sequence) So the answer would be No.
-(Correct Conclusion) The sequence includes: Delivering spearphishing emails with spyware disguised as a PDF file to gain initial access.
-Establishing communication with C&C (Command and Control) servers.
-Using VMProtect for obfuscation to evade detection.
-(Breakdown of Attack Sequence) The spearphishing emails with a PDF file are used to gain initial access to the target system.
-(T1566.002: Spearphishing Attachment).
-Once the initial access is gained, the malware typically establishes communication with the C&C servers to receive further instructions or exfiltrate data.
-(T1071: Application Layer Protocol).
-VMProtect is a tool used to obfuscate the malware to evade detection by security software.
-(T1027: Obfuscated Files or Information).
-(Knowledge Extraction) The logical sequence of events would be: 1.
-**Initial Access:**, 2.
-**C&C Communication:**, 3.
-**Obfuscation:**.
-(Incorrect Attack Sequence Order) Based on the logical sequence of TTPs, it is likely that the authors of Hacking Team’s spyware established communication with the C&C servers before using VMProtect for obfuscation to evade detection.
-Final Answer: Yes (Incorrect Conclusion)
-4.5. In-depth Analysis
-4.5.1. Impact of TTP Temporal Position
-To better understand the relationship between LLMs and attack sequences, we conduct fine-grained analysis on the performance of four representative LLMs under the Regular benchmark setting and across the ATT&CK tactics (refer to Figure 6).
-Additionally, we identify four categories to group the tactics into overarching attack phases as shown in Figure 6: Initial Intrusion, Exploitation, Stealth Expansion, Objective Orchestration.
-Our findings are: (1) Based on the standard deviation range across various tactics in each LLM, we observe that the range is the largest for AttackSeq-Tactic, indicating that the LLM’s performance largely varies for this task across different tactics.
-In contrast, the SD range is the smallest for AttackSeq-Procedure which suggests more stability in the LLMs’ performance.
-This highlights the challenge for LLMs to determine the correct tactics in attack sequences, as compared to procedural knowledge needed to correctly determine techniques and procedures.
-(2) GPT-4o consistently outperforms other models across most tasks, particularly in the in the Exploitation and Stealth Expansion phases.
-On the other hand, Llama-3.1-8B-Instruct and GPT o3-mini achieve comparable performance against GPT-4o in Initial Intrusion and Objective Orchestration phases, especially at the start and end of the sequence (i.e., Reconnaissance and Impact respectively).
-This inconsistency suggests that TTPs within certain ATT&CK tactics are more challenging for the LLM to interpret, likely due to the rarely observed tactics in real-world CTI reports.
-More detailed distribution of the performance can be found in Section B.2 of the Appendix.
-4.5.2. Inference Path Analysis on Slow-thinking Reasoning LLMs
-We perform a case study based on a question from AttackSeq-Procedure-No and compare the difference in the reasoning process in the output of two open-source LLMs: DeepSeek-R1-Distill-Qwen-32B and Qwen-2.5-72B-Instruct in the Zero-Shot setting.
-We choose not to use GPT-o3-mini in our analysis as the Batch API does not return the reasoning tokens.
-Based on Table 5, we observe that both models are able to breakdown the attack sequence and extract the relevant attack patterns.
-The reasoning output tokens suggest that DeepSeek-R1-Distill-Qwen-32B is able to retrieve relevant knowledge based on the attack sequence described in the question.
-However, Qwen-2.5-72B-Instruct outputs the incorrect order of attack patterns as stated in the question.
-Furthermore, it highlights the slow-thinking reasoning LLMs’ ability to perform Self-Correction (Kumar et al., 2024), enabling the model to elicit stronger reasoning capabilities and justify why the attack sequence stated in the question is incorrect.
-Notably, the slow-thinking reasoning LLM considers both incorrect and correct attack sequences before deciding on a conclusion, highlighting the potential in slow-thinking reasoning LLMs in zero-shot tasks.
-5. Outlook and Conclusion
-Outlook.
-AttackSeqBench extends beyond existing CRU-related benchmarks by emphasizing sequential patterns in cyber attacks, opening up numerous research avenues for related researchers.
-For instance, AttackSeqBench can be seamlessly utilized for the complex attack path modeling task as we provided in Section 4.5.1.
-Moreover, AttackSeqBench can also be leveraged to recognize the attack intention due to its temporal characteristic.
-Inspired by the ubiquitous causality in real-world scenarios, we believe that integrating the logical attack sequence with the causal inference is a promising research direction, discovering the critical causal relationships within the attack sequences and enabling more accurate predictions of the attackers’ potential action.
-Given the predicted attack intentions and next action, applying our AttackSeqBench to optimize the defense strategy is not only intuitive but also reasonable.
-Finally, the constructed dataset can be reformulated as the instruction-tuning data, facilitating the post-training phase of future security-related LLMs.
-Conclusion.
-In this work, we introduced AttackSeqBench, a novel benchmark designed to assess the LLMs’ ability to analyze attack sequences from real-world CTI reports.
-We propose an automated dataset construction pipeline, enabling our benchmark to be easily extended as cyber threats continue to evolve.
-Our benchmark incorporates three tasks that encompasses the varying aspects of adversarial behavior, allowing a comprehensive analysis of the LLMs’ understanding and reasoning capabilities in attack sequence analysis.
-We performed extensive experiments on a diverse set of fast-thinking and slow-thinking reasoning LLMs, offering crucial insights into the abilities and limitations of LLMs in understanding attack sequences.
-Our work opens up a meaningful new direction towards LLM-driven CRU in real-world cybersecurity operations.
+Zero-Shot Regular RAG Zero-Shot Regular RAG
+Llama-3.1-8B 0.9128 0.7572 0.8858 0.2434 0.6216 0.2111
+GPT-4o 0.9469 0.9567 0.8831 0.4426 0.2698 0.4143
+R1 (Llama-8B) 0.9332 0.8427 0.9191 0.1508 0.4417 0.1792
+GPT-o3-mini 0.7612 0.7048 0.7408 0.6303 0.6678 0.5552
+Comparison on Contextual Information.
+Comparing the performance of LLMs across three benchmark settings in Table 2, we can observe that: In general, the Context setting consistently outperforms Zero-Shot and RAG settings across most benchmark tasks, with the advantage more pronounced in larger LLMs.
+Taking the Qwen-2.5 series as an example: performance shifts from zero-shot being optimal in Qwen-2.5-3B (0.4467 vs.
+0.4614 vs.
+0.3296) to context-setting being optimal in Qwen-2.5-32B (0.6195 vs.
+0.5903 vs.
+0.5154) in the AttackSeq-Tactic task, with Qwen-2.5-14B showing the transition in between.
+This phenomenon is reasonable as larger LLMs possess more extensive internal knowledge, and task-specific context further enhances their effectiveness and robustness within the specific domain.
+Moreover, both LLMs and LRMs consistently fail to reach optimal performance under the RAG-empowered setting.
+This indicates that naive retrieval integration may introduce additional noise instead of enhancing results, underscoring the requirement for more advanced retrieval-augmented approaches.
+We further investigate its limitation in Section 3.5.2.
+3.4 Robustness Analysis
+Figure 4: Parameter sensitivity analysis on (a) Temperature and (b) Max Output Tokens in AttackSeq-Tactic under the zero-shot setting.
+3.4.1 Parameter Sensitivity Analysis
+Regarding the parameter sensitivity, we investigate the impact of temperature and maximum output tokens on LLMs and illustrate them in Figure 4(a) and Figure 4(b) respectively.
+Firstly, we observe that increasing the temperature from 0 to 1 causes a sharp performance drop in smaller LLMs, while larger LLMs remain relatively unaffected in attack sequence analysis.
+This may be because smaller LLMs lack discriminative power and oscillate among suboptimal answers, whereas larger ones generate more stable logits that preserve correct outputs even under smoothing, while aligning with the observations of previous work in general domain (Renze, 2024; Li et al., 2025).
+On the other hand, we also releaze that increasing the token budget yields stable performance and output length for LLMs, whereas LRMs achieve significant gains in both performance and output tokens.
+Specifically, when the Max Output Tokens increase from 1,024 to 4,096, R1 (LLaMA-8B) and R1 (Qwen-32B) improve accuracy by 13.29% and 16.74%, with average output tokens increasing by 37.35% and 43.28%, respectively.
+However, when the token budget is further increased to 8,192, LRMs exhibit diminishing returns: average output tokens increase by 13.27% and 9.83%, while accuracy improves only by 0.29% and 0.46%.
+This highlights the importance of carefully tuning the maximum output tokens parameter to optimize performance in LRMs while considering the associated costs incurred (Wang et al., 2024).
+3.4.2 Computational Complexity Analysis
+Figure 5: Computational complexity analysis of seven LLMs and four LRMs in AttackSeq-Tactic under the regular setting.
+The size of bubble represents inference time, where zigzag lines denote LLMs and cross hatch lines indicate LRMs.
+As illustrated in Figure 5, we extensively compare the performance, model size and inference cost of several open-source LLMs and LRMs in AttackSeq-Tactic under the context setting.
+We can observe the following findings: (1) The performance trend of LLMs follows scaling law in our AttackSeqBench (Kaplan et al., 2020), particularly among LLMs within the same series; (2) Comparing the performance trend of LLMs with different thinking types, LRMs can not consistently outperform LLMs, yet they require significantly more inference time.
+This phenomenon is more pronounced when comparing LLMs and LRMs of the same model size, where distilled LRMs perform significantly worse than their base LLMs.
+This finding on inference time are consistent with existing studies in the general domain, whereas the performance results show a starkly contrasting trend (DeepSeek-AI, 2025).
+We further investigate this issue in Section 3.5.1.
+Table 4: A sample of reasoning path analysis between Qwen-2.5-32B (LLM) and R1(Qwen-2.5-32B) (LRM) on AttackSeq-Procedure-No under the zero-shot setting.
+We utilize different color to differentiate the diverse phases within a given attack sequence.
+Question: After retrieving the final payload from update.bokts[.]com, is it likely that CosmicStrand will interact with the Transport Device Interface to generate network traffic only before redirecting the entry point of the CSMCORE DXE driver to attacker-supplied code in the .reloc section during system startup?
+Answer Choices: A) Yes B) No (Correct Answer) Attack Sequence in Question: Retrieve final payload $\rightarrow$ Interact with TDI $\rightarrow$ Redirect DXE
+[…]
+LLM: Qwen-2.5-32B LRM: R1 (Qwen-2.5-32B)
+Breakdown of Attack Sequence: 1.
+The attacker retrieves the final payload from a specific domain (update.bokts[.]com).
+2.
+The attacker interacts with the Transport Device Interface to generate network traffic.
+3.
+The attacker redirects the entry point of the CSMCORE DXE driver to attacker-supplied code in the .reloc section during system startup.
+Knowledge Retrieval: 1.
+Retrieving the Final Payload: This is a common step in many malware attacks where the payload is fetched from a command and control (C2) server or a malicious domain.
+2.
+Generating Network Traffic: Interacting with the Transport Device Interface to generate network traffic could be a technique used to exfiltrate data or to communicate with a C2 server.
+However, this step seems out of place if the primary goal is to redirect the entry point of a driver.
+3.
+Redirecting the Entry Point: This is a sophisticated technique used in firmware or UEFI-based attacks to gain persistence and control over the system at a very low level.
+Generating network traffic (step 2) before redirecting the entry point (step 3) does not logically follow the typical sequence of such an attack.
+Final Answer: B: No (Correct Conclusion)
+Breakdown of Attack Sequence: The question mentions retrieving a payload from a specific domain, update.bokts[.]com.
+Next, interacting with the Transport Device Interface (TDI) to generate network traffic.
+Then, during system startup, it redirects the entry point of the CSMCORE DXE driver to attacker-supplied code in the .reloc section.
+Knowledge Retrieval: TDI is part of the Windows kernel, used for handling network device drivers.
+If a malware interacts with TDI, it might be setting up a backdoor or a way to communicate over the network.
+DXE drivers are part of the UEFI (Unified Extensible Firmware Interface) system, which runs before the operating system starts.
+Usually, after gaining initial access (like downloading a payload), the next steps involve establishing persistence and communication.
+So, using TDI to set up network communication makes sense as a next step.
+(Incorrect Initial Approach)
+Wait
+(Transitional Word), but the question says ”only before redirecting.” So, does the TDI interaction happen only before the redirection, meaning it’s a one-time thing?
+Or is it part of the ongoing process?
+But wait (Transitional Word), could the redirection happen before the network traffic?
+(Questioning Alternatives) I think I’m leaning towards yes, so the answer is A: Yes.
+(Incorrect Conclusion)
+3.5 In-depth Analysis
+3.5.1 Reasoning Path Analysis
+To explore the reason why existing LRMs perform worser than LLMs in attack sequence understanding, we provide a sample of reasoning path between Qwen-2.5-32B (LLM) and R1 (Qwen-2.5-32B) (LRM) on AttackSeq-Procedure-No under the Zero-shot setting in Table 4.
+We observe that both LLM and LRM can decompose the attack sequence into granular TTPs and retrieve relevant knowledge.
+Here, LLM successfully aligns the retrieved knowledge with the logical order of TTPs, thereby recognizing that generating network traffic before redirecting the entry point contradicts the traditional attack sequence.
+It shows that LLM may rely on more direct sequence-matching between retrieved knowledge and procedural logic, enabling them to avoid unnecessary reasoning detours.
+In contrast, despite demonstrating reflective reasoning steps, LRM misinterprets the temporal constraint (“only before”) and overemphasizes the plausibility of the TDI interaction.
+This overthinking within LRMs are also more prone to construct redundant reasoning loops and further incur reasoning misalignment, which may amplify minor misunderstandings into incorrect conclusions.
+Figure 6: Error distribution of randomly-sampled 100 incorrect responses from GPT-4o in AttackSeq-Technique under the RAG-empowered setting.
+3.5.2 Effectiveness of RAG Strategies
+To investigate why LLMs underperform in the RAG-empowered setting of our AttackSeqBench, we collect a candidate set where GPT-4o answers correctly in the zero-shot setting but fails under this setting.
+Specifically, we randomly sample 100 incorrect responses in AttackSeq-Technique, and classify them into four categories as shown in Figure 6.
+These four categories are: (1) Factual Error, meaning that LLM’s prediction contradicts the ground truth despite the correct retrieved content; (2) Over-reliance (Xia et al., 2024), meaning that LLM excessively refers to the retrieved content and fails to synthesize the attack sequence in the given question; (3) Irrelevant Retrieved TTP, which refers to incorrect predictions due to irrelevant retrieval to the given question; (4) Incorrect Answer Format, which refers to LLM’s failure to follow the output format specified in the prompt template.
+Our analysis reveals that 59% of errors stem from Factual Error, where the primary cause is the model’s failure to effectively integrate retrieved evidence into the reasoning chain.
+Rather than enhancing the inference process, the retrieved knowledge functions as noise to distort the output distribution, thereby inducing faulty reasoning and incorrect answers.
+Moreover, around 32% of the errors occur because LLMs treat retrieved knowledge as the absolute authority without validating them against the question intent or their internal knowledge.
+Consequently, the model often relies solely on correct but incomplete retrieval chunks, which leads to faulty results.
+Within the ATT&CK KB, the nuances of TTP descriptions introduce several overlaps and ambiguities, which account for 8% of cases where the embedding model to retrieve incorrect tactics and techniques.
+For example, technique T1574 – Hijack Execution Flow  [^2] https://attack.mitre.org/techniques/T1574/ is associated with three distinct tactics (i.e., Persistence, Privilege Escalation, and Defense Evasion), leading the model to misinterpret attack sequences in the given question.
+Enhancing the integration of retrieved knowledge with question intent and internal knowledge in RAG scenario, or investigating embedding methods capable of capturing fine-grained TTP semantics, holds promise for improving the effectiveness of attack sequence analysis.
+4 Related Work
+Automating CTI Report Understanding.
+With the increasing demands of cybersecurity operations and the breakthrough of LLMs, researchers have progressively explored their applicability within CRU (Zhang et al., 2024a).
+For instance, prior works have showcased the remarkable capabilities of LLMs in interpreting TTPs from the ATT&CK KB, surpassing the performance of some of the fine-tuned LMs with cybersecurity data (Fayyazi et al., 2024).
+Meanwhile, another line of work proposes LLM-driven threat intelligence Knowledge Graph construction frameworks, which utilize the threat-related entities and relations to describe CTI reports in a structural manner (Huang & Xiao, 2024; Cheng et al., 2024).
+However, the extent to which LLMs can understand and reason about the precise relations between adversary behavioral sequences described in CTI reports remains largely under-explored.
+In our work, we perform a holistic evaluation on various pre-trained LLMs, LRMs and fine-tuned LLMs in attack sequence analysis, from deducing high-level tactics to detailed procedures described in CTI reports.
+Benchmarking LLMs in Cybersecurity.
+Inspired by the remarkable open-world knowledge and complex inference ability within LLMs, various benchmarks have been proposed to evaluate its general capabilities in language understanding (Hendrycks et al., 2021b), math reasoning (Cobbe et al., 2021), code generation (Chen et al., 2021).
+Regarding the cybersecurity domain, researchers start to benchmark the abilities of LLMs under such specialized setting, such as ethical hacking and compliance (Liu, 2023; Tihanyi et al., 2024; Garza et al., 2023).
+Targeting CRU-related tasks, SEvenLLM (Ji et al., 2024) explores the abilities of LLMs in threat-related entities extraction and summarizing reports from security vendors.
+SecBench (Jing et al., 2025) evaluates the knowledge retention and logical reasoning abilities of existing pre-trained LLMs from multiple languages and dimensions.
+Meanwhile, CTIBench (Alam et al., 2024) introduces five benchmark tasks to explore the threat entity attribution and cause-tracing abilities of LLMs within the security context.
+However, these studies primarily rely on authoritative sources (e.g., textbooks, open standards) while overlooking real-world sources such as CTI reports.
+For instance, CTIBench solely incorporates a small-scale set of CTI reports in its dataset construction process for only one of its five benchmark tasks.
+Furthermore, these benchmarks remain insufficient for providing a comprehensive evaluation towards the LLMs’ ability to understand relations among adversarial behaviors described in CTI reports, thereby failing to accurately reflect their reasoning capabilities over attack sequences containing domain-specific semantics.
+In this paper, we construct attack sequences based on an extensive set of CTI reports, while emphasizing on the practical aspects of CRU, inferring various aspects of adversarial behaviors, in our proposed benchmark tasks.
+5 Conclusion
+The breakthrough of LLMs has shown promising potential across the cybersecurity domain, particularly in CTI understanding.
+Despite this, the applicability of LLMs in analyzing adversarial sequences remains largely unexplored.
+In this work, we propose AttackSeqBench, a benchmark tailored for assessing LLMs’ ability in understanding how adversaries operate through inferring TTPs based on attack sequences from real-world CTI reports.
+To cater to the evolving threat landscape, we design an automated Q&A construction pipeline that enables the Extensibility of our benchmark to new CTI reports.
+We further conduct extensive experiments across three settings with varying context availability, evaluating diverse LLMs, LRMs, and post-training strategies to verify its Reasoning Scalability and Domain-Specific Epistemic Expandability and thoroughly analyze their ability boundaries in attack sequence analysis.
+Our work opens up a new direction towards LLM-driven CRU, enabling effective threat intelligence mining through automation.
+Ethics Statement
+Our work utilizes publicly available CTI reports, while ensuring that no proprietary information is used.
+The dataset generation pipeline is designed to maintain the integrity and accuracy of adversarial behavior sequences without fabricating or misrepresenting cyber threats.
+Furthermore, human evaluation is conducted with careful consideration of evaluator expertise and potential biases, ensuring fairness and reliability in assessment.
+Reproducibility Statement
+To promote reproducibility, we release an anonymous repository that contains all resources necessary to replicate our study, including the original CTI reports, the dataset construction pipeline, the complete datasets for the three tasks (i.e., AttackSeq-Tactic, AttackSeq-Technique, and AttackSeq-Procedure), and the code for running the three benchmark settings (i.e., Zero-Shot setting, Context setting, and RAG-empowered setting).
+In addition, we also provide the comprehensive implementation details about the four post-training strategies in Appendix.
+Together, these resources ensure that our AttackSeqBench can be reliably reproduced and the reported results independently verified.
+Appendices
 Appendix A Dataset
-A.1. Dataset Distribution
-The distribution of questions in AttackSeqBench is shown in Table 6, we construct an almost equal proportions of questions in each task to allow fair evaluation in our benchmark.
-Based on Figure 7, we observe that the top three most frequent tactics occur in the middle of attack sequences, while the bottom two tactics occurs at the start and end of the attack sequence.
-The most frequent technique belongs to the Command and Control tactic, this is likely due to the common operations of APTs.
-Similarly, the most frequent ATT&CK technique is T1071-Application Layer Protocol  [^3] https://attack.mitre.org/techniques/T1071/, which is associated with the Command and Control tactic.
-Figure 7. Visualization of the distribution of top 10 tactics and techniques in our AttackSeqBench.
-A.2. Dataset Evaluation
-Inspired by existing work (Fu et al., 2024), we evaluate the quality of our QA dataset based on six dimensions below:
-  • Answerability: We check if there is direct evidence in the CTI outline that supports the correct answer, while clearly standing out as the best answer choice.
+A.1 Dataset Distribution
+Based on Figure 7, we observe that the top three most frequent tactics (i.e., Command and Control, Defense Evasion and Execution) occur in the middle of attack sequences, while the bottom two tactics (i.e., Exfiltration and Reconnaissance) occurs at the start and the end of the attack sequence.
+Similarly, the most frequent ATT&CK technique is T1071-Application Layer Protocol  [^3] https://attack.mitre.org/techniques/T1071/, which is associated with the most common operations of APTs, Command and Control.
+Figure 7: Visualization of the distribution of top-10 tactics and techniques in AttackSeqBench.
+A.2 Dataset Evaluation Criteria
+Inspired by the existing work (Fu et al., 2024), we utilize the following six dimensions as the evaluation criteria to evaluate the quality of our constructed Q&A dataset:
+  • Answerability.
+We check if there is direct evidence in the CTI outline that supports the correct answer, while clearly standing out as the best answer choice.
 Within this aspect, we also check if the correct answer can be inferred even if the associated summary to the correct answer’s tactic is removed from the CTI outline.
@@
   • Answer Consistency. We check if the question can be fully answered by the correct answer, without any contradictions and inconsistencies.
-We give detailed instructions to the two human evaluators by defining each score in the 5-point Likert scale for each aspect (refer to Table 7).
-The same Likert scale definitions are used during automatic evaluations in the prompt.
-The detailed results of human evaluation and automatic evaluation are shown in Table 6 and Table 7, respectively.
-While the automatic evaluation results are lower than human evaluation, the human evaluation shows that most QA pairs in the dataset satisfy the requirements of all aspects.
+To quantitatively evaluate its quality, we first design the 5-point Likert scale for each aspect (refer to Table LABEL:tab.annotation_instructions), where each score corresponds to a different level of the given aspect.
+Then we instruct three cybersecurity experts and LLM to provide the score of each aspect to achieve the human evaluation and the automatic evaluation, respectively.
+The detailed results are shown in Table 1.
+While the automatic evaluation results are lower than human evaluation, the human evaluation shows that most Q&A pairs in the dataset satisfy the requirements of all aspects.
 This suggests that automatic evaluation is still limited in knowledge-intensive domains such as in cybersecurity.
-Note that for AttackSeq-Procedure-No, we only assess the questions in the following four aspects: Answerability; Clarity, Consistency; Answer Consistency.
-A.3. Ethics Statement
-This study adheres to ethical research practices by focusing on publicly available CTI reports and ensuring that no proprietary or sensitive information is used in the benchmark construction.
-The automated dataset generation process is designed to maintain the integrity and accuracy of adversarial behavior sequences without fabricating or misrepresenting cyber threats.
-Furthermore, human evaluation is conducted with careful consideration of annotator expertise and potential biases, ensuring fairness and reliability in assessment.
-Our work aims to contribute to the cybersecurity community by advancing LLM capabilities while upholding ethical standards in data usage, evaluation, and responsible AI development.
+Note that for the AttackSeq-Procedure-No, we evaluate questions only on four aspects—Answerability, Clarity, Consistency, and Answer Consistency—since it is derived from AttackSeq-Procedure-Yes through negation of temporal prepositions and replacement of procedures.
+A.3 Benchmark Tasks
+Inspired by existing LLM benchmarks in the general domain (Hendrycks et al., 2021a; Rein et al., 2023; Zhang et al., 2024b), we propose three tasks in the form of Multiple-Choice Questions and Yes-No Questions to evaluate the reasoning capabilities of LLMs in inferring TTPs present in attack sequences, where each task reflects a distinct aspect of adversarial behaviors.
+Table 5: Annotation instructions for each evaluation aspect.
+Aspects Instructions
+Answerability Score 1: The correct answer is not supported by the CTI outline.
+The information is either missing or contradicts the correct answer.
+Without the masked tactic paragraph, it is impossible to deduce the correct answer.
+Score 2: Some evidence in the CTI outline loosely supports the correct answer, but it does not clearly stand out as the best choice.
+Removing the masked tactic paragraph makes it highly difficult to deduce the answer, even when referring to the MITRE ATT&CK KB.
+Score 3: The correct answer has partial support in the CTI outline but is not explicitly stated.
+After removing the masked tactic paragraph, it is possible but challenging to infer the correct answer using the remaining information and MITRE ATT&CK KB.
+Score 4: The correct answer is well-supported by the CTI outline and is the most reasonable choice based on the provided information.
+If the masked tactic paragraph is removed, the answer remains largely deducible using remaining information, and MITRE ATT&CK KB.
+Score 5: The correct answer is directly supported by the CTI outline and is unambiguously the best choice.
+Even if the masked tactic paragraph is removed, the answer remains easily deducible based on the remaining CTI outline and MITRE ATT&CK KB.
+Clarity Score 1: The question is highly ambiguous, imprecise, or contains vague phrasing. It may directly state the correct answer, making inference unnecessary.
+Score 2: The question is somewhat unclear or contains minor ambiguities. It may hint too strongly at the correct answer, reducing the need for inference.
+Score 3: The question is fairly clear, but minor ambiguities exist. It does not directly state the correct answer, but slight rewording could improve precision.
+Score 4: The question is mostly clear and unambiguous. It requires inference and does not directly reveal the correct answer.
+Score 5: The question is precise, completely unambiguous, and free of vague phrasing. The correct answer is never directly mentioned, ensuring inference is required.
+Logical Score 1: The question does not align with the logical sequence of MITRE ATT&CK tactics in the CTI outline.
+Score 2: The question shows minimal alignment with the MITRE ATT&CK sequence. It may reference unrelated tactics.
+Score 3: The question has some logical alignment, but it may not reference preceding or subsequent tactics clearly.
+Score 4: The question follows the sequence of MITRE ATT&CK tactics and references preceding or subsequent TTPs in a logical manner.
+Score 5: The question perfectly aligns with the MITRE ATT&CK framework, referencing relevant TTPs in a way that naturally leads to the correct answer.
+Relevance Score 1: The question is completely unrelated to the CTI outline.
+Score 2: The question has only slight relevance to the CTI outline but is mostly off-topic.
+Score 3: The question is somewhat related to the CTI outline but could be refined to better fit the content.
+Score 4: The question is directly related to the CTI outline, with minor room for improvement.
+Score 5: The question fully aligns with the CTI outline and is highly relevant to the content.
+Consistency Score 1: The question contradicts the TTP description or is entirely misaligned with the provided details.
+Consistency Score 2: The question loosely aligns with the TTP description but has inconsistencies or inaccuracies.
+Score 3: The question mostly aligns with the TTP description but contains minor inconsistencies.
+Score 4: The question is highly consistent with the TTP description, with only minor areas for improvement.
+Score 5: The question fully aligns with the TTP description, with no inconsistencies or contradictions.
+Answer Consistency Score 1: The correct answer does not fully resolve the question, leaving contradictions or gaps.
+Score 2: The correct answer provides some resolution, but contradictions or inconsistencies remain.
+Score 3: The correct answer is mostly consistent, but minor contradictions exist.
+Score 4: The correct answer fully resolves the question with minimal inconsistencies.
+Score 5: The correct answer completely and unambiguously answers the question, with no contradictions or inconsistencies.
+AttackSeq-Tactic.
+This task evaluates the LLMs’ capability to infer a tactic $t_{k}\in T$.
+Given a question $Q$ that corresponds to tactic $t_{k}$ and four shuffled candidate tactics $C_{Tac}=\{c_{r}:r\in[1,4]\}$, the LLM will be instructed to select the correct tactic $c_{l}\in C_{Tac}$.
+AttackSeq-Technique.
+This task assesses the LLMs’ capability to infer a technique $e_{j,k}\in E(t_{k})$.
+Given a question $Q$ that corresponds to $e_{j,k}$ and four shuffled candidate techniques $C_{Tec}=\{c_{r}:r\in[1,4]\}$, the LLM will be instructed to select the correct technique $c_{l}\in C_{Tec}$.
+AttackSeq-Procedure.
+This task challenges the LLMs’ capability to determine the likelihood of procedures $p_{m,j,k}\in P(e_{j,k})$ in an attack sequence.
+Given a question $Q$ and two candidate choices $C_{Pro}=\{yes,no\}$, the LLM will be instructed to determine if the procedure $p_{m,j,k}$ is likely to occur in the given attack sequence $S$.
+We further divide AttackSeq-Procedure into two sub-tasks, namely AttackSeq-Procedure-Yes and AttackSeq-Procedure-No, based on the ground truth of the boolean question.
+This explores the LLMs’ ability in determining misleading procedures that are unlikely to occur in an attack sequence.
+A.4 Baselines
+To demonstrate the effectiveness and robustness of our proposed AttackSeqBench, we evaluate seven large language models, five large reasoning models and four post-training strategies across three tasks involving different levels of data and three benchmark settings with varying context completeness.
+We leverage vLLM (Kwon et al., 2023) to run all the open-source LLMs locally with two Nvidia H100 GPUs.
+For the colsed-source LLMs, we utilize OpenAI’s Batch API  [^4] https://platform.openai.com/docs/guides/batch to conduct inference in batches.
+In our experiments, we set the following sampling parameters while keeping the default value for the remaining parameters: temperature to 0, maximum output tokens to 2048, and top_p to 1.
+Below is the details of the utilized LLMs, LRMs and post-training strategies in our experiments:
+Large Language Models:
+  • LLaMa-3.1-8B (Grattafiori et al., 2024) is an instruction-tuned LLM from Meta, balancing performance and efficiency for textual understanding tasks.
+  • ChatGLM-4-9B (GLM et al., 2024) is pretrained on ten trillions of tokens and further achieve the high-quality alignment through supervised fine-tuning and human feedback learning.
+  • Qwen2.5-3B, Qwen2.5-14B, Qwen2.5-32B and Qwen2.5-72B (QwenTeam, 2024) represents Qwen-2.5 series LLMs with different parameter scales, demonstrating strong instruction-following and long-text generation capabilities.
+  • LLaMa-3.3-70B (Grattafiori et al., 2024) is an auto-regressive language model which is instruction-tuned in 70B with SFT and reinforcement learning with human feedback (RLHF).
+  • GPT-4o (OpenAI, 2024a) is one of the most advanced closed-source LLMs, which is a multi-lingual and multi-modal language model developed and functions well in real-time processing.
+Large Reasoning Models:
+  • DeepSeek-R1-Distill-Llama-8B (R1 (Llama-8B)), DeepSeek-R1-Distill-Qwen-14B (R1 (Qwen-14B)) and DeepSeek-R1-Distill-Qwen-32B (R1 (Qwen-32B)) (DeepSeek-AI, 2025) are fine-tuned from the Llama-3.1-8B, Qwen2.5-14B and Qwen2.5-32B with 800k samples curated with DeepSeek-R1, aiming to equip smaller models with reasoning capabilities like DeepSeek-R1.
+  • QWQ-32B-Preview (QWQ-32B) (Team, 2024) is a preview release which gives LLM time to ponder, to question, and to reflect, enabling the deeper insight into complex problems.
+  • GPT-o3-mini (OpenAI, 2025) is designed with a focus on enhancing LLMs’ reasoning capabilities.
+It leverages the Chain of Thought (CoT) to break down complex problems into several simpler steps to achieve this objective.
+Post-Training Strategies:
+  • Supervised Fine-tuning (SFT) (Zhang et al., 2023) is a critical process for adapting pre-trained LLMs to specific tasks by training them on a task-specific dataset with labeled examples.
+  • Reasoning Distillation (RD) (Huang et al., 2024) RD is a widely adopted approach for enhancing LLM reasoning, which collects reasoning samples with self-reflection from existing LRMs and distills them to guide LLMs in acquiring long-thought capabilities.
+  • Reinforcement Learning from Internal Feedback (RLIF) (Zhao et al., 2025) replaces the external rewards in Group Relative Policy Optimization (GRPO) with LLMs’ self-certainty, enabling unsupervised learning from intrinsic signals without relying on external rewards.
+  • Reinforcement Learning with Verifiable Rewards (RLVR)  (DeepSeek-AI, 2025) leverages rule-based verification functions to provide reward signals for tasks with clear correctness criteria, enabling the optimization of LLMs while avoiding the complexities and potential pitfalls of reward models within RLHF.
+A.5 Post-training Corpus Construction
+Considering the post-training strategies, we construct two diverse datasets for SFT and RD, RLIF and RLVR respectively.
+For the former, we utilize a subset of the Primus-Instruct dataset (Yu et al., 2025).
+Primus-Instruct is a cybersecurity corpus collected for instruction-tuning, containing diverse task types such as alert explanation, suspicious command analysis, security event query generation, retrieved security document QA, Terraform security mis-configuration repair, and general multi-turn instruction following.
+To mitigate the inherent bias from linguistic inconsistencies, we filter out non-English samples via the FastText language identification library (Joulin et al., 2016) and manually verify the results, yielding a subset of 710 samples for SFT.
+Regarding the latter, we use Primus-Reasoning, a cybersecurity reasoning distillation corpus constructed with DeepSeek-R1 (DeepSeek-AI, 2025) and GPT-o1-preview (OpenAI, 2024a).
+This dataset includes, but is not limited to, tasks such as Common Weakness Enumeration (CWE) mapping, Common Vulnerabilities and Exposures (CVE) analysis, and multiple-choice questions on general cybersecurity knowledge.
+Following (Zhang et al., 2025b), we leverage transitional words (i.e., “but”, “however”, “wait”, etc.) as the proxy for inferability, and retain only the 3,890 samples containing at least ten such words when constructing the corpus for RD, RLIF and RLVR.
+A.6 Implementation Details
+To examine the performance of LLMs on our AttackSeqBench after embedding cybersecurity knowledge, and considering GPU constraints, we evaluate existing post-training strategies on Qwen-2.5-3B (QwenTeam, 2024) and LLaMA-3.1-8B (Grattafiori et al., 2024) under both full-parameter fine-tuning and parameter-efficient fine-tuning paradigms across all benchmark tasks and settings.
+Retrieval Augmented Generation (RAG): We first crawl the description and the example procedures of each technique in the Enterprise ATT&CK Matrix v17  [^5] https://attack.mitre.org/versions/v17/matrices/enterprise/.
+Then, we split the textual data into text chunks and embed the chunks in a vector store (i.e., Chroma DB  [^6] https://www.trychroma.com/), where each chunk’s metadata contains the associated ATT&CK tactic and technique.
+We utilize a hybrid retriever with a re-ranker, by combining combine Okapi BM25 (Robertson & Zaragoza, 2009) and a dense retriever based on the more advanced text-embedding-3-large from OpenAI (OpenAI, 2024b).
+We set the chunk size to 512 and retrieved chunks to 3, and utilize LlamaIndex (Liu, 2022) to implement the retriever.
+Additionally, we also implement BGE-EN-ICL (Li et al., 2024) and ATT&CK-BERT (Abdeen et al., 2023a) within RAG to evaluate their effectiveness in attack sequence analysis.
+Supervised Fine-tuning (SFT): We fine-tune the backbone LLM on the first dataset within Appendix A.5 using the LLaMA-Factory (Zheng et al., 2024) framework.
+Specifically, we deliberately restricted SFT to one epoch with the learning rate of $3\times 10^{-6}$, leveraging DeepSpeed ZeRO Stage-3 with CPU offload for memory efficiency.
+In our preliminary experiments, extending training process to multiple epochs led to noticeable degradation in the LLMs’ general capabilities outside the cybersecurity domain.
+This effect can be attributed to “catastrophic forgetting”, where continued exposure to a narrow corpus may overwrite its previously acquired broad knowledge.
+Thus, a single epoch struck a balance between adapting the LLM to the cybersecurity tasks while preserving its pre-trained general-purpose performance.
+Reasoning Distillation (RD) refers to fine-tune LLM on a reasoning dataset distilled from the advanced LRMs (i.e., DeepSeek-R1  and GPT-o1-preview ), which enables the smaller LLM to inherit the reasoning behaviors of the above LRMs.
+For RD, we fine-tune our backbone LLM on the latter dataset within Appendix A.5 with the same parameter settings in SFT.
+Reinforcement Learning with Verifiable Rewards (RLVR) extends reinforcement learning by incorporating verifiable signals as rewards, such as correctness checks or logical consistency that can be programmatically validated.
+Specifically, we implement RLVR with Group Relative Policy Optimization (GRPO) (Shao et al., 2024) on the Volcano Engine Reinforcement Learning (verl) framework, where group-normalized rewards reduce variance and stabilize training.
+We conduct RLVR with the learning rate of $3\times 10^{-6}$ using Fully Sharded Data Parallel (FSDP).
+Reinforcement Learning with Internal Feedback (RLIF) (Zhao et al., 2025) enables LLMs to optimize policies using intrinsic signals without relying external supervision.
+In particular, it replaces the external rewards in GRPO with the self-certainty scores, which estimate the LLMs’ confidence from its outputs, to enable fully unsupervised learning while maintaining the stability benefits of GRPO.
+Here, we conduct RLIF on the same latter dataset within Appendix A.5 using the verl framework, adopting the same training hyperparameters and FSDP set-up as in RLVR.
+Figure 8: Comparison between existing CTI-related benchmarks and our AttackSeqBench based on a real-world CTI report (Koehl & Hannon, 2020).
+Our benchmark emphasizes on the sequences of adversary behaviors described in CTI reports.
+A.7 Related Benchmarks Comparison
+To demonstrate the uniqueness and novelty of our work, we illustrate the key differences between existing CTI-related benchmarks and our AttackSeqBench which significantly highlights attack sequence analyzing in Figure 1.
+Specifically, existing CTI-related benchmarks primarily focus on evaluating LLMs on three aspects: (1) CTI Classification, classifying malicious actions to known adversary behaviors (Alam et al., 2023); (2) CTI Extraction, extracting entities relevant to threat intelligence from the unstructured text (Bhusal et al., 2024); (3) CTI Inference, inferring the attributions of cyber attacks described in the real-world CTI reports (Alam et al., 2024).
+While these benchmarks preliminarily investigate the information extraction capabilities of LLMs within the CTI-related secnario, their ability to understand the sequential patterns of adversarial behavior remains largely unexplored.
+Besides, although KB (i.e., MITRE ATT&CK® (Strom et al., 2018)) document real-world adversary behaviors through the pre-defined attack patterns, analyzing the patterns individually is insufficient to fully capture the progression of cyber attacks as listed in CTI reports.
+The sophisticated and stealthy nature of APTs requires a comprehensive understanding of how adversaries transition between the different attack phases, which are orchestrated as an attack sequence.
+This raises the need to consider the sequential characteristics of a cyber attack within the given CTI report.
+A.8 The Use of Large Language Models
+We declare that LLMs were employed to assist with the refinement of this manuscript, specifically for grammar checking and language polishing.
+Additionally, LLMs were used in a limited capacity for minor debugging and syntactic correction of code snippets.
+Beyond these auxiliary roles, given that the primary purpose of this work is to explore the capability of diverse types of LLMs in understanding adversarial attack sequences, we also utilize a range of open-source and closed-source LLMs and LRMs during dataset construction, dataset refinement, and performance evaluation.
+All such uses are documented in the main paper and the appendix and were carefully controlled to ensure transparency and reproducibility.
+A.9 Limitations and Future works
+Limitations: While our work serves as a pioneering study into the LLMs’ reasoning capabilities in attack sequence analysis, several limitations should be acknowledged.
+Firstly, our study focuses on correctness of models’ responses through Multi-Choice Questions and Yes-No Questions, which may not fully capture the reasoning abilities of LLMs necessary for comprehensive evaluation.
+Secondly, although we have conducted extensive experiments with seven LLMs, five LRMs, and four post-training strategies across three benchmark tasks (AttackSeq-Tactic, AttackSeq-Technique, and AttackSeq-Procedure) and three benchmark settings (Zero-Shot setting, Context setting, and RAG-empowered setting), fully demonstrating the Reasoning Scalability and Domain-Specific Epistemic Expandability of our AttackSeqBench, the implementations of RAG and post-training strategies remains relatively basic and leave room for future refinement.
+Thirdly, our AttackSeqBench currently leverages 408 rigorously filtered CTI reports to extract attack sequences and generate Q&A pairs.
+Although this number substantially exceeds prior CTI-related studies (i.e., 12 in AttacKG+ (Zhang et al., 2025c), 12 in MM-AttacKG (Zhang et al., 2025d), and at most 71 in Attack Flow  [^7] https://center-for-threat-informed-defense.github.io/attack-flow/), the proposed dataset construction pipeline is flexible and can be readily extended to unseen CTI reports.
+This not only demonstrates the Extensibility of our AttackSeqBench but also highlights an important direction for continuously refining this benchmark in future work.
+Nevertheless, while it is important to be aware of these limitations, our AttackSeqBench serves are a valuable benchmark to systematically explore LLMs’ reasoning abilities across the tactical, technical and, procedural dimensions of adversarial behaviors.
+Future works: Building on the limitations, our future research on AttackSeqBench will proceed along three directions.
+Regarding evaluation, we plan to expand our evaluation methods from the simple Multiple-Choice Question tasks and Yes-No Question tasks to the more complex reasoning and completion tasks, thereby providing a more comprehensive assessment of model capabilities in CTI report understanding.
+In terms of methodology, we will build on AttackSeqBench to explore more fine-grained RAG approaches and advanced post-training strategies that account for the knowledge-extensive and high-stakes nature of CTI reports understanding, aiming to fully leverage model potential in complex cyber-attack scenarios.
+At the data level, we will continue to expand and dynamically update the CTI corpus to ensure our AttackSeqBench remains evolvable over time, thereby supporting the steady advancement of domain-specific foundation models for cybersecurity.
 Appendix B Experiments
-B.1. Baseline LLMs
-To demonstrate the robustness of our proposed datasets, we evaluate eight fast-thinking LLMs and four slow-thinking reasoning LLMs on our benchmark tasks.
-We leverage vLLM (Kwon et al., 2023) to run all open-source LLMs locally, utilizing two Nvidia H100 GPUs.
-For the proprietary LLMs, we utilize OpenAI’s Batch API  [^4] https://platform.openai.com/docs/guides/batch to reduce inference cost and perform inference in batches.
-In our experiments, we set the following sampling parameters while keeping the default value for the remaining parameters: temperature to 0, maximum output tokens to 2048, and top_p to 1.
-Below is the list of LLMs used in our experiments:
-Fast-thinking LLMs:
-  • Mistral-7B-Instruct-v0.3 Mistral-7B-Instruct-v0.3 (Jiang et al., 2023) is an instruct fine-tuned version of the Mistral-7B-v0.3 which exhibits extended vocabulary and supports the v3 tokenizer.
-  • Qwen2.5-7B-Instruct Qwen2.5-7B-Instruct (Team, 2024a) is one of the advanced Qwen series LLMs. It exhibits better instruction following and long text generating capability.
-  • LLaMa-3.1-8B-Instruct LLaMa-3.1-8B-Instruct (Grattafiori et al., 2024) is an instruction-tuned LLM from Meta, balancing performance and efficiency for textual understanding tasks.
-  • LLaMa-3.2-3B-Instruct LLaMa-3.2-3B-Instruct (Grattafiori et al., 2024) is a lightweight 3B version of LLaMA-3.2, prioritizing the rapid response and the deployment on low-resource devices.
-  • Qwen2.5-72B-Instruct Qwen2.5-72B-Instruct (Team, 2024a) is one of the latest LLMs from Qwen and serves as the biggest Qwen-2.5 series LLM with 72B parameters.
-  • LLaMa-3.3-70B-Instruct LLaMa-3.3-70B-Instruct (Grattafiori et al., 2024) is an auto-regressive language model which is instruction-tuned in 70B with supervised fine-tuning (SFT) and reinforcement learning with human feedback (RLHF).
-  • GPT-4o GPT-4o (OpenAI, 2024b) is one of the most advanced closed-source LLMs, which is a multi-lingual and multi-modal language model developed and functions well in real-time processing  [^5] To assess its superiority, we leverage the most recent GPT-4o-2024-11-20 and GPT-4o-mini-2024-07-18 (at the time point when we did this work) as the benchmark evaluator, respectively..
-  • GPT-4o-mini gpt-4o-mini (OpenAI, 2024a) is a compact variant of the GPT-4o (OpenAI, 2024b), which is optimized for faster inference and cost efficiency 5.
-Slow-thinking reasoning LLMs:
-  • DeepSeek-R1-Distill-Llama-8B DeepSeek-R1-Distill-Llama-8B (DeepSeek-AI, 2025) is fine-tuned from the open-source Llama-3.1-8B with 800k samples curated with DeepSeek-R1, aiming to equip smaller models with reasoning capabilities like DeepSeek-R1.
-  • DeepSeek-R1-Distill-Qwen-32B DeepSeek-R1-Distill-Qwen-32B (DeepSeek-AI, 2025) denotes the distilled version from the well-trained Qwen2.5-32B without any Reinforce Learning stage.
-  • QWQ-32B-Preview QWQ-32B-Preview (Team, 2024b) is a preview release which gives LLM time to ponder, to question, and to reflect, enabling the deeper insight into complex problems.
-  • GPT-o3-mini GPT-o3-mini (OpenAI, 2025) is designed with a focus on enhancing LLMs’ reasoning capabilities.
-It leverages the Chain of Thought (CoT) to break down complex problems into several simpler steps to achieve this objective.
-The retrievers in our experiments are built using LlamaIndex (Liu, 2022). We set the number of retrieved chunks $u$ and its size $l$ to be $3$ and $512$ respectively.
-B.2. TTP Temporal Position Analysis
-We conduct a detailed breakdown of the performance in four indicative LLMs based on each ATT&CK tactic associated to the questions in AttackSeqBench.
-Based on Table 10, the mean and standard deviation (SD) for each tactic and between each benchmark task is shown.
-We observe that the overall highest standard deviation occurs within the AttackSeq-Tactic task for all models, indicating that it is challenging for LLMs to correctly answer questions in AttackSeq-Tactic.
-Table 10. Results of performance of four representative models on all the 14 tactics, and the associative statistics of mean and standard deviation (SD).
+B.1 TTP Temporal Position Analysis
+To better understand the LLMs’ capabilities in attack sequence analysis, we conduct fine-grained analysis based on each stage within the attack sequences of MITRE ATT&CK®.
+We illustrate the performance of two LLMs and two LRMs across all benchmark tasks in the Regular setting in Figure 9 and show the corresponding values, the mean and standard deviation (SD) of these LLMs and LRMs on each tactic and benchmark task in Table 6.
+We identify four overachieving attack phases to categorize the ATT&CK tactics in attack sequences: (1) Initial Intrusion Phase; (2) Exploitation Phase; (3) Stealth Expansion Phase; (4) Objective Orchestration Phase.
+It is worth noting that our categorization follows Tactics, as each Technique and Procedure in MITRE ATT&CK® is uniquely mapped to a specific Tactic within a given attack sequence.
+Figure 9: The performance comparison between two LLMs and two LRMs on each tactic in each benchmark task.
+Table 6: Results of performance of two LLMs and two LRMs on all the 14 tactics, and the corresponding statistics of mean and standard deviation.
 Tactic
@@
 SD 0.0260 0.0796 0.1139 0.0462 0.0368 0.0684 0.0355 0.0848 0.0473 0.0467 0.0714 0.0467 0.0530 0.0829 - -
-B.3. Error Analysis
-To better understand the limitations of LLMs in analyzing attack sequences, we examine a specific cases where GPT-4o successfully and fails to answer the given question correctly (refer to Table 11).
-In the first example, the model incorrectly associates the technique in answer choice A, i.e., T1574-Hijack Execution Flow, with the Execution tactic.
-While this technique involves the execution of malicious payloads based on the ATT&CK KB, it is associated to only the following tactics, Persistence, Privilege Escalation and Defense Evasion.
-This incorrect interpretation suggests that the model struggles in distinguishing the inherent ambiguity found in TTP descriptions, which affects their ability to analyze attack sequences.
-Table 11. Error analysis of GPT-4o in AttackSeq-Technique. The LLM incorrectly associates technique T1574 with Execution tactic
-Question & Answer Choices LLM Response
+We summarize our findings into three key points: (1) Comparing the performance of three benchmark tasks, the overall highest standard deviation occurs within the AttackSeq-Tactic task for all LLMs and LRMs, suggesting that the LLMs’ and LRMs’ performance has the largest variance in this taskand further highlighting the challenge of inferring tactics within attack sequences; (2) In contrast, the standard deviation for AttackSeq-Procedure is the lowest for all LLMs, suggesting more consistency in the LLMs’ and LRMs’ performance among the different attack sequence stages; (3) GPT-4o consistently outperforms other LLMs and LRMs across most benchmark tasks, particularly in the Exploitation Phase and the Stealth Expansion Phase.
+However, the remaining LLMs achieve comparable performance in the remaining two phases, especially in tactics at the start and end of the attack sequences (e.g., Reconnaissance and Impact).
+This suggests that TTPs within rarely observed tactics are more challenging for the LLM and LRMs to interpret.
+B.2 Case Study
+To better understand the superiority and limitation of the advanced LLMs in analyzing attack sequences, we run GPT-4o multiple times to collect both successful and failed responses to the same questions under the identical experimental setting, which are presented in Table 7.
+For the Wrong case, GPT-4o incorrectly associates the technique of choice A, i.e., T1574-Hijack Execution Flow, with the Execution tactic.
+Although this technique indeed involves the execution of malicious payloads based on the ATT&CK knowledge base, it is only associated to the following three tactics: Persistence, Privilege Escalation, and Defense Evasion.
+This incorrect interpretation suggests that GPT-4o struggles in distinguishing the inherent ambiguity found in TTP descriptions, thereby affecting their ability to analyze attack sequences.
+Regarding the Right one, GPT-4o correctly identifies T1129-Shared Modules as the most plausible technique, which belongs to the Execution tactic  [^8] https://attack.mitre.org/techniques/T1129/ and serves as the executable files that are loaded into processes to provide access to execute malicious payloads.
+By selecting this option, GPT-4o demonstrates its ability to reason over the attack sequence: after WMI discovery detects the presence of antivirus, shared modules would facilitate the execution of “FinalBot.exe” to bypass basic defenses.
+This correct interpretation is beneficial to effectively link the ambiguous textual cues with the appropriate tactic/technique entities, thereby improving its reliability in analyzing attack sequences.
+Table 7: Error analysis of GPT-4o in AttackSeq-Technique, where GPT-4o incorrectly associates technique T1574-Hijack Execution Flow with the Execution tactic.
+Q & A Cases LLM Response
 Question: Before establishing persistence through creating a Run key in the registry, which ATT&CK technique most likely facilitated the execution of the backdoor FinalBot.exe after the WMI discovery process identified the presence of antivirus software?
-Answer Choices: A) T1574-Hijack Execution Flow B) T1003.001-LSASS Memory C) T1059.008-Network Device CLI D) T1129-Shared Modules (Correct Answer) **A: T1574-Hijack Execution Flow** This technique involves hijacking the execution flow of a legitimate process to execute malicious code.
+Options: A) T1574-Hijack Execution Flow B) T1003.001-LSASS Memory C) T1059.008-Network Device CLI D) T1129-Shared Modules (Correct Answer) Wrong A: T1574-Hijack Execution Flow This technique involves hijacking the execution flow of a legitimate process to execute malicious code.
 DLL sideloading is a common method for this technique.
@@
 This likely facilitated the execution of the backdoor ‘FinalBot.exe‘ after the WMI discovery process.
-**B: T1003.001-LSASS Memory**: This technique involves dumping credentials from the LSASS process memory.
+B: T1003.001-LSASS Memory: This technique involves dumping credentials from the LSASS process memory.
 There is no mention in the CTI report of credential dumping or any activity related to LSASS.
 This choice is irrelevant.
-**C: T1059.008-Network Device CLI**: This technique involves using a command-line interface (CLI) to interact with network devices.
+C: T1059.008-Network Device CLI: This technique involves using a command-line interface (CLI) to interact with network devices.
 The CTI report does not mention any interaction with network devices or the use of a CLI.
 This choice is irrelevant.
-**D: T1129-Shared Modules** This technique involves executing malicious code by loading shared modules.
+D: T1129-Shared Modules: This technique involves executing malicious code by loading shared modules.
 While DLL sideloading involves shared modules, the specific ATT&CK technique for DLL sideloading is T1574-Hijack Execution Flow.
 This choice is less precise than option A.
-The most plausible sequence of TTPs involves the use of **T1574-Hijack Execution Flow** (via DLL sideloading) to execute the backdoor ‘FinalBot.exe‘ after the WMI discovery process and before establishing persistence (Wrong Conclusion).
-Appendix C Prompts Strategy
-C.1. Question Generation Prompts
-We utilize few-shot prompting for each of our benchmark tasks, the prompts for AttackSeq-Tactic, AttackSeq-Technique, AttackSeq-Procedure-Yes, AttackSeq-Procedure-No is shown in Tables 12, 13, 14, 15 respectively.
-Table 12.
-Few-shot prompt for question generation for AttackSeq-Tactic.
-We insturct the LLM to generate a question based on the given CTI outline, where the answer to the question is the given ATT&CK TTP.
+Thus, the most plausible sequence of TTPs involves the use of T1574-Hijack Execution Flow (via DLL sideloading) to execute the backdoor ‘FinalBot.exe‘ after the WMI discovery process and before establishing persistence.
+(Wrong Conclusion)
+Right A: T1574-Hijack Execution Flow: This technique involves exploiting or redirecting the flow of execution to execute untrusted code.
+While this could potentially execute malware, it doesn’t directly relate to the sequence of WMI discovery leading to the execution of ‘FinalBot.exe‘.
+B: T1003.001-LSASS Memory: This technique involves accessing credentials in memory by dumping the LSASS process.
+This is unrelated to executing a backdoor like ‘FinalBot.exe‘ after WMI discovery.
+C: T1059.008-Network Device CLI: This technique involves command-line interfaces for network devices, which is unrelated to executing programs on a system post-discovery of antivirus software.
+D: T1129-Shared Modules: This technique involves executing malicious payloads using shared libraries or modules.
+If ‘FinalBot.exe‘ were executed via shared modules, this process could occur after determining system defenses, making it relevant to the scenario.
+Answer: D: T1129-Shared Modules.
+This technique likely facilitated the execution of ‘FinalBot.exe‘ after WMI discovery detected antivirus presence, allowing for a subtle execution method bypassing basic defenses.
+(Correct Conclusion)
+Table 8: Performance comparison between three embedding models (abbreviate as Emb. M.).
+#Emb. M. BGE OPENAI ATT&CK-BERT
+Tasks Tactic Technique Procedure Tactic Technique Procedure Tactic Technique Procedure
+Llama-3.1-8B 0.4751 0.5974 0.5243 0.4838 0.6103 0.5435 0.4820 0.5980 0.5334
+GPT-4o 0.5522 0.6860 0.6319 0.5616 0.6578 0.6482 0.5687 0.7016 0.6216
+R1 (Llama-8B) 0.4905 0.5740 0.5226 0.4932 0.5696 0.5150 0.4651 0.5804 0.5104
+GPT o3-mini 0.5115 0.5853 0.6474 0.5192 0.5827 0.6474 0.5245 0.5874 0.6414
+B.3 Impact of Embedding Models within RAG-empowered Setting
+The semantics within CTI reports contain a large volume of domain-specific technical terminologies, where the accuracy of retrievers in identifying the most relevant tactics, techniques and procedures critically influences LLMs’ performance in RAG-empowered settings.
+To further examine this issue and mitigate the potential knowledge bias introduced by BGE-EN-ICL, we incorporate two additional embedding models into our RAG-empowered setting, namely OpenAI’s text-embedding-3-large (OpenAI, 2024b) and a domain-adapted ATT&CK BERT  (Abdeen et al., 2023b) fine-tuned on the specific cybersecurity data.
+We conduct performance comparisons between two representative LLMs and two LRMs based on these above embedding models in Table 8.
+We observe that the three embedding models exhibit comparable performance across different benchmark tasks and settings, with only marginal differences.
+Among them, BGE-EN-ICL proves to be the most cost-efficient, generalizable, and effective choice, and thus we primarily report model performance based on this embedding throughout the paper.
+Notably, although ATT&CK-BERT contains far fewer parameters than BGE-EN-ICL (110M vs.
+7B), it achieves comparable performance, underscoring the importance of injecting domain-specific security knowledge into LLMs and pointing to a promising direction for future work.
+Appendix C Prompt Templates
+C.1 Question Generation Prompt Templates
+We utilize few-shot prompt templates for question generation of each of our benchmark tasks (i.e., AttackSeq-Tactic, AttackSeq-Technique, AttackSeq-Procedure-Yes, AttackSeq-Procedure-No), the corresponding prompts are shown in Box C.1, Box C.1, Box C.1, Box C.1 respectively.
+Box 1: Few-shot prompt for question generation for AttackSeq-Tactic.
 You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework.
-[Inputs]: You will receive two parts: 1.
-A **CTI Report** that describe a cyber attack ordered by MITRE ATT&CK tactics.
+[Inputs]:
+You will receive two parts:
+1.
+A CTI Report that describe a cyber attack ordered by MITRE ATT&CK tactics.
 Note that additional information labeled as “Others” provides context about the threat actor but is secondary.
 2.
-A **MITRE ATT&CK Tactic** present in the CTI report.
+A MITRE ATT&CK Tactic present in the CTI report.
 [End of Inputs]
@@
 The question should focus on inferring the given tactic by using the attack sequence based on the remaining tactics in the CTI report.
-[Steps]: Please follow these steps: 1.
+[Steps]:
+Please follow these steps:
+1.
 Analyze the CTI report: - Read the report carefully.
@@
 - Please refer to the example questions below for guidance.
-[Examples]: Example Questions: - Question: After gaining initial access through compromised VPN accounts, which ATT&CK tactic most likely occurred before Ke3chang achieved persistence by adding a Run key?
+[Examples]:
+Example Questions:
+- Question: After gaining initial access through compromised VPN accounts, which ATT&CK tactic most likely occurred before Ke3chang achieved persistence by adding a Run key?
 Answer: Execution - Question: Which ATT&CK tactic most likely occurred before Axiom gained initial access to the victim’s network using SQL injection?
@@
 Answer: Execution [End of Examples]
-3. Provide the Question-Answer Pair: - Please follow the output format: ”Question: ¡insert question here¿ Answer: ¡insert answer here¿”.
-[End of Steps] Following the steps above, please generate a question based on the CTI report and ATT&CK tactic given below.
-Table 13. Few-shot prompt for question generation for AttackSeq-Technique.
+3. Provide the Question-Answer Pair:
+- Please follow the output format: “Question: <insert question here> Answer: <insert answer here> [End of Steps] Following the steps above, please generate a question based on the CTI report and ATT&CK tactic given below.
+Box 2: Few-shot prompt for question generation for AttackSeq-Technique.
 You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework.
 [Inputs]: You will receive three parts: 1.
-A **CTI Report** that describe a cyber attack ordered by MITRE ATT&CK tactics.
+A CTI Report that describe a cyber attack ordered by MITRE ATT&CK tactics.
 Note that additional information labeled as “Others” provides context about the threat actor but is secondary.
 2.
-A **MITRE ATT&CK Tactic** present in the CTI report.
+A MITRE ATT&CK Tactic present in the CTI report.
 3.
-A **MITRE ATT&CK Technique** present in the CTI report.
+A MITRE ATT&CK Technique present in the CTI report.
 [End of Inputs]
@@
 Answer: T1569.002-Service Execution [End of Examples]
-3. Provide the Question-Answer Pair: - Please follow the output format: ”Question: ¡insert question here¿ Answer: ¡insert answer here¿”.
+3. Provide the Question-Answer Pair:
+- Please follow the output format: “Question: <insert question here> Answer: <insert answer here>”.
 [End of Steps] Following the steps above, please generate a question based on the CTI report and ATT&CK tactic and technique given below.
-Table 14. Few-shot prompt for question generation for AttackSeq-Procedure-Yes.
+Box 3: Few-shot prompt for question generation for AttackSeq-Procedure-Yes.
 You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework.
 [Inputs]: You will receive four parts: 1.
-A **CTI Report** that describe a cyber attack ordered by MITRE ATT&CK tactics.
+A CTI Report that describe a cyber attack ordered by MITRE ATT&CK tactics.
 Note that additional information labeled as “Others” provides context about the threat actor but is secondary.
 2.
-A **MITRE ATT&CK Tactic** present in the CTI report.
+A MITRE ATT&CK Tactic present in the CTI report.
 3.
-A **MITRE ATT&CK Technique** present in the CTI report.
+A MITRE ATT&CK Technique present in the CTI report.
 4.
 4.
-A list of **Procedures** present in the CTI report, where each procedure is represented as a (Subject, Relation, Object) triplet.
+A list of Procedures present in the CTI report, where each procedure is represented as a (Subject, Relation, Object) triplet.
 [End of Inputs]
@@
 Answer: Yes [End of Examples]
-3. Provide the Question-Answer Pair: - Please follow the output format: ”Question: ¡insert question here¿ Answer: ¡insert answer here¿”.
+3. Provide the Question-Answer Pair:
+- Please follow the output format: “Question: <insert question here> Answer: <insert answer here>”.
 [End of Steps] Following the steps above, please generate a question based on the CTI report and ATT&CK tactic, technique and procedures given below.
-Table 15. Few-shot prompt for question generation for AttackSeq-Procedure-No.
+Box 4: Few-shot prompt for question generation for AttackSeq-Procedure-No.
 You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework.
-[Inputs]: You will receive two parts: 1.
-A **Reference Question-Answer Pair** that focuses on the logical sequence of TTPs in a CTI report, note that the answer to the this question is always ”Yes”.
+[Inputs]:
+You will receive two parts:
+1.
+A Reference Question-Answer Pair that focuses on the logical sequence of TTPs in a CTI report, note that the answer to the this question is always ”Yes”.
 2.
-A **Reference MITRE TTP** that is NOT supported in the CTI report.
+A Reference MITRE TTP that is NOT supported in the CTI report.
 [End of Inputs]
@@
 Question 2 should replace the main procedures in the reference question with the provided reference MITRE TTP, such that the replaced main procedures are not found in the CTI report.
-[Steps]: Please follow these steps: 1.
+[Steps]:
+Please follow these steps:
+1.
 Analyze the Reference Question-Answer Pair: - Identify and outline the attack sequence in the order presented in the reference Question-Answer Pair.
@@
 - Please refer to the examples below for guidance.
-[Examples]: Example Questions: - Example 1: Reference Question: After gaining initial access through compromised VPN accounts, will the Ke3chang malware most likely run commands on the command-line interface before achieving persistence by adding a Run key?
+[Examples]:
+Example Questions:
+- Example 1: Reference Question: After gaining initial access through compromised VPN accounts, will the Ke3chang malware most likely run commands on the command-line interface before achieving persistence by adding a Run key?
 Reference Answer: Yes Reference TTP: Tactic: Initial Access, Technique: T1651-Cloud Administration Command, Example Procedures: AADInternals can execute commands on Azure virtual machines using the VM agent.
@@
 Answer: No [End of Examples]
-3.
-Provide the Question-Answer Pairs: - Please follow the output format: ”Question 1: ¡insert question 1 here¿ Answer: ¡insert answer to question 1 here¿ Question 2: ¡insert question 2 here¿ Answer: ¡insert answer to question 2 here¿”.
-[End of Steps] Following the steps above, please generate two questions based on the Reference Question-Answer Pair and Reference MITRE TTP given below.
+3. Provide the Question-Answer Pairs:
+- Please follow the output format: “Question 1: <insert question 1 here> Answer: <insert answer to question 1 here>.” “Question 2: <insert question 2 here> Answer: <insert answer to question 2 here>.” [End of Steps] Following the steps above, please generate two questions based on the Reference Question-Answer Pair and Reference MITRE TTP given below.
 Please only provide the final output of the two questions.
-C.2. Dataset Refinement Prompts
-The prompt templates to filter based on the Answerability criteria is in Table 16, while the feedback and refinement prompts are in Table 17 and Table 18 respectively.
-Table 16. Prompt template for verifying Answerability during Self-Refinement.
+C.2 Dataset Refinement Prompt Templates
+The prompt templates to filter based on the Answerability criteria is in Box C.2, while the feedback and refinement prompts are in Box C.2 and Box C.2 respectively.
+Box 5: Prompt template for verifying Answerability during Self-Refinement.
 You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework.
 [Inputs]: You will receive three parts: 1.
-**CTI Outline**: A structured account of a cyber attack, ordered by MITRE ATT&CK tactics.
+CTI Outline: A structured account of a cyber attack, ordered by MITRE ATT&CK tactics.
 Additional context under ”Others” provides background on the threat actor but is secondary.
 2.
-**TTP Description**: A reference description of the correct answer corresponding to the question.
+TTP Description: A reference description of the correct answer corresponding to the question.
 3.
-**Question with Answer Choices**: A question aimed at inferring a TTP from the attack sequence described in the CTI report, along with one correct answer and distractors among the answer choices.
+Question with Answer Choices: A question aimed at inferring a TTP from the attack sequence described in the CTI report, along with one correct answer and distractors among the answer choices.
 [End of Inputs] Your task is to evaluate the answerability of the given question using the provided information in the CTI Outline and TTP description.
@@
 3.
-Output evaluation result: - Output one of the following: - ”A”: Indicates that the question is answerable.
-- ”B”: Indicates that the question is not answerable.
-- ”C”: Indicates that you do not know/cannot determine if the question is answerable.
+Output evaluation result: - Output one of the following: - “A”: Indicates that the question is answerable.
+- “B”: Indicates that the question is not answerable.
+- “C”: Indicates that you do not know/cannot determine if the question is answerable.
 - Please also include a short and concise explanation of your evaluation result.
-- Please follow the output format: ”Explanation: ¡insert explanation here¿ Evaluation Result: ¡insert letter here¿”.
-[End of Steps] Following the steps above, please evaluate the question using the CTI report and description below and only output the evaluation result.
-Table 17. Prompt template for feedback on question quality based on evaluation criteria.
+- Please follow the output format: “Explanation: <insert explanation here> Evaluation Result: <insert letter here>.” [End of Steps]
+Following the steps above, please evaluate the question using the CTI report and description below and only output the evaluation result.
+Box 6: Prompt template for assessing question quality based on the evaluation criteria.
 You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the Tactics, Techniques and Procedures (TTPs) in MITRE ATT&CK framework.
 [Inputs]: You will receive three parts: 1.
-**CTI Outline**: A structured account of a cyber attack, ordered by MITRE ATT&CK tactics.
+CTI Outline: A structured account of a cyber attack, ordered by MITRE ATT&CK tactics.
 Additional context under ”Others” provides background on the threat actor but is secondary.
 2.
-**TTP Description**: A reference description of the correct answer to the question.
+TTP Description: A reference description of the correct answer to the question.
 3.
-**Question with Answer Choices**: A question aimed at inferring a TTP from the attack sequence described in the CTI outline, along with one correct answer and distractors among the answer choices.
+Question with Answer Choices: A question aimed at inferring a TTP from the attack sequence described in the CTI outline, along with one correct answer and distractors among the answer choices.
 [End of Inputs]
-[Evaluation Criteria]: Your task is to evaluate the QA pair and provide your feedback for each of the criteria defined below: 1.
-**Clarity**: Is the question precise, unambiguous, and free of vague phrasing?
+Your task is to evaluate the QA pair and provide your feedback for each of the criteria defined below: [Evaluation Criteria]:
+Please refer to the definition of each feedback criterion:
+1.
+Clarity: Is the question precise, unambiguous, and free of vague phrasing?
 Does it avoid directly mentioning the correct answer, ensuring the respondent must infer the correct answer rather than having it stated in the question?
 2.
-**Logical**: Does the question align with the logical sequence of MITRE ATT&CK tactics in the CTI outline?
+Logical: Does the question align with the logical sequence of MITRE ATT&CK tactics in the CTI outline?
 Does the question reference TTPs from the preceding or subsequent tactics in the CTI outline such that it logically leads to the correct answer?
 3.
-**Relevance**: Does the question directly relate to the CTI outline?
+Relevance: Does the question directly relate to the CTI outline?
 4.
-**Consistency**: Does the question align with the provided TTP Description?
+Consistency: Does the question align with the provided TTP Description?
 5.
-**Answer Consistency**: Can the question be fully answered using the correct answer, without any contradictions or inconsistencies?
+Answer Consistency: Can the question be fully answered using the correct answer, without any contradictions or inconsistencies?
 [End of Evaluation Criteria] [Steps]: Please follow these steps: 1.
@@
 4.
-Output Feedback Scores: - Please follow the output format: Feedback Scores: - Clarity: ¡Your feedback¿ (¡Score¿/5) - Logical: ¡Your feedback¿ (¡Score¿/5) - Relevance: ¡Your feedback¿ (¡Score¿/5) - Consistency: ¡Your feedback¿ (¡Score¿/5) - Answer Consistency: ¡Your feedback¿ (¡Score¿/5) Total Score: ¡Total Score¿/25 [End of Steps] Following the steps above, please evaluate the Question with Answer Choices below using the provided CTI report and TTP Description.
+Output Feedback Scores: - Please follow the output format: Feedback Scores: - Clarity: <Your feedback> (<Score>/5) - Logical: <Your feedback> (<Score>/5) - Relevance: <Your feedback> (<Score>/5) - Consistency: <Your feedback> (<Score>/5) - Answer Consistency: <Your feedback> (<Score>/5) Total Score: <Total Score>/25 [End of Steps] Following the steps above, please evaluate the Question with Answer Choices below using the provided CTI report and TTP Description.
 Please only output the Feedback Scores.
-Table 18. Prompt template for question refinement.
+Box 7: Prompt template for question refinement.
 You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework.
 [Inputs]: You will receive three parts: 1.
-**CTI Outline**: A structured account of a cyber attack, ordered by MITRE ATT&CK tactics.
+CTI Outline: A structured account of a cyber attack, ordered by MITRE ATT&CK tactics.
 Additional context under ”Others” provides background on the threat actor but is secondary.
 2.
-**Question with Answer Choices**: A question aimed at inferring a TTP from the attack sequence described in the CTI report, along with one correct answer and distractors among the answer choices.
+Question with Answer Choices: A question aimed at inferring a TTP from the attack sequence described in the CTI report, along with one correct answer and distractors among the answer choices.
 3.
-**Feedback Results**: A list of feedback scores and explanations for each desired criterion of the question defined below.
-[End of Inputs] [Evaluation Criteria]: Please refer to the definition of each feedback criterion: 1.
-**Clarity**: Is the question precise, unambiguous, and free of vague phrasing?
+Feedback Results: A list of feedback scores and explanations for each desired criterion of the question defined below.
+[End of Inputs] Your task is to iteratively refine the quality of the given question based on the feedback provided in the Feedback Results.
+[Evaluation Criteria]: Please refer to the definition of each feedback criterion: 1.
+Clarity: Is the question precise, unambiguous, and free of vague phrasing?
 Does it avoid directly mentioning the correct answer, ensuring the respondent must infer the correct answer rather than having it stated in the question?
 2.
-**Logical**: Does the question align with the logical sequence of MITRE ATT&CK tactics in the CTI outline?
+Logical: Does the question align with the logical sequence of MITRE ATT&CK tactics in the CTI outline?
 Does the question reference TTPs from the preceding or subsequent tactics in the CTI outline such that it logically leads to the correct answer?
 3.
-**Relevance**: Does the question directly relate to the CTI outline?
+Relevance: Does the question directly relate to the CTI outline?
 4.
-**Consistency**: Does the question align with the provided TTP Description?
+Consistency: Does the question align with the provided TTP Description?
 5.
-**Answer Consistency**: Can the question be fully answered using the correct answer, without any contradictions or inconsistencies?
-Your task is to iteratively refine the quality of the given question based on the feedback provided in the Feedback Results.
+Answer Consistency: Can the question be fully answered using the correct answer, without any contradictions or inconsistencies?
 [End of Evaluation Criteria] [Steps]: Please follow these steps: 1.
@@
 4.
-Output the Refined Question: - Please follow the output format: ”Refined Question: ¡Your refined question here¿”.
-[End of Steps] Following the steps above, please refine the question based on the Feedback Results and CTI Outline provided below.
+Output the Refined Question: - Please follow the output format: “Refined Question: <Your refined question here>.” [End of Steps] Following the steps above, please refine the question based on the Feedback Results and CTI Outline provided below.
 Please only output the refined question.
-C.3. Automatic Evaluation Prompt
-We utilize the definitions in the evaluation criteria in Table 7 to create prompts. We show an example prompt template for evaluating the Logical aspect of the question shown in Table 19.
-Table 19. Prompt template for evaluating the Logical aspect.
+C.3 Automatic Evaluation Prompt Templates
+We utilize the definitions in the evaluation criteria in Table LABEL:tab.annotation_instructions to create prompts.
+We show an example prompt template for evaluating the Logical aspect of the question shown in Box C.3.
+Box 8: Prompt template for evaluating the Logical aspect.
 You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework.
 [Inputs]: You will receive three parts: 1.
-**CTI Outline**: A structured account of a cyber attack, ordered by MITRE ATT&CK tactics.
+CTI Outline: A structured account of a cyber attack, ordered by MITRE ATT&CK tactics.
 Additional context under ”Others” provides background on the threat actor but is secondary.
 2.
-**Question with Answer Choices**: A question aimed at inferring a TTP from the attack sequence described in the CTI report, along with one correct answer and distractors among the answer choices.
+Question with Answer Choices: A question aimed at inferring a TTP from the attack sequence described in the CTI report, along with one correct answer and distractors among the answer choices.
 3.
-**Description to Correct Answer**: The description of the correct answer from the MITRE ATT&CK framework.
+Description to Correct Answer: The description of the correct answer from the MITRE ATT&CK framework.
 [End of Inputs] Your task is to rate the question on one metric below.
-[Definition of Logical Aspect] Evaluation Criteria: Logical (1-5): Does the question align with the logical sequence of MITRE ATT&CK tactics in the CTI outline?
+[Definition]
+Evaluation Criteria:
+Logical (1-5): Does the question align with the logical sequence of MITRE ATT&CK tactics in the CTI outline?
 Does the question reference TTPs from the preceding and/or subsequent tactics in the CTI outline such that it logically leads to the correct answer?
@@
 5 - Perfect Logical Alignment: The question perfectly aligns with the MITRE ATT&CK framework, referencing relevant TTPs in a way that naturally leads to the correct answer.
+[End of the Definition]
 [Steps:] Evaluation Steps: 1.
@@
 [End of Steps] Following the steps above, please evaluate the question and only output the numerical evaluation score.
-C.4. Answering Prompts
-The prompt templates for the three benchmarking tasks, Regular, Zero-Shot and RAG-empowered are shown in Tables 20, 21, 22 respectively.
-Table 20. Prompt template for Regular setting.
+C.4 Answering Prompt Templates
+The prompt templates for the three benchmark settings (i.e., Context setting, Zero-Shot setting and RAG-empowered setting) are shown in Box C.4, Box C.4, Box C.4 respectively.
+Box 9: Prompt template for the Context setting.
 You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework.
 [Inputs]: You will receive two parts: 1.
-A **CTI Report** that describe a cyber attack ordered by MITRE ATT&CK tactics.
+A CTI Report that describe a cyber attack ordered by MITRE ATT&CK tactics.
 Note that additional information labeled as “Others” provides context about the threat actor but is secondary.
 2.
-A **Question** about a sequence of TTPs with several answer choices.
+A Question about a sequence of TTPs with several answer choices.
 [End of Inputs]
@@
 Provide a Step-by-Step Reasoning and Final Answer: - Outline your reasoning step-by-step.
-- Conclude with the final answer in the following format: ”Final Answer: ¡insert answer choice here¿”.
-[End of Steps] Following the steps above, please answer the question below using the provided CTI report.
-Table 21. Prompt template for Zero-Shot setting.
+- Conclude with the final answer in the following format: “Final Answer: <insert answer choice here>.” [End of Steps] Following the steps above, please answer the question below using the provided CTI report.
+Box 10: Prompt template for the Zero-Shot setting.
 You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework.
 [Inputs]: You will receive a question about a sequence of TTPs with several answer choices.
+[End of Inputs]
 Your task is to determine which answer choice forms the most plausible sequence of TTPs based on the attack sequence described in the question.
-[End of Inputs]
 [Steps:] Please follow these steps: 1.
@@
 Provide a Step-by-Step Reasoning and Final Answer: - Outline your reasoning step-by-step.
-- Conclude with the final answer in the following format: ”Final Answer: ¡insert answer choice here¿”.
-[End of Steps] Following the steps above, please answer the question below.
-Table 22. Prompt template for RAG-empowered setting.
+- Conclude with the final answer in the following format: “Final Answer: <insert answer choice here>.” [End of Steps] Following the steps above, please answer the question below.
+Box 11: Prompt template for the RAG-empowered setting.
 You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework.
 [Inputs]: You will receive two parts: 1.
-A **Question** about a sequence of TTPs with several answer choices.
+A Question about a sequence of TTPs with several answer choices.
 2.
-A list of **Related TTPs** that are relevant to the question.
+A list of Related TTPs that are relevant to the question.
 [End of Inputs]
@@
 Provide a Step-by-Step Reasoning and Final Answer: - Outline your reasoning step-by-step.
-- Conclude with the final answer in the following format: ”Final Answer: ¡insert answer choice here¿”.
-[End of Steps] Following the steps above, please answer the question below.
+- Conclude with the final answer in the following format: “Final Answer: <insert answer choice here>.” [End of Steps] Following the steps above, please answer the question below.

### v3

@@
-AttackSeqBench: Benchmarking Large Language Models in Analyzing Attack Sequences within Cyber Threat Intelligence
-Abstract
+AttackSeqBench: Benchmarking the Capabilities of LLMs for Attack Sequences Understanding
+Abstract.
 Cyber Threat Intelligence (CTI) reports document observations of cyber threats, synthesizing evidence about adversaries’ actions and intent into actionable knowledge that informs detection, response, and defense planning.
@@
 To address this, we introduce AttackSeqBench, a benchmark designed to systematically evaluate LLMs’ reasoning abilities across the tactical, technical, and procedural dimensions of adversarial behaviors, while satisfying Extensibility, Reasoning Scalability, and Domain-dpecific Epistemic Expandability.
-We further benchmark 7 LLMs, 5 LRMs and 4 post-training strategies across the proposed 3 benchmark settings and 3 benchmark tasks within our AttackSeqBench to identify their advantages and limitations in such specific domain.
+We further benchmark 7 LLMs, 5 LRMs and 4 post-training strategies across 3 benchmark settings and 3 benchmark tasks within our AttackSeqBench to identify their advantages and limitations in such specific domain.
 Our findings contribute to a deeper understanding of LLM-driven CTI report understanding and foster its application in cybersecurity operations.
-1 Introduction
+Our code of benchmark construction and evaluation and the corresponding dataset are available at: https://github.com/hulkima/AttackSeqBench.
+Cyber Security, Cyber Threat Intelligence, Large Language Models
+copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXccs: Security and privacyccs: Computing methodologies Artificial intelligenceccs: Computing methodologies Machine learning approaches
+1. Introduction
 Amid rapid digital transformation, the increasing sophistication and diversity of cyber attacks have become a pervasive concern for cybersecurity globally (Duo et al., 2022).
@@
 Such ability is crucial in anticipating future malicious attack actions, particularly in multi-stage cyber attacks launched by Advanced Persistent Threats (APTs) (Li et al., 2022).
-As illustrated in Figure 1, we define the sequence of adversary behaviors as attack sequence (Al-Sada et al., 2025) to represent the execution flow of malicious actions across different stages of a cyber attack under the MITRE ATT&CK® framework (Strom et al., 2018).
-Building on this definition and the following key perspectives, we further delve into the suitability of LLMs in analyzing attack sequences.
-1) Extensibility: To address the ever-evolving threat landscape and the advancements of LLMs, the proposed benchmark must be extensible to incorporate attack sequences from newly observed CTI reports.
+Figure 1. Illustration of an example cyber attack sequence and our AttackSeqBench.
+[…]
+Against this backdrop, we examine the suitability of LLMs for analyzing attack sequences from the following three perspectives.
+1) Extensibility: To address the ever-evolving threat landscape and the advancements of LLMs, the proposed benchmark must be extensible to incorporate the attack sequences from newly observed CTI reports.
 2) Reasoning Scalability: Recently, Large Reasoning Models (LRMs) have demonstrated substantial advantages over conventional LLMs in multi-step reasoning tasks, such as coding and mathematical reasoning.
-However, existing CRU works have primarily focused on addressing CTI–related tasks via LLM, leaving the necessity of reasoning for attack sequence analysis largely unexplored.
-3) Domain-Specific Epistemic Expandability: LLMs exhibited limitations in factual reliability on knowledge-intensive tasks (Xu et al., 2024b), analogously, LLM-driven CRU, which requires specialized cybersecurity knowledge, is also subject to such limitations.
+However, existing CRU works have primarily focused on addressing CTI–related tasks via LLMs, leaving the necessity of reasoning within LRMs for attack sequence analysis largely unexplored.
+3) Domain-Specific Epistemic Expandability: LLMs have exhibited significant limitations in factual reliability on knowledge-intensive tasks (Xu et al., 2024c), analogously, LLM-driven CRU, which requires specialized cybersecurity knowledge, is also subject to such limitations.
 This requirement becomes particularly pronounced in attack sequence analysis, which necessitates a comprehensive understanding of adversarial behaviors to effectively reason multi-stage cyber attacks.
-Figure 1: Illustration an example cyber attack sequence and our AttackSeqBench.
-Building upon these perspectives, we introduce AttackSeqBench, a novel benchmark designed for comprehensive evaluation of LLMs in attack sequence analysis.
+Building upon these aforementioned perspectives, we introduce AttackSeqBench, a novel benchmark designed for comprehensive evaluation of LLMs in attack sequence analysis.
 Catering to Extensibility, we first construct attack sequences based on extensive real-world CTI reports, ensuring that the benchmark accurately reflects the complexity and diversity of Tactics, Techniques, and Procedures (TTPs) in cyber attacks performed by APTs.
@@
 Moreover, we also extend beyond the conventional Zero-Shot setting by introducing Context-based and RAG-empowered settings, which pertinently assess LLMs’ epistemic expandability when injecting domain-specific cybersecurity knowledge at the semantic and representation levels.
-Our contribution are as follows: (I) We introduce AttackSeqBench, the pioneering benchmark that systematically evaluates the ability of existing LLMs, LRMs, and post-training strategies to analyze attack sequences across diverse inference settings and multi-level tasks.
+Our contributions can be summarized as follows: (I) We introduce AttackSeqBench, a pioneering benchmark that systematically evaluates the ability of existing LLMs, LRMs, and post-training strategies to understand attack sequences across diverse settings and multi-level tasks.
 (cf.
@@
 Section 4.4 and Section 4.5)
-2 Dataset Construction and Verification
+2. Related Work
+Automating CTI Report Understanding.
+With the increasing demands of cybersecurity operations and the breakthrough of LLMs, researchers have progressively explored their applicability within CRU (Zhang et al., 2024a).
+For instance, prior works have showcased the remarkable capabilities of LLMs in interpreting TTPs from the ATT&CK Knowledge Base (KB), surpassing the performance of some of the fine-tuned LMs with cybersecurity data (Fayyazi et al., 2024).
+Meanwhile, another line of work proposes LLM-driven threat intelligence Knowledge Graph construction frameworks, which utilize the threat-related entities and relations to describe CTI reports in a structural manner (Huang and Xiao, 2024; Cheng et al., 2024).
+However, the extent to which LLMs can understand and reason about the precise relations between adversary behavioral sequences described in CTI reports remains largely under-explored.
+In our work, we perform a holistic evaluation on various pre-trained LLMs, LRMs and fine-tuned LLMs in attack sequence analysis, from deducing high-level tactics to procedures described in CTI reports.
+Benchmarking LLMs in Cybersecurity.
+Inspired by the remarkable open-world knowledge and complex inference ability within LLMs, various benchmarks have been proposed to evaluate its general capabilities in language understanding (Hendrycks et al., 2021b), math reasoning (Cobbe et al., 2021), code generation (Chen et al., 2021).
+Regarding the cybersecurity domain, researchers start to benchmark the abilities of LLMs under such specialized setting, such as ethical hacking and compliance (Liu, 2023; Tihanyi et al., 2024; Garza et al., 2023).
+Targeting CRU-related tasks, SEvenLLM (Ji et al., 2024) explores LLMs’ abilities in threat-related entities extraction and summarizes reports from security vendors.
+SecBench (Jing et al., 2025) evaluates the knowledge retention and logical reasoning abilities of existing pre-trained LLMs from multiple languages and dimensions.
+Meanwhile, CTIBench (Alam et al., 2024) introduces five benchmark tasks to explore the threat entity attribution and cause-tracing abilities of LLMs within the security context.
+However, these studies primarily rely on authoritative sources (e.g., textbooks, open standards) while overlooking real-world sources such as CTI reports.
+For instance, CTIBench solely incorporates a small-scale set of CTI reports in its dataset construction process for only one of its five benchmark tasks.
+Furthermore, these benchmarks remain insufficient for providing a comprehensive evaluation towards the LLMs’ ability to understand relations among adversarial behaviors described in CTI reports, thereby failing to accurately reflect their reasoning capabilities over attack sequences containing domain-specific semantics.
+In this paper, we construct attack sequences based on an extensive set of CTI reports, while emphasizing on the practical aspects of CRU, inferring various aspects of adversarial behaviors, in our proposed benchmark tasks.
+Figure 2. Overview of our automated Q&A dataset construction pipeline.
+3. Dataset Construction and Verification
+3.1. Problem Definitions
+Attack sequence understanding aims to convert the unstructural report into the structural formulation and further comprehend the sequential attack patterns of the structured threat intelligence knowledge.
+To achieve this, we define the attack sequence $S$ as the progression of adversarial behaviors described in a given CTI report, characterized by the logical order of TTPs based on their associated tactics within the ATT&CK KB.
+Formally, we utilize a 4-tuple to represent $S$ as $S=(T,E,P,O)$, where:
 Goal of our AttackSeqBench Our goal is to explore the capability of diverse types of LLMs in attack sequence understanding.
 Through comprehensive evaluation across various tasks and settings, we emphasize the strengths and limitations of existing LLMs, offering the promising yet underexplored directions.
-2.1 Problem Definitions
-CTI report understanding aims to convert the unstructural report into the structural formulation and further comprehend the sequential attack patterns of the structured threat intelligence knowledge.
-To achieve this, we define the attack sequence $S$ as the progression of adversarial behaviors described in a given CTI report, characterized by the logical order of TTPs based on their associated tactics within the ATT&CK KB.
-Formally, we utilize a 4-tuple to represent $S$ as $S=(T,E,P,O)$, where:
     Tactic Sequence $T$: An ordered list of ATT&CK tactics, such that $T=(t_{1},\ldots,t_{n})$, where $t_{k}$ is the $k$-th tactic in the sequence.
@@
     CTI Outline $O$: A textual summary of the organized TTPs based on the order of Tactic Sequence $T$, such that $O=(o_{1},\ldots,o_{n})$, where $o_{k}$ refers to the summarized text associated with tactic $t_{k}$.
-Figure 2: Overview of our automated QA dataset construction pipeline.
-2.2 Dataset Construction
+3.2. Dataset Construction
 As illustrated in Figure 2, we first construct attack sequences using the extracted TTPs and CTI outline from CTI reports.
-Then, we generate Q&A pairs based on the constructed attack sequences and refine them based on a tailored evaluate criteria before populating the QA dataset.
+Then, we generate Q&A pairs based on the constructed attack sequences and refine them based on a tailored evaluate criteria before populating the Q&A dataset of our AttackSeqBench.
 Attack Sequence Construction.
 To efficiently and massively extract threat intelligence from these unstructured reports, we utilize a set of 408 CTI reports from various security vendors (Cisco Talos Intelligence Group, 2025; Microsoft, 2025) to construct attack sequences that accurately reflect the behaviors of real-world APTs.
-Specifically, we utilize a LLM-based KG construction framework (Zhang et al., 2025c) to automatically parse CTI reports, extract TTPs from each chunk into three level, generate CTI outlines, and combin them to construct the attack sequences $S$.
+The gold standard for real-world attack sequences is elusive because CTI reports vary in quality and completeness and practitioners with different expertise workflows and threat modeling frameworks reconstruct sequences in inherently-divergent ways.
+Thjrerfore, we utilize a LLM-based Knowledge Graph (KG) construction framework (Zhang et al., 2025c) to automatically parse CTI reports, extract TTPs from each chunk into three level, generate CTI outlines, and combine them to construct the attack sequences $S$.
 Notably, we exclude CTI outlines which contains less than two ATT&CK tactics in attack sequence construction as they are unlikely to detail attack patterns observed in real-world cyber attacks.
@@
 For the Multiple-Choice Question (MCQ) tasks, we adopt a rule-based approach to select three choices as distractors.
-Specifically, we select a adjacent tactic of tactic $t_{k}$ within the Tactic Sequence $T$ (i.e., $t_{k+1}$ or $t_{k-1}$) and randomly select two tactics from the ATT&CK KB in AttackSeq-Tactic.
-Regarding AttackSeq-Technique, we follow the STARC annotation framework (Berzak et al., 2020) to define the selection rules with the given technique $e_{j,k}$: (1) The first technique belongs to the same tactic $t_{k}$ but not present in the given attack sequence, i.e., $e_{i,k}\notin E(t_{k})$; (2) The second technique is supported by the given attack sequence but belongs to another tactic $e_{j,j}\notin E(t_{k})$; (3) The third technique comes from a randomly chosen tactic that is not supported by the given attack sequence.
-Regarding the Yes-No Question tasks, we first instruct LLM to generate questions for each group of procedures within the attack sequence to construct the AttackSeq-Procedure-Yes.
-Next, we randomly sample its 70% questions to generate the negative question samples.
-Specifically, we design two types of Yes-to-No transferring strategies as follows: (1) Negation of temporal prepositions, i.e., changing “before” to “only after” and/or “after” to “only before”, such that the modified question contradicts the given attack sequence (Rajpurkar et al., 2018); (2) Replacement of the procedures in the question with another procedures that is not supported by the given attack sequence.
+Specifically, we select an adjacent tactic of tactic $t_{k}$ within the Tactic Sequence $T$ (i.e., $t_{k-1}$ or $t_{k+1}$) and randomly select two tactics from the ATT&CK KB in AttackSeqBench-Tactic.
+Regarding AttackSeqBench-Technique, we follow the STARC annotation framework (Berzak et al., 2020) to define the selection rules with the given technique $e_{j,k}$: (1) The first wrong option $e_{i,k}$ belongs to the same tactic $t_{k}$ but not present in the given attack sequence, i.e., $e_{i,k}\in E(t_{k})$; (2) The second wrong option $e_{j,l}$ is supported by the attack sequence but belongs to another tactic $e_{j,l}\notin E(t_{k})$; (3) The third wrong option comes from a randomly chosen tactic that is not supported by the attack sequence.
+Regarding the Yes-No Question tasks, we first instruct LLM to generate questions for each group of procedures within the attack sequence to construct the AttackSeqBench-Procedure-Yes.
+Next, we randomly sample its 70% questions to generate the negative samples.
+Specifically, we design two Yes-to-No transferring strategies as follows: (1) Negation of temporal prepositions, i.e., changing “before” to “only after” or “after” to “only before”, such that the modified question contradicts the given attack sequence (Rajpurkar et al., 2018); (2) Replacement of the procedures in the question with another procedures that is not supported by the given attack sequence.
+Table 1.
+Evaluation results of four sub-tasks on human evaluation and automatic evaluation.
+Here, Hum.
+Perf.
+= Human Performance, Ans.
+= Answerability, Con.
+= Consistency, Ans.
+Con.
+= Answer Consistency.
+Task Human Evaluation Automatic Evaluation
+Hum. Perf. Clarity Ans. Relevance Con. Ans. Con. Logic Clarity Ans. Relevance Con. Ans. Con. Logic
+AttackSeqBench-Tactic 0.51 4.36 4.30 4.56 4.46 4.44 4.45 4.65 4.52 4.84 4.65 4.76 4.79
+AttackSeqBench-Technique 0.71 4.21 4.09 4.45 4.44 4.41 4.40 4.40 4.10 4.63 4.39 4.59 4.62
+AttackSeqBench-Procedure 0.64 4.78 4.70 - 4.82 4.79 - 3.85 3.63 - 3.24 3.55 -
+Average 0.63 4.51 4.43 4.50 4.63 4.60 4.42 4.19 3.97 4.72 3.90 4.13 4.69
 Q&A Refinement.
@@
 Appendix A.2).
-Considering the foundational role of Answerability in benchmark design, we first instruct the LLM to assess whether each question satisfies this criterion—specifically, whether the CTI report provides direct evidence supporting a correct answer that is clearly preferable to alternatives.
+Considering the foundational role of Answerability, we first instruct the LLM to assess whether each question satisfies this criterion—specifically, whether the CTI report provides direct evidence supporting a correct answer that is clearly preferable to alternatives.
 Questions failing this requirement are discarded from the next step.
@@
 Appendix C.2).
-We repeat this three-step process once more to improve the quality of the questions, the questions with full numerical scores are added to our final QA dataset.
-After the Q&A refinement, the data volume of four sub-tasks in our AttackSeqBench reduce from 2,158/2,937/1,393/3,249 to 1,697/1,917/1,223/1,412, filtering out 35.82% of the original samples that cannot satisfy the defined selection criteria.
+We repeat this three-step process once more to improve the quality of the questions, the questions with full numerical scores are added to our final Q&A dataset.
+After the Q&A refinement, the data volume of four sub-tasks in our AttackSeqBench reduce from 2,158/2,937/4,642 to 1,697/1,917/2,635, filtering out 35.82% of the original samples that cannot satisfy the defined selection criteria.
 Additionally, we further illustrate the top-10 ATT&CK tactics and techniques within our dataset in Figure 7 (a) and 7 (b) respectively.
 The most frequent tactic and technique in the figure is associated with a key objective of APTs, highlighting the relevance of our Q&A dataset in capturing attack sequences based on real-world cyber attacks.
-Table 1: Evaluation results of four sub-tasks on human evaluation and automatic evaluation.
-Task Human Evaluation Automatic Evaluation
-Hum. Perf. Ans. Cla. Log Rel. Con. Ans. Con. Ans. Cla. Log Rel. Con. Ans. Con.
-Tactic 0.51 4.30 4.36 4.45 4.56 4.46 4.44 4.52 4.65 4.79 4.84 4.65 4.76
-Technique 0.71 4.09 4.21 4.40 4.45 4.44 4.41 4.10 4.40 4.62 4.63 4.39 4.59
-Procedure-Yes 0.74 4.88 4.70 4.88 5.00 4.81 4.94 4.02 4.06 4.61 4.47 3.78 3.89
-Procedure-No 0.56 4.55 4.84 - - 4.82 4.66 3.29 3.66 - - 2.77 3.25
-Average 0.63 4.45 4.53 4.57 4.67 4.63 4.61 3.98 4.19 4.67 4.64 3.90 4.12
-2.3 Dataset Evaluation
-LLMs demonstrate strong potential in solving complex tasks, but they inevitably exhibit even severer hallucinations, which has become a widely recognized concern in the research community.
-To address this, we adopt a hybrid approach towards evaluating the quality of the constructed QA dataset using the criteria defined in our Q&A refinement (cf.
-Section 2.2).
+$3.$ Dataset Evaluation
+LLMs show potential in solving complex tasks, but they inevitably exhibit severer hallucinations, which has become a widely recognized concern in the research community.
+To address this, we adopt a hybrid approach towards evaluating the quality of the constructed Q&A dataset using the criteria defined in our Q&A refinement (cf.
+Section 1).
 We design 5-point Likert scales for each of the evaluation criterion, where higher scores indicate better alignment.
@@
 We first randomly sample 35 questions from each sub-task to construct a question set for human evaluation.
-Three cybersecurity experts are then invited to answer and evaluate the quality of our Q&A dataset based on the six aspects defined in Section 2.2.
-Based on Table 1, we observe that the average Human Performance (abbreviated as Hum.
-Perf.) equals 0.63, suggesting that these questions is challenging and deducible even for individuals with domain expertise.
-Notably, AttackSeq-Procedure-No is derived from AttackSeq-Procedure-Yes through Yes-to-No transferring strategies, that is, its Logical and Relevance are inherently misaligned with the attack sequence, and we therefore do not evaluate these two aspects.
-Furthermore, the human evaluation shows consistently high average scores across all aspects, ranging from $4.45$ to $4.61$ out of $5$, indicating that the generated Q&A are easy to comprehend and well aligned with the attack sequences.
+Three cybersecurity experts are then invited to answer and evaluate the quality of our Q&A dataset based on the six aspects defined in Section 1.
+Based on Table 1, we observe that the average Human Performance equals 0.63, suggesting that these questions is challenging and deducible even for individuals with domain expertise.
+Notably, AttackSeqBench-Procedure-No in AttackSeqBench-Procedure is derived from AttackSeqBench-Procedure-Yes via Yes-to-No transferring strategies, where its Logical and Relevance are inherently deviate from the given attack sequence, and we therefore do not evaluate these two aspects.
+Furthermore, the human evaluation shows consistently high average scores across all aspects, ranging from $4.43$ to $4.63$ out of 55, indicating that the generated Q&A are easy to comprehend and well aligned with the attack sequences.
 Automatic Evaluation.
-To alleviate the laborious task of human evaluation, recent works (Zheng et al., 2023; Yao et al., 2024) have shown considerable effectiveness of LLM-as-a-Judge framework in aligning with human preferences within specific domain, including cybersecurity (Xu et al., 2024a).
+To alleviate the laborious task of human evaluation, recent works (Zheng et al., 2023; Yao et al., 2024) have shown considerable effectiveness of LLM-as-a-Judge framework in aligning with human preferences within specific domain (Xu et al., 2024b).
 We incorporate G-Eval (Liu et al., 2023), a Chain-of-Thought (CoT) (Wei et al., 2022) and form-filling paradigm, to systematically assess the quality of generated Q&A pairs.
 Specifically, we design individual prompts for each aspect in the evaluation criteria that includes its definition and the scoring guideline based on the same 5-point Likert scale in Human Evaluation (cf.
-Table LABEL:tab.annotation_instructions in Appendix A.2).
+Table 5 in Appendix A.2).
 Then we instruct GPT-4o rate the Q&A for each aspect based on the evaluation criterion and the correct answer from the ATT&CK KB (cf.
 Appendix C.3).
-Based on Table 1, we observe that Logical and Relevance are the highest rated aspects, reinforcing the LLM’s ability to construct questions that follow the logical order of attack sequences (cf.
-).
+Based on Table 1, we observe that Logical and Relevance are the highest rated aspects, reinforcing the LLM’s ability to construct questions that follow the logical order of attack sequences.
 The fact that automatic evaluation scores are lower than human evaluation scores further indicates that answering questions correctly in our AttackSeqBench is more challenging for LLMs than for domain experts.
-3 Benchmark and Experiments
-Figure 3: Overview of the three benchmark settings that exhibit varying levels of contextual information given to the LLM.
-3.1 Benchmark Settings
+4. Benchmark and Experiments
+4.1. Benchmark Settings
 As illustrated in Figure 3, we elaborate on three benchmark settings with varying levels of contextual information: (1) Zero-Shot setting, (2) Context setting, and (3) RAG-empowered setting.
 Zero-Shot Setting.
-Motivated by the significant zero-shot reasoning abilities of LLMs across several downstream tasks (Hou et al., 2024; Kojima et al., 2022), we directly utilize the system prompt and the Q&A pairs to evaluate the LLMs’ performance on three tasks based on their inherent knowledge.
+Motivated by the significant Zero-Shot reasoning ability of LLMs across several downstream tasks (Hou et al., 2024; Kojima et al., 2022), we directly utilize the system prompt and Q&A pairs to evaluate LLMs’ performance on three tasks based on their inherent knowledge.
+Figure 3. Overview of benchmark settings that exhibit varying levels of contextual information given to LLM.
 Context Setting.
-Considering the existing context-aware work (Ma et al., 2023; Jin et al., 2024), we also organize this context setting to evaluate LLMs’ Domain-Specific Epistemic Scalability at the semantic level.
-Here, we remove the corresponding summary of the ground truth tactic $t_{k}$ from the CTI outline $O$ to construct the masked CTI outline $O_{m}$, where $O_{m}\!=\!O\setminus o_{k}$.
+Considering the existing context-aware work (Ma et al., 2023; Jin et al., 2024), we also organize context setting to evaluate LLMs’ Domain-Specific Epistemic Scalability at the semantic level.
+Here, we remove the corresponding summarized text $o_{k}$ of the ground truth tactic $t_{k}$ from the CTI outline $O$ to construct the masked CTI outline $O_{m}$, where $O_{m}\!=\!O\setminus o_{k}$.
 Afterwards, the LLM will be instructed to use the corresponding $O_{m}$ to answer the question, highlighting its potential to perform abductive reasoning to determine the most plausible TTP in attack sequence.
 RAG-empowered Setting.
-Previous studies have demonstrated the Retrieval Augmented Generation (RAG) can significantly enhance the reliability pf LLMs and mitigate hallucinations (Zhang et al., 2025a).
+Previous studies have demonstrated the Retrieval Augmented Generation (RAG) can significantly enhance the reliability of LLMs and mitigate hallucinations (Zhang et al., 2025a).
 Here, we also design the RAG-empowered setting to evaluate the Domain-Specific Epistemic Scalability of LLMs at the representation level.
 This leverages the LLMs’ in-context learning ability to learn the associations between the entity in the question’s body and the relevant TTPs, thereby decomposing the problem and eliciting its stronger reasoning ability (Wu et al., 2022).
-3.2 Implementation Details
-To investigate the CRU capability of existing models, we evaluate seven LLMs (i.e., LLaMa-3.1-8B (Grattafiori et al., 2024), ChatGLM-4-9B (GLM et al., 2024), Qwen-2.5-3B, Qwen-2.5-14B, Qwen-2.5-32B (QwenTeam, 2024), Llama-3.3-70B (Grattafiori et al., 2024), and GPT-4o (OpenAI, 2024a) ) and five LRMs (i.e., R1 (Llama-3.1-8B), R1(Qwen-2.5-14B), R1(Qwen-2.5-32B) (DeepSeek-AI, 2025), QWQ-32B (Team, 2024) and GPT-o3-mini (OpenAI, 2025)) on AttackSeqBench.
-We also utilize four post-training strategies (i.e., SFT (Zhang et al., 2023), RD (Huang et al., 2024), RLIF (Zhao et al., 2025) and RLVR (DeepSeek-AI, 2025)) to embed security knowledge into LLMs to evaluate the Domain-Specific Epistemic Scalability of our AttackSeqBench (cf.
+4.2. Implementation Details
+To investigate the CRU ability of existing models, we evaluate seven LLMs (i.e., LLaMa3.1-8B (Grattafiori et al., 2024), ChatGLM4-9B (GLM et al., 2024), Qwen2.5-3B, Qwen2.5-14B, Qwen2.5-32B (QwenTeam, 2024), LLaMa3.3-70B (Grattafiori et al., 2024), and GPT-4o (OpenAI, 2024a) ) and five LRMs (i.e., R1 (LLaMa3.1-8B), R1 (Qwen2.5-14B), R1 (Qwen2.5-32B) (DeepSeek-AI, 2025), QWQ-32B (Team, 2024) and GPT-o3-mini (OpenAI, 2025)) on AttackSeqBench.
+We use four post-training strategies (i.e., SFT (Zhang et al., 2023), RD (Huang et al., 2024), RLIF (Zhao et al., 2025) and RLVR (DeepSeek-AI, 2025)) to embed cybersecurity knowledge into LLMs to evaluate the Domain-Specific Epistemic Scalability of AttackSeqBench (cf.
 Appendix A.4).
 Here, we measure the performance with accuracy Acc=n/MAcc\!=\!n/M, where $n$ is the correctly-answered number of questions and $M$ is the total number.
-[^1] Due to the page limitation, we introduce the complete implementation details in Appendix A.5 and A.6.
-Table 2: Performance comparison of various LLMs, LRMs and post-training strategies across three benchmark tasks and settings. Bold and underlined denote best and second-best in each column.
-LLMs AttackSeq-Tactic AttackSeq-Technique AttackSeq-Procedure
+[^2] We introduce the complete implementation details in Appendix A.5 and A.6.
+Table 2. Performance comparison of various LLMs, LRMs and post-training strategies across three benchmark tasks and settings. Bold and underlined denote best and second-best in each column.
+Versions Models AttackSeqBench-Tactic AttackSeqBench-Technique AttackSeqBench-Procedure
 Zero-Shot Context RAG Zero-Shot Context RAG Zero-Shot Context RAG
@@
 […]
-3.3 Performance Comparison
-Comparison between diverse groups of LLMs.
-As shown in Table 2, we notice that: Although LLMs generally follow the scaling laws in our AttackSeqBench, none of the LLMs consistently outperforms the others, and the optimal LLM varies diverse tasks.
-For instance, the best-performing models under the zero-shot setting across the three benchmark tasks are Qwen-2.5-32B, Qwen-2.5-32B, and GPT-o3-mini, respectively.
-This suggests that current models may not possess explicit security-specific knowledge, as relevant information in pretraining corpus is likely overshadowed by general-domain content.
-Moreover, most models consistently perform worst in AttackSeq-Tactic compared to the other two tasks, mirroring the human evaluation results in Section 2.3 and underscoring the common challenge faced by both human experts and LLMs in tactical inference.
-Furthermore, from Table 3, we can observe that compared to the zero-shot setting, all models exhibit substantial performance gains on AttackSeq-Procedure-No under the context setting, indicating the importance of contextual information in identifying highly implausible actions within attack sequences.
-As defined in Appendix A.3, AttackSeq-Procedure-No is inherently more complex and reasoning-demanding than AttackSeq-Procedure-Yes, as it requires models to overcome the helpful-only bias and explicitly answer ‘No’ to disprove the plausibility of procedures occurring within the attack sequence.
-This explains why LRMs with stronger reasoning ability outperform in AttackSeq-Procedure-No compared to other tasks, underscoring the benchmark’s emphasis on Reasoning Scalability.
-Finally, most post-training strategies substantially improve the performance of its base LLM, particularly in zero-shot settings that rely solely on internal knowledge.
+4.3. Performance Comparison
+Comparison between diverse groups of models.
+As shown in Table 2, we notice that: Although LLMs generally follow the scaling laws in our AttackSeqBench, none of them consistently outperforms the others, and the optimal LLM varies diverse tasks.
+For instance, the best-performing models under the Zero-Shot setting across three benchmark tasks are Qwen2.5-32B, Qwen2.5-32B, and GPT-o3-mini, respectively.
+This suggests that current models may not possess explicit security-specific knowledge, as relevant information in pre-training corpus is likely overshadowed by general-domain content.
+Moreover, most models consistently perform worst in AttackSeqBench-Tactic compared to other two tasks, mirroring the human evaluation results in Section 3.3 and underscoring the common challenge faced by both human experts and general LLMs in tactical inference.
+Furthermore, we can observe that compared to the Zero-Shot setting, all models exhibit substantial performance gains on AttackSeqBench-Procedure-No under the context setting, indicating the importance of contextual information in identifying highly implausible actions within attack sequences.
+As defined in Appendix A.3, AttackSeqBench-Procedure-No is inherently more complex and reasoning-demanding than AttackSeqBench-Procedure-Yes, as it requires models to overcome the helpful-only bias and explicitly answer ‘No’ to disprove the plausibility of procedures occurring within the attack sequence.
+This explains why LRMs with stronger reasoning ability outperform in AttackSeqBench-Procedure-No compared to other tasks, underscoring the benchmark’s emphasis on Reasoning Scalability.
+Finally, most post-training strategies substantially improve the performance of its LLM backbone, particularly in Zero-Shot settings that rely solely on internal knowledge.
 However, their performance still lags behind instructive LLMs equipped with task-adapted prompts.
-This highlights a promising direction: Designing specialized post-training strategy to embed security-related knowledge into existing LLMs, thereby advancing the development of domain-specific models for cybersecurity.
-Table 3: Performance comparison of two LLMs and two LRMs on AttackSeq-Procedure-Yes and AttackSeq-Procedure-No in each benchmark setting, where the bold value indicates the best performance of each column.
-LLMs AttackSeq-Procedure-Yes AttackSeq-Procedure-No
+This highlights a promising direction: Designing a specialized post-training strategy to embed security-related knowledge into existing LLMs, thereby advancing the development of domain-specific models for cybersecurity domain.
+Table 3. Performance comparison of four representative LLMs and LRMs under each setting, where the bold value indicates the best performance of each column.
+LLMs AttackSeqBench-Procedure-Yes AttackSeqBench-Procedure-No
 Zero-Shot Regular RAG Zero-Shot Regular RAG
-Llama-3.1-8B 0.9128 0.7572 0.8858 0.2434 0.6216 0.2111
+LLaMA3.1-8B 0.9128 0.7572 0.8858 0.2434 0.6216 0.2111
 GPT-4o 0.9469 0.9567 0.8831 0.4426 0.2698 0.4143
@@
 Comparing the performance of LLMs across three benchmark settings in Table 2, we can observe that: In general, the Context setting consistently outperforms Zero-Shot and RAG settings across most benchmark tasks, with the advantage more pronounced in larger LLMs.
-Taking the Qwen-2.5 series as an example: performance shifts from zero-shot being optimal in Qwen-2.5-3B (0.4467 vs.
+Taking the Qwen2.5 series as an example: performance shifts from Zero-Shot being optimal in Qwen2.5-3B (0.4467 vs.
 0.4614 vs.
-0.3296) to context-setting being optimal in Qwen-2.5-32B (0.6195 vs.
+0.3296) to context-setting being optimal in Qwen2.5-32B (0.6195 vs.
 0.5903 vs.
-0.5154) in the AttackSeq-Tactic task, with Qwen-2.5-14B showing the transition in between.
+0.5154) in AttackSeqBench-Tactic, with Qwen2.5-14B showing the transition in between.
 This phenomenon is reasonable as larger LLMs possess more extensive internal knowledge, and task-specific context further enhances their effectiveness and robustness within the specific domain.
@@
 We further investigate its limitation in Section 4.5.2.
-3.4 Robustness Analysis
-Figure 4: Parameter sensitivity analysis on (a) Temperature and (b) Max Output Tokens in AttackSeq-Tactic under the zero-shot setting.
-3.4.1 Parameter Sensitivity Analysis
-Regarding the parameter sensitivity, we investigate the impact of temperature and maximum output tokens on LLMs and illustrate them in Figure 4(a) and Figure 4(b) respectively.
+$4.$ Robustness Analysis
+4.4.1. Parameter Sensitivity Analysis
+Regarding the parameter sensitivity, we investigate the impact of temperature and maximum output tokens on LLMs and LRMs and illustrate them in Figure 4 (a) and Figure 4 (b) respectively.
 Firstly, we observe that increasing the temperature from 0 to 1 causes a sharp performance drop in smaller LLMs, while larger LLMs remain relatively unaffected in attack sequence analysis.
@@
 This highlights the importance of carefully tuning the maximum output tokens parameter to optimize performance in LRMs while considering the associated costs incurred (Wang et al., 2024).
-3.4.2 Computational Complexity Analysis
-Figure 5: Computational complexity analysis of seven LLMs and four LRMs in AttackSeq-Tactic under the regular setting.
-The size of bubble represents inference time, where zigzag lines denote LLMs and cross hatch lines indicate LRMs.
-As illustrated in Figure 5, we extensively compare the performance, model size and inference cost of several open-source LLMs and LRMs in AttackSeq-Tactic under the context setting.
-We can observe the following findings: (1) The performance trend of LLMs follows scaling law in our AttackSeqBench (Kaplan et al., 2020), particularly among LLMs within the same series; (2) Comparing the performance trend of LLMs with different thinking types, LRMs can not consistently outperform LLMs, yet they require significantly more inference time.
+Figure 4. Parameter sensitivity analysis on (a) Temperature and (b) Max Output Tokens in AttackSeqBench-Tactic.
+4.4.2. Computational Complexity Analysis
+As illustrated in Figure 5, we extensively compare the performance, model size and inference cost of several open-source LLMs and LRMs in AttackSeqBench-Tactic under the context setting.
+We can observe the following findings: (1) The performance trend of LLMs follows scaling law in our AttackSeqBench (Kaplan et al., 2020), particularly among LLMs within the same series.
+(2) Comparing the performance trend of LLMs with different thinking types, LRMs can not consistently outperform LLMs, yet they require significantly more inference time.
 This phenomenon is more pronounced when comparing LLMs and LRMs of the same model size, where distilled LRMs perform significantly worse than their base LLMs.
@@
 We further investigate this issue in Section 4.5.1.
-Table 4: A sample of reasoning path analysis between Qwen-2.5-32B (LLM) and R1(Qwen-2.5-32B) (LRM) on AttackSeq-Procedure-No under the zero-shot setting.
+Table 4.
+A sample of reasoning path analysis between Qwen2.5-32B and R1 (Qwen2.5-32B) on AttackSeqBench-Procedure under the Zero-Shot setting.
 We utilize different color to differentiate the diverse phases within a given attack sequence.
@@
 […]
-LLM: Qwen-2.5-32B LRM: R1 (Qwen-2.5-32B)
+LLM: Qwen2.5-32B LRM: R1 (Qwen2.5-32B)
 Breakdown of Attack Sequence: 1.
@@
 2.
-Generating Network Traffic: Interacting with the Transport Device Interface to generate network traffic could be a technique used to exfiltrate data or to communicate with a C2 server.
-However, this step seems out of place if the primary goal is to redirect the entry point of a driver.
+Generating Network Traffic: Interacting with the Transport Device Interface to generate network traffic could be a technique used to exfiltrate data or communicate with a C2 server.
+However, this step seems out of place if the goal is to redirect the entry point of a driver.
 3.
@@
 Wait
-(Transitional Word), but the question says ”only before redirecting.” So, does the TDI interaction happen only before the redirection, meaning it’s a one-time thing?
+(Transitional Word), but the question says "only before redirecting." So, does the TDI interaction happen only before the redirection, meaning it’s a one-time thing?
 Or is it part of the ongoing process?
@@
 (Incorrect Conclusion)
-3.5 In-depth Analysis
-3.5.1 Reasoning Path Analysis
-To explore the reason why existing LRMs perform worser than LLMs in attack sequence understanding, we provide a sample of reasoning path between Qwen-2.5-32B (LLM) and R1 (Qwen-2.5-32B) (LRM) on AttackSeq-Procedure-No under the Zero-shot setting in Table 4.
+4.5. In-depth Analysis
+4.5.1. Reasoning Path Analysis
+To explore the reason why LRMs perform worser than LLMs in our AttackSeqBench, we provide a sample of reasoning path between Qwen2.5-32B and R1 (Qwen2.5-32B) in Table 4.
 We observe that both LLM and LRM can decompose the attack sequence into granular TTPs and retrieve relevant knowledge.
@@
 This overthinking within LRMs are also more prone to construct redundant reasoning loops and further incur reasoning misalignment, which may amplify minor misunderstandings into incorrect conclusions.
-Figure 6: Error distribution of randomly-sampled 100 incorrect responses from GPT-4o in AttackSeq-Technique under the RAG-empowered setting.
-3.5.2 Effectiveness of RAG Strategies
-To investigate why LLMs underperform in the RAG-empowered setting of our AttackSeqBench, we collect a candidate set where GPT-4o answers correctly in the zero-shot setting but fails under this setting.
-Specifically, we randomly sample 100 incorrect responses in AttackSeq-Technique, and classify them into four categories as shown in Figure 6.
+Figure 5.
+Computational complexity analysis of seven LLMs and four LRMs in AttackSeqBench-Tactic under the regular setting.
+The size of bubble represents inference time, where zigzag lines denote LLMs and cross hatch lines indicate LRMs.
+Figure 6. Error distribution of randomly-sampled 100 incorrect responses from GPT-4o in AttackSeqBench-Technique under the RAG-empowered setting.
+4.5.2. Effectiveness of RAG Strategies
+To investigate why LLMs underperform in the RAG-empowered setting of our AttackSeqBench, we collect some candidates where GPT-4o answers correctly in Zero-Shot setting but fails under this setting.
+Specifically, we randomly sample 100 incorrect responses in AttackSeqbench-Technique, and classify them into four categories in Figure 6.
 These four categories are: (1) Factual Error, meaning that LLM’s prediction contradicts the ground truth despite the correct retrieved content; (2) Over-reliance (Xia et al., 2024), meaning that LLM excessively refers to the retrieved content and fails to synthesize the attack sequence in the given question; (3) Irrelevant Retrieved TTP, which refers to incorrect predictions due to irrelevant retrieval to the given question; (4) Incorrect Answer Format, which refers to LLM’s failure to follow the output format specified in the prompt template.
-Our analysis reveals that 59% of errors stem from Factual Error, where the primary cause is the model’s failure to effectively integrate retrieved evidence into the reasoning chain.
+Figure 6 reveals that 59% of errors stem from Factual Error, where the primary cause is the model’s failure to effectively integrate retrieved evidence into the reasoning chain.
 Rather than enhancing the inference process, the retrieved knowledge functions as noise to distort the output distribution, thereby inducing faulty reasoning and incorrect answers.
-Moreover, around 32% of the errors occur because LLMs treat retrieved knowledge as the absolute authority without validating them against the question intent or their internal knowledge.
-Consequently, the model often relies solely on correct but incomplete retrieval chunks, which leads to faulty results.
+Moreover, 32% of errors occur because LLMs treat retrieved knowledge as the absolute authority without validating them against the question intent.
+Consequently, the model often relies solely on correct but incomplete retrieved chunks, which leads to faulty results.
 Within the ATT&CK KB, the nuances of TTP descriptions introduce several overlaps and ambiguities, which account for 8% of cases where the embedding model to retrieve incorrect tactics and techniques.
-For example, technique T1574 – Hijack Execution Flow  [^2] https://attack.mitre.org/techniques/T1574/ is associated with three distinct tactics (i.e., Persistence, Privilege Escalation, and Defense Evasion), leading the model to misinterpret attack sequences in the given question.
-Enhancing the integration of retrieved knowledge with question intent and internal knowledge in RAG scenario, or investigating embedding methods capable of capturing fine-grained TTP semantics, holds promise for improving the effectiveness of attack sequence analysis.
-4 Related Work
-Automating CTI Report Understanding.
-With the increasing demands of cybersecurity operations and the breakthrough of LLMs, researchers have progressively explored their applicability within CRU (Zhang et al., 2024a).
-For instance, prior works have showcased the remarkable capabilities of LLMs in interpreting TTPs from the ATT&CK KB, surpassing the performance of some of the fine-tuned LMs with cybersecurity data (Fayyazi et al., 2024).
-Meanwhile, another line of work proposes LLM-driven threat intelligence Knowledge Graph construction frameworks, which utilize the threat-related entities and relations to describe CTI reports in a structural manner (Huang & Xiao, 2024; Cheng et al., 2024).
-However, the extent to which LLMs can understand and reason about the precise relations between adversary behavioral sequences described in CTI reports remains largely under-explored.
-In our work, we perform a holistic evaluation on various pre-trained LLMs, LRMs and fine-tuned LLMs in attack sequence analysis, from deducing high-level tactics to detailed procedures described in CTI reports.
-Benchmarking LLMs in Cybersecurity.
-Inspired by the remarkable open-world knowledge and complex inference ability within LLMs, various benchmarks have been proposed to evaluate its general capabilities in language understanding (Hendrycks et al., 2021b), math reasoning (Cobbe et al., 2021), code generation (Chen et al., 2021).
-Regarding the cybersecurity domain, researchers start to benchmark the abilities of LLMs under such specialized setting, such as ethical hacking and compliance (Liu, 2023; Tihanyi et al., 2024; Garza et al., 2023).
-Targeting CRU-related tasks, SEvenLLM (Ji et al., 2024) explores the abilities of LLMs in threat-related entities extraction and summarizing reports from security vendors.
-SecBench (Jing et al., 2025) evaluates the knowledge retention and logical reasoning abilities of existing pre-trained LLMs from multiple languages and dimensions.
-Meanwhile, CTIBench (Alam et al., 2024) introduces five benchmark tasks to explore the threat entity attribution and cause-tracing abilities of LLMs within the security context.
-However, these studies primarily rely on authoritative sources (e.g., textbooks, open standards) while overlooking real-world sources such as CTI reports.
-For instance, CTIBench solely incorporates a small-scale set of CTI reports in its dataset construction process for only one of its five benchmark tasks.
-Furthermore, these benchmarks remain insufficient for providing a comprehensive evaluation towards the LLMs’ ability to understand relations among adversarial behaviors described in CTI reports, thereby failing to accurately reflect their reasoning capabilities over attack sequences containing domain-specific semantics.
-In this paper, we construct attack sequences based on an extensive set of CTI reports, while emphasizing on the practical aspects of CRU, inferring various aspects of adversarial behaviors, in our proposed benchmark tasks.
-5 Conclusion
-The breakthrough of LLMs has shown promising potential across the cybersecurity domain, particularly in CTI understanding.
-Despite this, the applicability of LLMs in analyzing adversarial sequences remains largely unexplored.
+For example, technique T1574 – Hijack Execution Flow  [^3] https://attack.mitre.org/techniques/T1574/ is associated with three distinct tactics (i.e., Persistence, Privilege Escalation, and Defense Evasion), leading the model to misinterpret the given attack sequences.
+Enhancing the integration of retrieved knowledge with question intent, identifying SAE-derived feature to generate answers that aligned with retrieved evidence, or tuning the embedding models to capture more accurate TTP semantics, holds promise for improving the effectiveness of RAG in attack sequence understanding.
+5. Conclusion
+The breakthrough of LLMs has shown promising potential across the cybersecurity domain, particularly in CTI analysis.
+Despite this, the applicability of LLMs in understanding attack sequences remains largely unexplored.
 In this work, we propose AttackSeqBench, a benchmark tailored for assessing LLMs’ ability in understanding how adversaries operate through inferring TTPs based on attack sequences from real-world CTI reports.
@@
 Our work opens up a new direction towards LLM-driven CRU, enabling effective threat intelligence mining through automation.
+6. Limitations and Future works
+Limitations: While our work serves as a pioneering study in benchmarking LLMs’ capability in attack sequence understanding, several limitations should be acknowledged.
+Firstly, our study focuses on the correctness of LLMs’ responses through Multi-Choice Questions and Yes-No Questions, which may not fully capture the reasoning abilities of LLMs.
+Secondly, although we have conducted extensive experiments with seven LLMs, five LRMs, and four post-training strategies across three benchmark tasks (AttackSeqBench-Tactic, AttackSeqBench-Technique, and AttackSeqBench-Procedure) and three benchmark settings (Zero-Shot setting, Context setting, and RAG-empowered setting), fully demonstrating the Reasoning Scalability and Domain-Specific Epistemic Expandability of our AttackSeqBench, the implementations of RAG and post-training strategies remains relatively basic and leave room for future refinement.
+Thirdly, our AttackSeqBench currently leverages 408 rigorously filtered CTI reports to extract attack sequences and generate Q&A pairs.
+Although this number substantially exceeds prior CTI-related studies (i.e., 12 in AttacKG+ (Zhang et al., 2025c)  [^4] Notably, AttacKG+ just evaluates on only 12 CTI reports, the “500 CTI reports” are merely used to show extracted TTP distributions rather than method effectiveness., 12 in $M$-AttacKG (Zhang et al., 2026), and at most 71 in Attack Flow  [^5] https://center-for-threat-informed-defense.github.io/attack-flow/), the proposed Q&A dataset construction pipeline (source codes are publicly-available in our GitHub repo) is flexible and can be readily extended to unseen CTI reports.
+This not only demonstrates the Extensibility of our AttackSeqBench but also highlights an important direction for continuously refining this benchmark in future work.
+Nevertheless, while it is important to be aware of these limitations, our AttackSeqBench serves as a valuable benchmark to systematically explore LLMs’ reasoning abilities across the tactical, technical and, procedural dimensions of adversarial behaviors.
+Future works: Building on the limitations, our future research on AttackSeqBench will proceed along three directions.
+Regarding evaluation, we plan to expand our evaluation methods from the simple Multiple-Choice Question tasks and Yes-No Question tasks to the more complex reasoning and completion tasks, thereby providing a more comprehensive assessment of LLM capability in attack sequence understanding.
+In terms of methodology, we will build on AttackSeqBench to explore more fine-grained RAG approaches and advanced post-training strategies that account for the knowledge-extensive and high-stakes nature of CTI reports, aiming to fully leverage model potential in complex cyber-attack scenarios.
+At the data level, we will continue to expand and dynamically update the CTI corpus to ensure our AttackSeqBench remains evolvable over time, thereby supporting the steady advancement of domain-specific foundation models for cybersecurity.
 Ethics Statement
@@
 Furthermore, human evaluation is conducted with careful consideration of evaluator expertise and potential biases, ensuring fairness and reliability in assessment.
-Reproducibility Statement
-To promote reproducibility, we release an anonymous repository that contains all resources necessary to replicate our study, including the original CTI reports, the dataset construction pipeline, the complete datasets for the three tasks (i.e., AttackSeq-Tactic, AttackSeq-Technique, and AttackSeq-Procedure), and the code for running the three benchmark settings (i.e., Zero-Shot setting, Context setting, and RAG-empowered setting).
-In addition, we also provide the comprehensive implementation details about the four post-training strategies in Appendix.
-Together, these resources ensure that our AttackSeqBench can be reliably reproduced and the reported results independently verified.
-Appendices
 Appendix A Dataset
-A.1 Dataset Distribution
-Based on Figure 7, we observe that the top three most frequent tactics (i.e., Command and Control, Defense Evasion and Execution) occur in the middle of attack sequences, while the bottom two tactics (i.e., Exfiltration and Reconnaissance) occurs at the start and the end of the attack sequence.
+A.1. Dataset Distribution
+Based on Figure 7, we observe that the top-3 most frequent tactics (i.e., Command and Control, Defense Evasion and Execution) occur in the middle of attack sequences, while the bottom two tactics (i.e., Exfiltration and Reconnaissance) occurs at the start and the end of the attack sequence.
 Similarly, the most frequent ATT&CK technique is T1071-Application Layer Protocol  [^6] https://attack.mitre.org/techniques/T1071/, which is associated with the most common operations of APTs, Command and Control.
-Figure 7: Visualization of the distribution of top-10 tactics and techniques in AttackSeqBench.
-A.2 Dataset Evaluation Criteria
+Figure 7. Visualization of the distribution of top-10 tactics and techniques in AttackSeqBench.
+A.2. Dataset Evaluation Criteria
 Inspired by the existing work (Fu et al., 2024), we utilize the following six dimensions as the evaluation criteria to evaluate the quality of our constructed Q&A dataset:
@@
   • Answer Consistency. We check if the question can be fully answered by the correct answer, without any contradictions and inconsistencies.
-To quantitatively evaluate its quality, we first design the 5-point Likert scale for each aspect (refer to Table LABEL:tab.annotation_instructions), where each score corresponds to a different level of the given aspect.
+To quantitatively evaluate its quality, we first design the 5-point Likert scale for each aspect (refer to Table 5), where each score corresponds to a different level of the given aspect.
 Then we instruct three cybersecurity experts and LLM to provide the score of each aspect to achieve the human evaluation and the automatic evaluation, respectively.
@@
 This suggests that automatic evaluation is still limited in knowledge-intensive domains such as in cybersecurity.
-Note that for the AttackSeq-Procedure-No, we evaluate questions only on four aspects—Answerability, Clarity, Consistency, and Answer Consistency—since it is derived from AttackSeq-Procedure-Yes through negation of temporal prepositions and replacement of procedures.
-A.3 Benchmark Tasks
+Note that for the AttackSeqBench-Procedure-No, we evaluate questions only on four aspects—Answerability, Clarity, Consistency, and Answer Consistency—since it is derived from AttackSeqBench-Procedure-Yes through negation of temporal prepositions and replacement of procedures.
+A.3. Benchmark Tasks
 Inspired by existing LLM benchmarks in the general domain (Hendrycks et al., 2021a; Rein et al., 2023; Zhang et al., 2024b), we propose three tasks in the form of Multiple-Choice Questions and Yes-No Questions to evaluate the reasoning capabilities of LLMs in inferring TTPs present in attack sequences, where each task reflects a distinct aspect of adversarial behaviors.
-Table 5: Annotation instructions for each evaluation aspect.
+Table 5. Annotation instructions for each evaluation aspect.
 Aspects Instructions
@@
 Consistency Score 1: The question contradicts the TTP description or is entirely misaligned with the provided details.
-Consistency Score 2: The question loosely aligns with the TTP description but has inconsistencies or inaccuracies.
+Score 2: The question loosely aligns with the TTP description but has inconsistencies or inaccuracies.
 Score 3: The question mostly aligns with the TTP description but contains minor inconsistencies.
@@
 Score 5: The correct answer completely and unambiguously answers the question, with no contradictions or inconsistencies.
-AttackSeq-Tactic.
+AttackSeqBench-Tactic.
 This task evaluates the LLMs’ capability to infer a tactic $t_{k}\in T$.
 Given a question $Q$ that corresponds to tactic $t_{k}$ and four shuffled candidate tactics $C_{Tac}=\{c_{r}:r\in[1,4]\}$, the LLM will be instructed to select the correct tactic $c_{l}\in C_{Tac}$.
-AttackSeq-Technique.
+AttackSeqBench-Technique.
 This task assesses the LLMs’ capability to infer a technique $e_{j,k}\in E(t_{k})$.
 Given a question $Q$ that corresponds to $e_{j,k}$ and four shuffled candidate techniques $C_{Tec}=\{c_{r}:r\in[1,4]\}$, the LLM will be instructed to select the correct technique $c_{l}\in C_{Tec}$.
-AttackSeq-Procedure.
+AttackSeqBench-Procedure.
 This task challenges the LLMs’ capability to determine the likelihood of procedures $p_{m,j,k}\in P(e_{j,k})$ in an attack sequence.
 Given a question $Q$ and two candidate choices $C_{Pro}=\{yes,no\}$, the LLM will be instructed to determine if the procedure $p_{m,j,k}$ is likely to occur in the given attack sequence $S$.
-We further divide AttackSeq-Procedure into two sub-tasks, namely AttackSeq-Procedure-Yes and AttackSeq-Procedure-No, based on the ground truth of the boolean question.
+Given the AttackSeqBench-Procedure, we further divide it into two sub-tasks based on the ground truth of the boolean question, namely AttackSeqBench-Procedure-Yes and AttackSeqBench-Procedure-No.
 This explores the LLMs’ ability in determining misleading procedures that are unlikely to occur in an attack sequence.
-A.4 Baselines
+A.4. Baselines
 To demonstrate the effectiveness and robustness of our proposed AttackSeqBench, we evaluate seven large language models, five large reasoning models and four post-training strategies across three tasks involving different levels of data and three benchmark settings with varying context completeness.
@@
 In our experiments, we set the following sampling parameters while keeping the default value for the remaining parameters: temperature to 0, maximum output tokens to 2048, and top_p to 1.
-Below is the details of the utilized LLMs, LRMs and post-training strategies in our experiments:
+Below are the details of the utilized LLMs, LRMs and post-training strategies in our experiments:
 Large Language Models:
-  • LLaMa-3.1-8B (Grattafiori et al., 2024) is an instruction-tuned LLM from Meta, balancing performance and efficiency for textual understanding tasks.
-  • ChatGLM-4-9B (GLM et al., 2024) is pretrained on ten trillions of tokens and further achieve the high-quality alignment through supervised fine-tuning and human feedback learning.
-  • Qwen2.5-3B, Qwen2.5-14B, Qwen2.5-32B and Qwen2.5-72B (QwenTeam, 2024) represents Qwen-2.5 series LLMs with different parameter scales, demonstrating strong instruction-following and long-text generation capabilities.
-  • LLaMa-3.3-70B (Grattafiori et al., 2024) is an auto-regressive language model which is instruction-tuned in 70B with SFT and reinforcement learning with human feedback (RLHF).
+  • LLaMa3.1-8B (Grattafiori et al., 2024) is an instruction-tuned LLM from Meta, balancing performance and efficiency for textual understanding tasks.
+  • ChatGLM4-9B (GLM et al., 2024) is pretrained on ten trillions of tokens and further achieve the high-quality alignment through supervised fine-tuning and human feedback learning.
+  • Qwen2.5-3B, Qwen2.5-14B, Qwen2.5-32B and Qwen2.5-72B (QwenTeam, 2024) represents Qwen2.5 series LLMs with different parameter scales, demonstrating strong instruction-following and long-text generation capabilities.
+  • LLaMa3.3-70B (Grattafiori et al., 2024) is an auto-regressive language model which is instruction-tuned in 70B with SFT and reinforcement learning with human feedback (RLHF).
   • GPT-4o (OpenAI, 2024a) is one of the most advanced closed-source LLMs, which is a multi-lingual and multi-modal language model developed and functions well in real-time processing.
 Large Reasoning Models:
-  • DeepSeek-R1-Distill-Llama-8B (R1 (Llama-8B)), DeepSeek-R1-Distill-Qwen-14B (R1 (Qwen-14B)) and DeepSeek-R1-Distill-Qwen-32B (R1 (Qwen-32B)) (DeepSeek-AI, 2025) are fine-tuned from the Llama-3.1-8B, Qwen2.5-14B and Qwen2.5-32B with 800k samples curated with DeepSeek-R1, aiming to equip smaller models with reasoning capabilities like DeepSeek-R1.
+  • DeepSeek-R1-Distill-Llama-8B (R1 (LLaMa-8B)), DeepSeek-R1-Distill-Qwen-14B (R1 (Qwen-14B)) and DeepSeek-R1-Distill-Qwen-32B (R1 (Qwen-32B)) (DeepSeek-AI, 2025) are fine-tuned from the LLaMa3.1-8B, Qwen2.5-14B and Qwen2.5-32B with 800k samples curated with DeepSeek-R1, aiming to equip smaller models with reasoning capabilities like DeepSeek-R1.
   • QWQ-32B-Preview (QWQ-32B) (Team, 2024) is a preview release which gives LLM time to ponder, to question, and to reflect, enabling the deeper insight into complex problems.
@@
   • Reinforcement Learning with Verifiable Rewards (RLVR)  (DeepSeek-AI, 2025) leverages rule-based verification functions to provide reward signals for tasks with clear correctness criteria, enabling the optimization of LLMs while avoiding the complexities and potential pitfalls of reward models within RLHF.
-A.5 Post-training Corpus Construction
-Considering the post-training strategies, we construct two diverse datasets for SFT and RD, RLIF and RLVR respectively.
+A.5. Post-training Corpus Construction
+Considering the post-training strategies, we construct two diverse datasets for instruction-tuning (i.e., SFT) and reinforcement learning (i.e., RLVR, RD and RLIF) respectively.
 For the former, we utilize a subset of the Primus-Instruct dataset (Yu et al., 2025).
-Primus-Instruct is a cybersecurity corpus collected for instruction-tuning, containing diverse task types such as alert explanation, suspicious command analysis, security event query generation, retrieved security document QA, Terraform security mis-configuration repair, and general multi-turn instruction following.
+Primus-Instruct is a cybersecurity corpus collected for instruction-tuning, containing diverse task types such as alert explanation, suspicious command analysis, security event query generation, retrieved security document Q&A, Terraform security mis-configuration repair, and general multi-turn instruction following.
 To mitigate the inherent bias from linguistic inconsistencies, we filter out non-English samples via the FastText language identification library (Joulin et al., 2016) and manually verify the results, yielding a subset of 710 samples for SFT.
@@
 Following (Zhang et al., 2025b), we leverage transitional words (i.e., “but”, “however”, “wait”, etc.) as the proxy for inferability, and retain only the 3,890 samples containing at least ten such words when constructing the corpus for RD, RLIF and RLVR.
-A.6 Implementation Details
+A.6. Implementation Details
 To examine the performance of LLMs on our AttackSeqBench after embedding cybersecurity knowledge, and considering GPU constraints, we evaluate existing post-training strategies on Qwen-2.5-3B (QwenTeam, 2024) and LLaMA-3.1-8B (Grattafiori et al., 2024) under both full-parameter fine-tuning and parameter-efficient fine-tuning paradigms across all benchmark tasks and settings.
 Retrieval Augmented Generation (RAG): We first crawl the description and the example procedures of each technique in the Enterprise ATT&CK Matrix v17  [^8] https://attack.mitre.org/versions/v17/matrices/enterprise/.
-Then, we split the textual data into text chunks and embed the chunks in a vector store (i.e., Chroma DB  [^6] https://www.trychroma.com/), where each chunk’s metadata contains the associated ATT&CK tactic and technique.
-We utilize a hybrid retriever with a re-ranker, by combining combine Okapi BM25 (Robertson & Zaragoza, 2009) and a dense retriever based on the more advanced text-embedding-3-large from OpenAI (OpenAI, 2024b).
+Although alternative threat modeling approaches such as the Diamond Model and the Lockheed Martin Kill Chain exist, they are fundamentally abstractions that constrain cyber attacks, differing primarily in their modeling emphases.
+Here, we adopt the ATT&CK framework, as it is continuously maintained and underpins the majority of existing TTP extraction efforts.
+Then, we split the textual data into several chunks and embed them in a vector store (i.e., Chroma DB  [^9] https://www.trychroma.com/), where each chunk’s metadata contains the associated ATT&CK tactic and technique.
+We utilize a hybrid retriever with a re-ranker, by combining BM25 (Robertson and Zaragoza, 2009) as the sparse retriever and a dense retriever based on the more advanced text-embedding-3-large from OpenAI (OpenAI, 2024b).
 We set the chunk size to 512 and retrieved chunks to 3, and utilize LlamaIndex (Liu, 2022) to implement the retriever.
-Additionally, we also implement BGE-EN-ICL (Li et al., 2024) and ATT&CK-BERT (Abdeen et al., 2023a) within RAG to evaluate their effectiveness in attack sequence analysis.
+Additionally, we also implement BGE-EN-ICL (Li et al., 2024) and ATT&CK-BERT (Abdeen et al., 2023b) within RAG to evaluate their effectiveness in attack sequence analysis.
 Supervised Fine-tuning (SFT): We fine-tune the backbone LLM on the first dataset within Appendix A.5 using the LLaMA-Factory (Zheng et al., 2024) framework.
@@
 Thus, a single epoch struck a balance between adapting the LLM to the cybersecurity tasks while preserving its pre-trained general-purpose performance.
-Reasoning Distillation (RD) refers to fine-tune LLM on a reasoning dataset distilled from the advanced LRMs (i.e., DeepSeek-R1  and GPT-o1-preview ), which enables the smaller LLM to inherit the reasoning behaviors of the above LRMs.
+Reasoning Distillation (RD) refers to fine-tune LLM on a reasoning dataset distilled from the advanced LRMs (i.e., DeepSeek-R1 (DeepSeek-AI, 2025) and GPT-o1-preview (OpenAI, 2024a)), which enables the smaller LLM to inherit the reasoning behaviors of the above LRMs.
 For RD, we fine-tune our backbone LLM on the latter dataset within Appendix A.5 with the same parameter settings in SFT.
-Reinforcement Learning with Verifiable Rewards (RLVR) extends reinforcement learning by incorporating verifiable signals as rewards, such as correctness checks or logical consistency that can be programmatically validated.
-Specifically, we implement RLVR with Group Relative Policy Optimization (GRPO) (Shao et al., 2024) on the Volcano Engine Reinforcement Learning (verl) framework, where group-normalized rewards reduce variance and stabilize training.
-We conduct RLVR with the learning rate of $3\times 10^{-6}$ using Fully Sharded Data Parallel (FSDP).
 Reinforcement Learning with Internal Feedback (RLIF) (Zhao et al., 2025) enables LLMs to optimize policies using intrinsic signals without relying external supervision.
@@
 Here, we conduct RLIF on the same latter dataset within Appendix A.5 using the verl framework, adopting the same training hyperparameters and FSDP set-up as in RLVR.
-Figure 8: Comparison between existing CTI-related benchmarks and our AttackSeqBench based on a real-world CTI report (Koehl & Hannon, 2020).
+Reinforcement Learning with Verifiable Rewards (RLVR) extends reinforcement learning by incorporating verifiable signals as rewards, such as correctness checks or logical consistency that can be programmatically validated.
+Specifically, we implement RLVR with Group Relative Policy Optimization (GRPO) (Shao et al., 2024) on the Volcano Engine Reinforcement Learning (verl) framework, where the utilized reward combines an accuracy reward and a format reward for each query, and group-normalized rewards reduce variance and stabilize training.
+We conduct RLVR with the learning rate of $3\times 10^{-6}$ using Fully Sharded Data Parallel (FSDP).
+Figure 8.
+Comparison between existing CTI-related benchmarks and our AttackSeqBench based on a real-world CTI report (Koehl and Hannon, 2020).
 Our benchmark emphasizes on the sequences of adversary behaviors described in CTI reports.
-A.7 Related Benchmarks Comparison
+A.7. Related Benchmarks Comparison
 To demonstrate the uniqueness and novelty of our work, we illustrate the key differences between existing CTI-related benchmarks and our AttackSeqBench which significantly highlights attack sequence analyzing in Figure 1.
@@
 This raises the need to consider the sequential characteristics of a cyber attack within the given CTI report.
-A.8 The Use of Large Language Models
-We declare that LLMs were employed to assist with the refinement of this manuscript, specifically for grammar checking and language polishing.
-Additionally, LLMs were used in a limited capacity for minor debugging and syntactic correction of code snippets.
-Beyond these auxiliary roles, given that the primary purpose of this work is to explore the capability of diverse types of LLMs in understanding adversarial attack sequences, we also utilize a range of open-source and closed-source LLMs and LRMs during dataset construction, dataset refinement, and performance evaluation.
-All such uses are documented in the main paper and the appendix and were carefully controlled to ensure transparency and reproducibility.
-A.9 Limitations and Future works
-Limitations: While our work serves as a pioneering study into the LLMs’ reasoning capabilities in attack sequence analysis, several limitations should be acknowledged.
-Firstly, our study focuses on correctness of models’ responses through Multi-Choice Questions and Yes-No Questions, which may not fully capture the reasoning abilities of LLMs necessary for comprehensive evaluation.
-Secondly, although we have conducted extensive experiments with seven LLMs, five LRMs, and four post-training strategies across three benchmark tasks (AttackSeq-Tactic, AttackSeq-Technique, and AttackSeq-Procedure) and three benchmark settings (Zero-Shot setting, Context setting, and RAG-empowered setting), fully demonstrating the Reasoning Scalability and Domain-Specific Epistemic Expandability of our AttackSeqBench, the implementations of RAG and post-training strategies remains relatively basic and leave room for future refinement.
-Thirdly, our AttackSeqBench currently leverages 408 rigorously filtered CTI reports to extract attack sequences and generate Q&A pairs.
-Although this number substantially exceeds prior CTI-related studies (i.e., 12 in AttacKG+ (Zhang et al., 2025c), 12 in MM-AttacKG (Zhang et al., 2025d), and at most 71 in Attack Flow  [^7] https://center-for-threat-informed-defense.github.io/attack-flow/), the proposed dataset construction pipeline is flexible and can be readily extended to unseen CTI reports.
-This not only demonstrates the Extensibility of our AttackSeqBench but also highlights an important direction for continuously refining this benchmark in future work.
-Nevertheless, while it is important to be aware of these limitations, our AttackSeqBench serves are a valuable benchmark to systematically explore LLMs’ reasoning abilities across the tactical, technical and, procedural dimensions of adversarial behaviors.
-Future works: Building on the limitations, our future research on AttackSeqBench will proceed along three directions.
-Regarding evaluation, we plan to expand our evaluation methods from the simple Multiple-Choice Question tasks and Yes-No Question tasks to the more complex reasoning and completion tasks, thereby providing a more comprehensive assessment of model capabilities in CTI report understanding.
-In terms of methodology, we will build on AttackSeqBench to explore more fine-grained RAG approaches and advanced post-training strategies that account for the knowledge-extensive and high-stakes nature of CTI reports understanding, aiming to fully leverage model potential in complex cyber-attack scenarios.
-At the data level, we will continue to expand and dynamically update the CTI corpus to ensure our AttackSeqBench remains evolvable over time, thereby supporting the steady advancement of domain-specific foundation models for cybersecurity.
 Appendix B Experiments
-B.1 TTP Temporal Position Analysis
-To better understand the LLMs’ capabilities in attack sequence analysis, we conduct fine-grained analysis based on each stage within the attack sequences of MITRE ATT&CK®.
-We illustrate the performance of two LLMs and two LRMs across all benchmark tasks in the Regular setting in Figure 9 and show the corresponding values, the mean and standard deviation (SD) of these LLMs and LRMs on each tactic and benchmark task in Table 6.
-We identify four overachieving attack phases to categorize the ATT&CK tactics in attack sequences: (1) Initial Intrusion Phase; (2) Exploitation Phase; (3) Stealth Expansion Phase; (4) Objective Orchestration Phase.
-It is worth noting that our categorization follows Tactics, as each Technique and Procedure in MITRE ATT&CK® is uniquely mapped to a specific Tactic within a given attack sequence.
-Figure 9: The performance comparison between two LLMs and two LRMs on each tactic in each benchmark task.
-Table 6: Results of performance of two LLMs and two LRMs on all the 14 tactics, and the corresponding statistics of mean and standard deviation.
-Tactic
+B.1. TTP Temporal Position Analysis
+Figure 9. The performance comparison between two LLMs and two LRMs on each tactic in each benchmark task.
+Table 6. Results of performance of two LLMs and two LRMs on all the 14 tactics, and the corresponding statistics of mean and standard deviation.
+Tactics
 Reconna-
@@
 Exfiltr. Impact Mean SD
-Tactic Llama-3.1-8B 0.6170 0.3077 0.4406 0.5573 0.4260 0.3023 0.3867 0.4211 0.2551 0.5778 0.4390 0.6444 0.5915 0.4194 0.4561 0.1239
+AttackSeqBench -Tactic LLaMA3.1-8B 0.6170 0.3077 0.4406 0.5573 0.4260 0.3023 0.3867 0.4211 0.2551 0.5778 0.4390 0.6444 0.5915 0.4194 0.4561 0.1239
 GPT-4o 0.6383 0.5000 0.4368 0.7470 0.5207 0.4884 0.5430 0.5000 0.7245 0.7111 0.6951 0.7155 0.7042 0.6774 0.6144 0.1095
@@
 SD 0.0786 0.1132 0.0454 0.0961 0.0938 0.1434 0.1325 0.0792 0.2149 0.1066 0.1643 0.0362 0.0971 0.1697 - -
-Technique Llama-3.1-8B 0.6556 0.4878 0.5327 0.5890 0.7324 0.4667 0.5914 0.5625 0.6839 0.5179 0.6557 0.6884 0.7500 0.7241 0.6170 0.0943
+AttackSeqBench -Technique LLaMA3.1-8B 0.6556 0.4878 0.5327 0.5890 0.7324 0.4667 0.5914 0.5625 0.6839 0.5179 0.6557 0.6884 0.7500 0.7241 0.6170 0.0943
 GPT-4o 0.6222 0.6341 0.6262 0.6568 0.7254 0.6444 0.6512 0.6000 0.7513 0.6429 0.6148 0.7329 0.7632 0.6897 0.6682 0.0541
@@
 SD 0.0591 0.0755 0.0802 0.0624 0.0764 0.0907 0.1031 0.0927 0.0896 0.1153 0.0638 0.0557 0.0170 0.1072 - -
-Procedure Llama-3.1-8B 0.6634 0.5319 0.6489 0.6552 0.6667 0.5556 0.6809 0.6331 0.6987 0.6633 0.6027 0.6606 0.6906 0.6140 0.6404 0.0489
+AttackSeqBench -Procedure LLaMA3.1-8B 0.6634 0.5319 0.6489 0.6552 0.6667 0.5556 0.6809 0.6331 0.6987 0.6633 0.6027 0.6606 0.6906 0.6140 0.6404 0.0489
 GPT-4o 0.7030 0.6170 0.8351 0.7586 0.7469 0.6852 0.7325 0.7554 0.7067 0.7041 0.6712 0.7515 0.777 0.7193 0.7260 0.0523
@@
 SD 0.0260 0.0796 0.1139 0.0462 0.0368 0.0684 0.0355 0.0848 0.0473 0.0467 0.0714 0.0467 0.0530 0.0829 - -
-We summarize our findings into three key points: (1) Comparing the performance of three benchmark tasks, the overall highest standard deviation occurs within the AttackSeq-Tactic task for all LLMs and LRMs, suggesting that the LLMs’ and LRMs’ performance has the largest variance in this taskand further highlighting the challenge of inferring tactics within attack sequences; (2) In contrast, the standard deviation for AttackSeq-Procedure is the lowest for all LLMs, suggesting more consistency in the LLMs’ and LRMs’ performance among the different attack sequence stages; (3) GPT-4o consistently outperforms other LLMs and LRMs across most benchmark tasks, particularly in the Exploitation Phase and the Stealth Expansion Phase.
+To better understand the LLMs’ capabilities in attack sequence understanding, we conduct fine-grained analysis based on each stage within the attack sequences of MITRE ATT&CK®.
+We illustrate the performance of two LLMs and two LRMs across all benchmark tasks in the Regular setting in Figure 9 and show the corresponding values, the mean and standard deviation (SD) of these LLMs and LRMs on each tactic and benchmark task in Table 6.
+We identify four overachieving attack phases to categorize the ATT&CK tactics in attack sequences: (1) Initial Intrusion Phase; (2) Exploitation Phase; (3) Stealth Expansion Phase; (4) Objective Orchestration Phase.
+It is worth noting that our categorization follows Tactics, as each Technique and Procedure in MITRE ATT&CK® is uniquely mapped to a specific Tactic within a given attack sequence.
+We summarize our findings into three key points: (1) Comparing the performance of three benchmark tasks, the overall highest standard deviation occurs within the AttackSeqBench-Tactic task for all LLMs and LRMs, suggesting that the LLMs’ and LRMs’ performance has the largest variance in this task and further highlighting the challenge of inferring tactics within attack sequences; (2) In contrast, the standard deviation for AttackSeqBench-Procedure is the lowest for all LLMs, suggesting more consistency in the LLMs’ and LRMs’ performance among the different attack sequence stages; (3) GPT-4o consistently outperforms other LLMs and LRMs across most benchmark tasks, particularly in the Exploitation Phase and the Stealth Expansion Phase.
 However, the remaining LLMs achieve comparable performance in the remaining two phases, especially in tactics at the start and end of the attack sequences (e.g., Reconnaissance and Impact).
 This suggests that TTPs within rarely observed tactics are more challenging for the LLM and LRMs to interpret.
-B.2 Case Study
-To better understand the superiority and limitation of the advanced LLMs in analyzing attack sequences, we run GPT-4o multiple times to collect both successful and failed responses to the same questions under the identical experimental setting, which are presented in Table 7.
-For the Wrong case, GPT-4o incorrectly associates the technique of choice A, i.e., T1574-Hijack Execution Flow, with the Execution tactic.
-Although this technique indeed involves the execution of malicious payloads based on the ATT&CK knowledge base, it is only associated to the following three tactics: Persistence, Privilege Escalation, and Defense Evasion.
-This incorrect interpretation suggests that GPT-4o struggles in distinguishing the inherent ambiguity found in TTP descriptions, thereby affecting their ability to analyze attack sequences.
-Regarding the Right one, GPT-4o correctly identifies T1129-Shared Modules as the most plausible technique, which belongs to the Execution tactic  [^8] https://attack.mitre.org/techniques/T1129/ and serves as the executable files that are loaded into processes to provide access to execute malicious payloads.
-By selecting this option, GPT-4o demonstrates its ability to reason over the attack sequence: after WMI discovery detects the presence of antivirus, shared modules would facilitate the execution of “FinalBot.exe” to bypass basic defenses.
-This correct interpretation is beneficial to effectively link the ambiguous textual cues with the appropriate tactic/technique entities, thereby improving its reliability in analyzing attack sequences.
-Table 7: Error analysis of GPT-4o in AttackSeq-Technique, where GPT-4o incorrectly associates technique T1574-Hijack Execution Flow with the Execution tactic.
+Table 7. Error analysis of GPT-4o in AttackSeqBench-Technique, where GPT-4o incorrectly associates technique T1574-Hijack Execution Flow with the Execution tactic.
 Q & A Cases LLM Response
@@
 (Correct Conclusion)
-Table 8: Performance comparison between three embedding models (abbreviate as Emb. M.).
-#Emb. M. BGE OPENAI ATT&CK-BERT
+B.2. Case Study
+To better understand the superiority and limitation of the advanced LLMs in analyzing attack sequences, we run GPT-4o multiple times to collect both successful and failed responses to the same questions under the identical experimental setting, which are presented in Table 7.
+For the Wrong case, GPT-4o incorrectly associates the technique of choice A, i.e., T1574-Hijack Execution Flow, with the Execution tactic.
+Although this technique indeed involves the execution of malicious payloads based on the ATT&CK knowledge base, it is only associated to the following three tactics: Persistence, Privilege Escalation, and Defense Evasion.
+This incorrect interpretation suggests that GPT-4o struggles in distinguishing the inherent ambiguity found in TTP descriptions, thereby affecting their ability to analyze attack sequences.
+Regarding the Right one, GPT-4o correctly identifies T1129-Shared Modules as the most plausible technique, which belongs to the Execution tactic  [^10] https://attack.mitre.org/techniques/T1129/ and serves as the executable files that are loaded into processes to provide access to execute malicious payloads.
+By selecting this option, GPT-4o demonstrates its ability to reason over the attack sequence: after WMI discovery detects the presence of antivirus, shared modules would facilitate the execution of “FinalBot.exe” to bypass basic defenses.
+This correct interpretation is beneficial to effectively link the ambiguous textual cues with the appropriate tactic/technique entities, thereby improving its reliability in analyzing attack sequences.
+Table 8. Performance comparison between three embedding models (abbreviate as Emb. M.).
+#Emb. M. BGE-EN-ICL OPENAI ATT&CK-BERT
 Tasks Tactic Technique Procedure Tactic Technique Procedure Tactic Technique Procedure
-Llama-3.1-8B 0.4751 0.5974 0.5243 0.4838 0.6103 0.5435 0.4820 0.5980 0.5334
+LLaMA3.1-8B 0.4751 0.5974 0.5243 0.4838 0.6103 0.5435 0.4820 0.5980 0.5334
 GPT-4o 0.5522 0.6860 0.6319 0.5616 0.6578 0.6482 0.5687 0.7016 0.6216
-R1 (Llama-8B) 0.4905 0.5740 0.5226 0.4932 0.5696 0.5150 0.4651 0.5804 0.5104
-GPT o3-mini 0.5115 0.5853 0.6474 0.5192 0.5827 0.6474 0.5245 0.5874 0.6414
-B.3 Impact of Embedding Models within RAG-empowered Setting
+R1 (LLaMA3.1-8B) 0.4905 0.5740 0.5226 0.4932 0.5696 0.5150 0.4651 0.5804 0.5104
+GPT-o3-mini 0.5115 0.5853 0.6474 0.5192 0.5827 0.6474 0.5245 0.5874 0.6414
+B.3. Impact of Embedding Models within RAG-empowered Setting
 The semantics within CTI reports contain a large volume of domain-specific technical terminologies, where the accuracy of retrievers in identifying the most relevant tactics, techniques and procedures critically influences LLMs’ performance in RAG-empowered settings.
-To further examine this issue and mitigate the potential knowledge bias introduced by BGE-EN-ICL, we incorporate two additional embedding models into our RAG-empowered setting, namely OpenAI’s text-embedding-3-large (OpenAI, 2024b) and a domain-adapted ATT&CK BERT  (Abdeen et al., 2023b) fine-tuned on the specific cybersecurity data.
+To further examine this issue and mitigate the potential knowledge bias introduced by BGE-EN-ICL, we incorporate two additional embedding models into our RAG-empowered setting, namely OpenAI’s text-embedding-3-large (OpenAI, 2024b) and a domain-adapted ATT&CK BERT  (Abdeen et al., 2023a) fine-tuned on the specific cybersecurity data.
 We conduct performance comparisons between two representative LLMs and two LRMs based on these above embedding models in Table 8.
@@
 Appendix C Prompt Templates
-C.1 Question Generation Prompt Templates
-We utilize few-shot prompt templates for question generation of each of our benchmark tasks (i.e., AttackSeq-Tactic, AttackSeq-Technique, AttackSeq-Procedure-Yes, AttackSeq-Procedure-No), the corresponding prompts are shown in Box C.1, Box C.1, Box C.1, Box C.1 respectively.
-Box 1: Few-shot prompt for question generation for AttackSeq-Tactic.
+C.1. Question Generation Prompt Templates
+We utilize few-shot prompt templates for question generation of each of our benchmark tasks (i.e., AttackSeqBench-Tactic, AttackSeqBench-Technique, AttackSeqBench-Procedure-Yes, AttackSeqBench- Procedure-No), the corresponding prompts are shown in Box C.4, Box C.4, Box C.4, Box C.4 respectively.
+C.2. Dataset Refinement Prompt Templates
+The prompt templates to filter based on the Answerability criteria is in Box C.4, while the feedback and refinement prompts are in Box C.4 and Box C.4 respectively.
+C.3. Automatic Evaluation Prompt Templates
+We utilize the definitions in the evaluation criteria in Table 5 to create prompts. We show an example prompt template for evaluating the Logical aspect of the question shown in Box C.4.
+C.4. Answering Prompt Templates
+The prompt templates for the three benchmark settings (i.e., Context setting, Zero-Shot setting and RAG-empowered setting) are shown in Box C.4, Box C.4, Box C.4 respectively.
+Box 1: Few-shot prompt for question generation for AttackSeqBench-Tactic.
 You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework.
@@
 - Please follow the output format: “Question: <insert question here> Answer: <insert answer here> [End of Steps] Following the steps above, please generate a question based on the CTI report and ATT&CK tactic given below.
-Box 2: Few-shot prompt for question generation for AttackSeq-Technique.
+Box 2: Few-shot prompt for question generation for AttackSeqBench-Technique.
 You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework.
@@
 [End of Steps] Following the steps above, please generate a question based on the CTI report and ATT&CK tactic and technique given below.
-Box 3: Few-shot prompt for question generation for AttackSeq-Procedure-Yes.
+Box 3: Few-shot prompt for question generation for AttackSeqBench-Procedure-Yes.
 You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework.
@@
 Your task is to generate a question about the attack sequence based on the MITRE ATT&CK tactics found in the CTI report, the question should focus on inferring the given list of procedures based on the given MITRE ATT&CK tactic and technique.
-The answer to the question is ”Yes”, indicating that the given list of procedures is likely to occur in the attack sequence.
+The answer to the question is "Yes", indicating that the given list of procedures is likely to occur in the attack sequence.
 [Steps]: Please follow these steps: 1.
@@
 Instead, include the TTPs in the tactic that precedes before and/or follows after the given MITRE ATT&CK tactic based on the order of tactics in the CTI report.
-- Ensure that the answer to the question is ”Yes”.
+- Ensure that the answer to the question is "Yes".
 - The question should be concise, clear, and targeted towards experienced cybersecurity professionals.
@@
 [End of Steps] Following the steps above, please generate a question based on the CTI report and ATT&CK tactic, technique and procedures given below.
-Box 4: Few-shot prompt for question generation for AttackSeq-Procedure-No.
+Box 4: Few-shot prompt for question generation for AttackSeqBench-Procedure-No.
 You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework.
@@
 1.
-A Reference Question-Answer Pair that focuses on the logical sequence of TTPs in a CTI report, note that the answer to the this question is always ”Yes”.
+A Reference Question-Answer Pair that focuses on the logical sequence of TTPs in a CTI report, note that the answer to the this question is always "Yes".
 2.
@@
 [End of Inputs]
-Your task is to generate two questions based on the given reference question, such that attack sequence described in the question is modified and the correct answer to the two questions is ”No”.
+Your task is to generate two questions based on the given reference question, such that attack sequence described in the question is modified and the correct answer to the two questions is "No".
 The definitions of the two questions are as follows: 1.
-Question 1 should negate the ”before” and/or ”after” clauses of the reference question, such that the attack sequence contradicts the original sequence of TTPs in the reference question.
+Question 1 should negate the "before" and/or "after" clauses of the reference question, such that the attack sequence contradicts the original sequence of TTPs in the reference question.
 2.
@@
 Construct the Questions: - Design two questions that modify the attack sequence described in the reference question.
-- Ensure that the answer to the questions is ”No”.
+- Ensure that the answer to the questions is "No".
 - The question should be concise, clear, and targeted towards experienced cybersecurity professionals.
@@
 Please only provide the final output of the two questions.
-C.2 Dataset Refinement Prompt Templates
-The prompt templates to filter based on the Answerability criteria is in Box C.2, while the feedback and refinement prompts are in Box C.2 and Box C.2 respectively.
-Box 5: Prompt template for verifying Answerability during Self-Refinement.
+Box 5: Prompt template for verifying Answerability via Self-Refinement.
 You are a cybersecurity expert with deep knowledge of Cyber Threat Intelligence (CTI) reports and the MITRE ATT&CK framework.
@@
 CTI Outline: A structured account of a cyber attack, ordered by MITRE ATT&CK tactics.
-Additional context under ”Others” provides background on the threat actor but is secondary.
+Additional context under "Others" provides background on the threat actor but is secondary.
 2.
@@
 CTI Outline: A structured account of a cyber attack, ordered by MITRE ATT&CK tactics.
-Additional context under ”Others” provides background on the threat actor but is secondary.
+Additional context under "Others" provides background on the threat actor but is secondary.
 2.
@@
 [End of Inputs]
-Your task is to evaluate the QA pair and provide your feedback for each of the criteria defined below: [Evaluation Criteria]:
+Your task is to evaluate the Q&A pair and provide your feedback for each of the criteria defined below: [Evaluation Criteria]:
 Please refer to the definition of each feedback criterion:
@@
 CTI Outline: A structured account of a cyber attack, ordered by MITRE ATT&CK tactics.
-Additional context under ”Others” provides background on the threat actor but is secondary.
+Additional context under "Others" provides background on the threat actor but is secondary.
 2.
@@
 Please only output the refined question.
-C.3 Automatic Evaluation Prompt Templates
-We utilize the definitions in the evaluation criteria in Table LABEL:tab.annotation_instructions to create prompts.
-We show an example prompt template for evaluating the Logical aspect of the question shown in Box C.3.
 Box 8: Prompt template for evaluating the Logical aspect.
@@
 CTI Outline: A structured account of a cyber attack, ordered by MITRE ATT&CK tactics.
-Additional context under ”Others” provides background on the threat actor but is secondary.
+Additional context under "Others" provides background on the threat actor but is secondary.
 2.
@@
 [End of Steps] Following the steps above, please evaluate the question and only output the numerical evaluation score.
-C.4 Answering Prompt Templates
-The prompt templates for the three benchmark settings (i.e., Context setting, Zero-Shot setting and RAG-empowered setting) are shown in Box C.4, Box C.4, Box C.4 respectively.
 Box 9: Prompt template for the Context setting.